### v1

# Partial Regularity for the three-dimensional Stochastic Ericksen–Leslie equations

Abstract.
In this article, we investigate the existence of martingale suitable weak solutions to 3D stochastic Ericksen–Leslie equations driven by an external random force. Despite the inherent white noise, we prove that paths of martingale suitable weak solutions exhibit singular points of one-dimensional parabolic Hausdorff measure zero. Our approach is mainly based on two ideas: a local energy-preserving approximation which serves to establish the existence of suitable weak solutions, and a blow-up argument employed to achieve the partial regularity.
1. Introduction
Liquid crystal, an anisotropic phase situated between the isotropic liquid phase and solid phase, has long intrigued scientists from various disciplines. Among many types of liquid crystal states, the nematic phase is particularly notable for its characteristic that in this kind of material, the rod-shaped molecules display the orientational order (tending to point in the same direction) without positional order. The hydrodynamics of nematic liquid crystals has been a subject of extensive research, see for instance [25, 38, 49, 50]. A fundamental mathematical model for describing hydrodynamics in nematic liquid crystal flows is provided by the Ericksen–Leslie equations [16, 17, 33].
The Ericksen–Leslie equations combine momentum balance, typically captured by the Navier–Stokes equation [21, 45, 46], with the evolution of molecular orientation, characterized by the heat flow of harmonic maps [36]. In this article, we focus on the simplified version of the full Ericksen–Leslie equations, which was initially studied by Lin and Liu [34, 35], driven by a random force:
(1.1) $\left\{\begin{array}[]{l}\,\mathrm{d}{{\bf{u}}}-\Delta{{\bf{u}}}\,\mathrm{d}t+{{\bf{u}}}\cdot\nabla{{\bf{u}}}\,\mathrm{d}t+\nabla P\,\mathrm{d}t=-\mathop{\rm div}\nolimits(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}})\,\mathrm{d}t+\,\mathrm{d}W,\\ \mathop{\rm div}\nolimits{{\bf{u}}}=0,\\ \,\mathrm{d}{{\bf{d}}}-\Delta{{\bf{d}}}\,\mathrm{d}t+{{\bf{u}}}\cdot\nabla{{\bf{d}}}\,\mathrm{d}t=-f_{\epsilon}({{\bf{d}}})\,\mathrm{d}t.\end{array}\right.$
Here the unknown triple $({{\bf{u}}},P,{{\bf{d}}})$ represents the fluid velocity field, the pressure function, and the director field which characterizes the mean orientation of molecules, and $f_{\epsilon}({{\bf{d}}})=\nabla_{{\bf{d}}}F_{\epsilon}({{\bf{d}}})$, where
$F_{\epsilon}({{\bf{d}}})=\frac{1}{\epsilon^{2}}(|{{\bf{d}}}|^{2}-1)^{2}.$
It is evident that (1.1)${}_{1,2}$ is a forced Stokes system with the so-called Ericksen stress tensor $(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}})_{ij}=\partial_{x_{i}}{{\bf{d}}}\cdot\partial_{x_{j}}{{\bf{d}}}$ and a random force $W$ which will be specified later, while (1.1)${}_{3}$ is a transported semi-linear heat flow associated to the Ginzburg–Landau potential function $F_{\epsilon}({{\bf{d}}})$. The ultimate goal is to study the singular limit as $\epsilon\to 0$ in (1.1)${}_{3}$, which leads to the following transported heat flow of harmonic maps for the director field ${\bf{d}}$:
$\,\mathrm{d}{{\bf{d}}}-\Delta{{\bf{d}}}\,\mathrm{d}t+{{\bf{u}}}\cdot\nabla{{\bf{d}}}\,\mathrm{d}t=|\nabla{{\bf{d}}}|^{2}{{\bf{d}}}\,\mathrm{d}t.$
It is well-established that the critical dimension for heat flow of harmonic maps is two [8, 9]. Consequently, in 2D, it has been confirmed that solutions to the deterministic (limited) Ericksen–Leslie equation are partially regular [37], and singularities could occur [26, 30]. Moreover, the singular limit as $\epsilon\to 0$ has been justified through concentration-cancellation compactness [29], as also discussed in [14]. However, in 3D, the well-posedness for the deterministic (limited) Ericksen–Leslie equation remains an open problem with only a few partial results available, such as those presented in [39, 47].
Despite vast success in studies of the deterministic Ericksen–Leslie equations, it is worth noting that real-world scenarios introduce complexities that a deterministic framework could possibly overlook, including environmental disturbances [10], measurement uncertainties, and inherent thermal fluctuations [3, 31]. To address these intricacies, a typical way is to introduce a stochastic component into the Ericksen–Leslie equations, which leads to the formulation of the stochastic Ericksen–Leslie (SEL) equations. A probable consequence of the stochastic force in the nematic liquid crystal flows is its potential impact on the emergence of singularities, which can manifest as either turbulence in fluid flow or defects in molecular alignments. These phenomena carry substantial implications, impacting our understanding of theory and the practical applications of liquid crystal materials.
To this end, in this article, we study the SEL equations on the three-dimensional tori $\mathbb{T}^{3}$ driven by additive white noise. Despite a recent growing interest in the SEL equations, to our best knowledge, there have been no reported results regarding the regularity of the SEL equations in three dimensions. This research gap persists even when $\epsilon$ in (1) is held fixed rather than tending towards zero, as mentioned in [4, 5]. It’s worth mentioning that well-posedness results concerning the (limited) SEL equations in two dimensions can be found in [12] (with additive noise), and in [48] (with multiplicative noise); for further details, one may refer to [6, 16, 44]. Here our main objectives are two-fold. Firstly, we introduce the concept and establish the existence of martingale suitable weak solutions to (1.1). Since $\epsilon$ plays no role in our analysis, from now on we fix $\epsilon=1$ and omit it from $F_{\epsilon}$ and $f_{\epsilon}$. It is worth noting that the concept of martingale solutions is equivalent to that of statistical solutions, a topic extensively studied in statistical fluid dynamics (see for instance [22, 23, 24]). Secondly, we demonstrate that, with certain assumptions on the white noise, the set of singular points, where solutions became unbounded, cannot have a positive one-dimensional parabolic Hausdorff measure.
The article is organized as follows. In Section 2, we specify notations and stochastic framework to set up the SEL equations. Then in Section 3 we introduce a notion of suitable weak solutions, drawing inspiration from [20, 42] in the context of the stochastic Navier–Stokes equation. This approach involves splitting solutions to the SEL equations into two parts. The first part is governed by linear stochastic Stokes equations, and the second part is subject to the remaining nonlinear coupled system. Then in Section 4, we prove the existence of martingale suitable weak solutions. Finally in Section 5 we establish the partial regularity of martingale suitable weak solutions.
2. Mathematical settings and notations
2.1. Functional spaces
We introduce two fundamental spaces that take into account the incompressibility condition (1.1)${}_{2}$ (c.f. [46]):
$H:=\{{\bf{u}}\in L^{2}(\mathbb{T}^{3};\mathbb{R}^{3})|\mathop{\rm div}\nolimits{\bf{u}}=0\},$
and
$V:=\{{\bf{u}}\in H^{1}(\mathbb{T}^{3};\mathbb{R}^{3})|\mathop{\rm div}\nolimits{\bf{u}}=0\}.$
The operator $A:D(A)\subset H\to H$ is defined as $Au=-P_{L}\Delta{\bf{u}}$, where $P_{L}$ is the Helmholtz–Leray projection from $L^{2}(\mathbb{T}^{3};\mathbb{R}^{3})$ onto $H$ and $D(A)=H^{2}(\mathbb{T}^{3};\mathbb{R}^{3})\cap V$. It is well-known that $A^{-1}$ is well-defined from $H$ to $D(A)$. It is also self-adjoint. By the spectral theory of compact self-adjoint operators in a Hilbert space, there exist a sequence of eigenvalues $0<\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{m}\leq\cdots$, $\lambda_{m}\to\infty$ as $m\to\infty$ and an orthonormal basis $\{e_{i}\}_{i=1}^{\infty}$ such that $Ae_{i}=\lambda_{i}e_{i}$. The fractional power $A^{\alpha}$ of $A$, $\alpha\geq 0$, are simply defined by
$A^{\alpha}{\bf{u}}=\sum_{i=1}^{\infty}\lambda_{i}^{\alpha}e_{i}({\bf{u}},e_{i})_{L^{2}(\mathbb{T}^{3})},$
with domain
$D(A^{\alpha})=\left\{{\bf{u}}\in H\Big{|}\|{\bf{u}}\|_{D(A^{\alpha})}^{2}=\sum_{i=1}^{\infty}\lambda^{2\alpha}_{i}({\bf{u}},e_{i})_{L^{2}(\mathbb{T}^{3})}^{2}=\|A^{\alpha}{\bf{u}}\|_{L^{2}(\mathbb{T}^{3})}^{2}<\infty\right\}.$
Also we note that $D(A^{\alpha})\subset H^{2\alpha}(\mathbb{T}^{3};\mathbb{R}^{3})$. For the definition of $H^{2\alpha}(\mathbb{T}^{3};\mathbb{R}^{3})$, we refer to [41, Definition 1.10].
2.2. Stochastic framework
In order to define the noise term, we first recall some basic concepts and notations of probability and stochastic processes. For more details see for instance [11, 18, 19, 20, 40, 42].
We fix a probability space $(\Omega,\mathcal{F},\mathbb{P})$, where $\Omega$ is a sample space equipped with a $\sigma$-algebra $\mathcal{F}$, and a probability measure $\mathbb{P}$. Then we define a stochastic basis
(2.1) $\mathfrak{B}:=\left(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P},\left(W(t)\right)_{t\geq 0}\right),$
which is a filtered probability space with a Brownian motion $\left(W(t)\right)_{t\geq 0}$ adapted to the filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$. In order to avoid unnecessary complications below we may assume that $\{\mathcal{F}_{t}\}_{t\geq 0}$ is complete and right continuous (see [40]). We then assume that
(2.2) $\left(W(t)\right)_{t\geq 0}$ takes value in $D(A^{\delta})$ for some $\delta>0$ with a covariance operator $\mathcal{L}$.
Here $\mathcal{L}\in\mathcal{B}(D(A^{\delta}))$ and is non-negative and of trace class. Let $\{b_{i}\}_{i\geq 1}$ be the orthonormal basis of $D(A^{\delta})$ consisting of eigenvectors of $\mathcal{L}$ subject to non-negative eigenvalues $\{\gamma_{i}\}_{i\geq 1}$. Then we have the following expansion (c.f. [40, Theorem 4.20])
$W(t)=\sum_{i}W_{i}(t)b_{i},$
where the real-valued Brownian motions
$W_{i}(t)=(W(t),b_{i})_{D(A^{\alpha})}$
are independent and their covariances are
$\mathbb{E}[W_{i}(t)W_{i}(s)]=(t\wedge s)\gamma_{i}.$
Moreover, by Kolmogorov’s continuity criterion (see for instance [43]), we have that for every $\beta\in(0,1/2)$, it holds that
(2.3) $W(t)\in C^{\beta}\left([0,T];D(A^{\delta})\right)\quad\mathbb{P}-a.s..$
3. Martingale suitable weak solution
We will assume throughout the paper the following assumptions for the initial data
(3.1) $({{\bf{u}}}_{0},{{\bf{d}}}_{0})\in H\times H^{1}(\mathbb{T}^{3};\mathbb{R}^{3}),\quad|{{\bf{d}}}_{0}(x)|=1\text{ for all }x\in\mathbb{T}^{3}.$
As a preparation to introduce notions of deterministic and stochastic suitable weak martingale solutions, we will first split the solutions ${{\bf{u}}}$ and $P$ properly. We introduce new variables ${{\bf{u}}}={{\bf{z}}}+{{\bf{v}}}$, $P={Q}+{\pi}$ where $({{\bf{z}}},{Q})$ solves the linear stochastic Stokes system as follows
(3.2) $\left\{\begin{array}[]{l}{\,\mathrm{d}}{{\bf{z}}}-\Delta{{\bf{z}}}{\,\mathrm{d}}t+\nabla{Q}{\,\mathrm{d}}t={\,\mathrm{d}}W,\\ \mathop{\rm div}\nolimits{{\bf{z}}}=0,\\ {{\bf{z}}}(0)=0.\end{array}\right.$
While the triple $({{\bf{v}}},{\pi},{{\bf{d}}})$ solves the modified Ericksen–Leslie equations as follows
(3.3) $\left\{\begin{array}[]{l}\partial_{t}{{\bf{v}}}-\Delta{{\bf{v}}}+({{\bf{z}}}+{{\bf{v}}})\cdot\nabla({{\bf{z}}}+{{\bf{v}}})+\nabla{\pi}=-\mathop{\rm div}\nolimits\left(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}|\nabla{{\bf{d}}}|^{2}{\rm I}_{3}-F({{\bf{d}}}){\rm I}_{3}\right),\\ \mathop{\rm div}\nolimits{{\bf{v}}}=0,\\ \partial_{t}{{\bf{d}}}-\Delta{{\bf{d}}}+({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{d}}}=-f({{\bf{d}}}),\\ {{\bf{v}}}(0)={{\bf{u}}}_{0},{{\bf{d}}}(0)={{\bf{d}}}_{0}.\end{array}\right.$
Remark 3.1.
Due to the incompressibility condition (3.3)${}_{2}$, the right-hand side of (3.3)${}_{1}$ is equivalent to (${}_{1}$, since for the extra term we have
$\displaystyle\mathop{\rm div}\nolimits\left(\frac{1}{2}|\nabla{{\bf{d}}}|^{2}{\rm I}_{3}+F({{\bf{d}}}){\rm I}_{3}\right)=\nabla\left(\frac{1}{2}|\nabla{{\bf{d}}}|^{2}+F({{\bf{d}}})\right),$
which then can be treated as part of the pressure gradient. On the other hand, direct computations yield
(3.4) $-\mathop{\rm div}\nolimits\left(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}|\nabla{{\bf{d}}}|^{2}{\rm I}_{3}-F({{\bf{d}}}){\rm I}_{3}\right)=-\nabla{\bf{d}}\cdot(\Delta{\bf{d}}-f({\bf{d}})).$
Let ${{\bf{z}}}(t)$ be the stochastic process that solves (3.2). In the mild form we have
${{\bf{z}}}(t)=\int_{0}^{t}e^{-(t-s)A}\,\mathrm{d}W(s)=\sum_{i}^{\infty}\int_{0}^{t}e^{-(t-s)A}b_{i}\,\mathrm{d}W_{i}(s).$
By the Doob inequality and the Burkholder–Davis–Gundy inequality, for $p>1$, $\delta>0$, and $\varepsilon>0$ we can estimate
(3.5) $\mathbb{E}\left[\sup_{0\leq t\leq T}\|{{\bf{z}}}(t)\|_{D(A^{\delta+\frac{1}{2}-\varepsilon})}^{p}\right]$ $=\mathbb{E}\left[\sup_{0\leq t\leq T}\|A^{\delta+\frac{1}{2}-\varepsilon}{{\bf{z}}}(t)\|_{L^{2}(\mathbb{T}^{3})}^{p}\right]$
$\leq C(p)\mathbb{E}\left(\sum_{i=1}^{\infty}\gamma_{i}\int_{0}^{T}\|A^{\frac{1}{2}-\varepsilon}e^{-(T-s)A}A^{\delta}b_{i}\|_{L^{2}(\mathbb{T}^{3})}^{2}\,\mathrm{d}s\right)^{\frac{p}{2}}$
$\leq C(p)\mathbb{E}\left(\sum_{i=1}^{\infty}\gamma_{i}\int_{0}^{T}\|A^{\frac{1}{2}-\varepsilon}e^{-(T-s)A}\|^{2}_{\mathcal{B}(L^{2}(\mathbb{T}^{3}))}\|A^{\delta}b_{i}\|_{L^{2}(\mathbb{T}^{3})}^{2}\,\mathrm{d}s\right)^{\frac{p}{2}}$
[…]
where $C(p)$ is a constant that only depends on $p$. This implies that for a.s. $\omega$, ${{\bf{z}}}(\omega)\in L^{\infty}(0,T;D(A^{\delta+\frac{1}{2}-\varepsilon}))$. In particular, if we set $\delta=\frac{1}{4}+\beta$, $\beta>\varepsilon$, then by the Sobolev embedding
$D(A^{\frac{3}{4}+\beta-\varepsilon})\subset H^{\frac{3}{2}+2\beta-2\varepsilon}(\mathbb{T}^{3})\subset L^{\infty}(\mathbb{T}^{3}),$
based on which we have the following lemma which ensures the boundedness of ${{\bf{z}}}(\omega)$:
Lemma 3.2.
[20, Lemma 2.2] Fix $\omega\in\Omega$ with $W(\omega)\in C^{\frac{1}{2}-\varepsilon}([0,\infty);D(A^{\frac{1}{4}+\beta}))$ for some $\beta>\varepsilon>0$, then
(3.6) ${{\bf{z}}}(\omega)\in L_{\rm loc}^{\infty}(\mathbb{T}^{3}\times[0,\infty)).$
Definition 3.1.
Let $T>0$ and $\omega\in\Omega$ such that ${\bf{z}}(\omega)$ satisfies (3.6). Then we say $({{\bf{v}}},{{\bf{d}}})$ is a weak solution to (3.3) in the sense of distribution on $Q_{T}:=\mathbb{T}^{3}\times(0,T]$ subject to ${{\bf{v}}}(0)={{\bf{u}}}_{0},{{\bf{d}}}(0)={{\bf{d}}}_{0}$, and for any $\phi_{1}\in C_{0}^{\infty}(\mathbb{T}^{3},\mathbb{R}^{3})$, $\mathop{\rm div}\nolimits\phi_{1}=0$, $\phi_{2}\in C_{0}^{\infty}(\mathbb{T}^{3},\mathbb{R}^{3})$, and $t>0$, the following weak formulation of (3.3) is satisfied:
$\int_{\mathbb{T}^{3}\times{\{t\}}}{{\bf{v}}}\cdot\phi_{1}-$ $\int_{0}^{t}\int_{\mathbb{T}^{3}}{{\bf{v}}}\cdot\Delta\phi_{1}+({{\bf{z}}}+{{\bf{v}}})\otimes({{\bf{z}}}+{{\bf{v}}}):\nabla\phi_{1}$
$=\int_{\mathbb{T}^{3}}{\bf{u}}_{0}\cdot\phi_{1}+\int_{0}^{t}\int_{\mathbb{T}^{3}}\left(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}|\nabla{{\bf{d}}}|^{2}{\rm I}_{3}-F({{\bf{d}}}){\rm I}_{3}\right):\nabla\phi_{1},$
$\int_{\mathbb{T}^{3}\times{\{t\}}}{{\bf{d}}}\cdot\phi_{2}-$ $\int_{0}^{t}\int_{\mathbb{T}^{3}}{{\bf{d}}}\cdot\Delta\phi_{2}+({{\bf{z}}}+{{\bf{v}}})\otimes{\bf{d}}:\nabla\phi_{2}=\int_{\mathbb{T}^{3}}{\bf{d}}_{0}\cdot\phi_{2}-\int_{0}^{t}\int_{\mathbb{T}^{3}}f({\bf{d}})\cdot\phi_{2}.$
Now we are ready to introduce the definition of suitable weak solutions:
Definition 3.2.
Let $T>0$ and $\omega\in\Omega$ such that ${\bf{z}}(\omega)$ satisfies (3.6). Then a suitable weak solution to (1.1) is a triple $({{\bf{u}}},P,{{\bf{d}}})$ such that if ${{\bf{v}}}={{\bf{u}}}-{{\bf{z}}}$ and ${\pi}=P-{Q}$ where $({{\bf{z}}},{Q})$ is the solution to (3.2), then
  1. (1)
    $({{\bf{v}}},{{\bf{d}}})$ is weakly continuous with respect to time,
  2. (2)
    ${{\bf{v}}}\in L^{\infty}(0,T;H)\cap L^{2}(0,T;V)$, ${{\bf{d}}}\in L^{\infty}(0,T;H^{1}(\mathbb{T}^{3};\mathbb{R}^{3}))\cap L^{2}(0,T;H^{1}(\mathbb{T}^{3};\mathbb{R}^{3}))$ and ${\pi}\in L^{5/3}(0,T;L^{5/3}(\mathbb{T}^{3};\mathbb{R})),$
  3. (3)
    $({{\bf{v}}},{{\bf{d}}})$ is a weak solution to (3.3) in the sense of distribution on $Q_{T}$ as specified in Def. 3.1,
  4. (4)
    for all $t\leq T$ and almost every $s<t$,
    $\displaystyle\int_{\mathbb{T}^{3}\times\{t\}}\left(\frac{1}{2}|{{\bf{v}}}|^{2}+\frac{1}{2}|\nabla{{\bf{d}}}|^{2}+F({{\bf{d}}})\right)+\int_{s}^{t}\int_{\mathbb{T}^{3}}|\nabla{{\bf{v}}}|^{2}+|\Delta{{\bf{d}}}-f({{\bf{d}}})|^{2}$
    $\displaystyle\leq\int_{\mathbb{T}^{3}\times\{s\}}\left(\frac{1}{2}|{{\bf{v}}}|^{2}+\frac{1}{2}|\nabla{{\bf{d}}}|^{2}+F({{\bf{d}}})\right)$
    $\displaystyle+\int_{s}^{t}\int_{\mathbb{T}^{3}}{{\bf{z}}}\cdot[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{v}}}]+({{\bf{z}}}\cdot\nabla){{\bf{d}}}\cdot(\Delta{{\bf{d}}}-f({{\bf{d}}})).$
  5. (5)
    for any $\varphi\in C_{0}^{\infty}(Q_{T})$, $\varphi\geq 0$,
    $\displaystyle\int_{\mathbb{T}^{3}\times\{T\}}\left(\frac{|{{\bf{v}}}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}|^{2}}{2}+F({{\bf{d}}})\right)+\int_{Q_{T}}(|\nabla{{\bf{v}}}|^{2}+|\Delta{{\bf{d}}}|^{2}+|f({{\bf{d}}})|^{2})\varphi$
    $\displaystyle\leq\int_{Q_{T}}\left(\frac{|{{\bf{v}}}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}|^{2}}{2}+F({{\bf{d}}})\right)\partial_{t}\varphi+\int_{Q_{T}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{\bf{v}}]\cdot{\bf{z}}\varphi$
    $\displaystyle+\int_{Q_{T}}\left(\frac{|{{\bf{v}}}|^{2}}{2}+{\bf{v}}\cdot{\bf{z}}\right)[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla\varphi]+{\pi}{{\bf{v}}}\cdot\nabla\varphi+\left(\frac{|{{\bf{v}}}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}|^{2}}{2}\right)\Delta\varphi$
    $\displaystyle+\int_{Q_{T}}\left(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}|\nabla{\bf{d}}|^{2}{\rm I}_{3}\right):\nabla^{2}\varphi$
    $\displaystyle+\int_{Q_{T}}({{\bf{z}}}\cdot\nabla{{\bf{d}}})\cdot(\Delta{{\bf{d}}}-f({{\bf{d}}}))\varphi+[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{d}}}]\cdot(\nabla\varphi\cdot\nabla){{\bf{d}}}$
    $\displaystyle-\int_{Q_{T}}f({{\bf{d}}})\cdot(\nabla\varphi\cdot\nabla){{\bf{d}}}-2\int_{Q_{T}}\nabla f({{\bf{d}}}):\nabla{{\bf{d}}}\varphi.$
Next, we introduce the notions of martingale suitable weak solutions.
Definition 3.3.
A martingale suitable weak solution to (1.1) is a quadruple $(\widetilde{\mathfrak{B}},\widetilde{{\bf{u}}},\widetilde{P},\widetilde{{\bf{d}}})$, where $\widetilde{\mathfrak{B}}$ is a stochastic basis
(3.7) $\widetilde{\mathfrak{B}}=(\widetilde{\Omega},\widetilde{\mathcal{F}},\{\widetilde{\mathcal{F}}_{t}\}_{t\geq 0},\widetilde{\mathbb{P}},(\widetilde{W}(t))_{t\geq 0}),$
and $(\widetilde{W}(t))_{t\geq 0}$ is a Brownian motion adapted to the filtration {~$\{\widetilde{\mathcal{F}}_{t}\}_{t\geq 0}$ with values in $D(A^{\delta})$ and covariance operator $\mathcal{L}$, such that
$(\widetilde{{\bf{u}}}(\cdot),\widetilde{P}(\cdot),\widetilde{{\bf{d}}}(\cdot)):\widetilde{\Omega}\times[0,\infty)\rightarrow H\times\mathbb{R}\times H^{1}(\mathbb{T}^{3};\mathbb{R}^{3})$
is an ~$\widetilde{\mathcal{F}}_{t}$ adapted process, and
$\widetilde{\omega}\in\widetilde{\Omega}\to(\widetilde{{\bf{u}}}(\widetilde{\omega}),\widetilde{P}(\widetilde{\omega}),\widetilde{{\bf{d}}}(\widetilde{\omega}))\in L^{2}(0,T;H)\times L^{5/3}(Q_{T})\times L^{2}(0,T;H^{1}(\mathbb{T}^{3};\mathbb{R}^{3}))$
is a measurable mapping, and there exists a set $\widetilde{\Omega}_{0}\subset\widetilde{\Omega}$ of full probability such that on this set the triple $(\widetilde{{\bf{u}}}(\cdot,\widetilde{\omega}),\widetilde{P}(\cdot,\widetilde{\omega}),\widetilde{{\bf{d}}}(\cdot,\widetilde{\omega}))$ is a suitable weak solution in the sense of Definition 3.2.
To state the main result of our paper, we introduce the notion of singular points for a suitable weak solution.
Definition 3.4.
We call a point $z=(x,t)\in\mathbb{T}^{3}\times(0,\infty)$ regular if there exists a neighborhood of $z$ where $({{\bf{v}}},\nabla{{\bf{d}}})$ is essentially bounded. The other points will be called singular. The set of singular points will be denoted by $\Sigma$.
Theorem 3.3.
With the assumptions (3.1), suppose furthermore that the Brownian motion $\left(W(t)\right)_{t\geq 0}$ takes value in $D(A^{\frac{1}{4}+\delta})$ for some $\delta>0$. Then there exists a martingale suitable weak solution to (1.1) such that with full probability,
$\mathcal{P}^{1}(\Sigma)=0,$
where $\mathcal{P}^{k}$, $0\leq k\leq 4$, denotes the $k$-dimensional Hausdorff measure on $\mathbb{T}^{3}\times\mathbb{R}_{+}$ with respect to the parabolic distance $\delta_{p}((x,t),(y,s)):=\max\{|x-y|,\sqrt{|t-s|}\},\forall(x,t),(y,s)\in\mathbb{T}^{3}\times\mathbb{R}_{+}.$
4. Existence
In this section, we will establish the existence of martingale suitable weak solutions to (1.1) in the sense of Definition 3.3.
4.1. Formal derivation of energy inequalities
The following a priori energy estimate will play an essential role in subsequent analysis.
Lemma 4.1 (Local energy equality).
Fix $\omega\in\Omega$ with ${{\bf{z}}}(\omega)\in L^{\infty}(0,T;H)\cap L^{2}(0,T;V)$ and let the triple $({{\bf{v}}}(\omega),{\pi}(\omega),{{\bf{d}}}(\omega))$ be a classical solution to (3.3), then for any $\varphi\in C_{0}^{\infty}(Q_{T})$, $\varphi\geq 0$, it holds that
(4.1) $\frac{{\,\mathrm{d}}}{{\,\mathrm{d}}t}\int_{\mathbb{T}^{3}}\left(\frac{|{{\bf{v}}}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}|^{2}}{2}+F({{\bf{d}}})\right)\varphi+\int_{\mathbb{T}^{3}}\left(|\nabla{{\bf{v}}}|^{2}+|\Delta{{\bf{d}}}|^{2}+|f({{\bf{d}}})|^{2}\right)\varphi$
$=\int_{\mathbb{T}^{3}}\left(\frac{|{{\bf{v}}}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}|^{2}}{2}+F({{\bf{d}}})\right)\partial_{t}\varphi+\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla]{{{\bf{v}}}}\cdot{{{\bf{z}}}}\varphi$
$+\int_{\mathbb{T}^{3}}\left(\frac{|{{\bf{v}}}|^{2}}{2}+{{\bf{v}}}\cdot{{\bf{z}}}\right)({{\bf{z}}}+{{\bf{v}}})\cdot\nabla\varphi+{\pi}{{\bf{v}}}\cdot\nabla\varphi+\left(\frac{|{{\bf{v}}}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}|^{2}}{2}\right)\Delta\varphi$
$+\int_{\mathbb{T}^{3}}\left(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}|\nabla{{\bf{d}}}|^{2}{\rm I}_{3}\right):\nabla^{2}\varphi$
$+\int_{\mathbb{T}^{3}}({\bf{z}}\cdot\nabla){{\bf{d}}}\cdot(\Delta{{\bf{d}}}-f({{\bf{d}}}))\varphi+\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{d}}}]\cdot(\nabla\varphi\cdot\nabla){{\bf{d}}}$
$-\int_{\mathbb{T}^{3}}f({{\bf{d}}})\cdot(\nabla\varphi\cdot\nabla){{\bf{d}}}-2\int_{\mathbb{T}^{3}}\nabla f({{\bf{d}}}):\nabla{{\bf{d}}}\varphi$
Proof.
Taking the inner product of (3.3)${}_{1}$ with respect to ${{\bf{v}}}\varphi$ and integrating over $\mathbb{T}^{3}$ gives
$\int_{\mathbb{T}^{3}}\partial_{t}{{\bf{v}}}\cdot{{\bf{v}}}\varphi=\frac{\,\mathrm{d}}{{\,\mathrm{d}t}}\int_{\mathbb{T}^{3}}\frac{|{{\bf{v}}}|^{2}}{2}\varphi-\int_{\mathbb{T}^{3}}\frac{|{{\bf{v}}}|^{2}}{2}\partial_{t}\varphi.$
Using integration by parts we can derive
$\displaystyle\int_{\mathbb{T}^{3}}-\Delta{{\bf{v}}}\cdot{{\bf{v}}}\varphi$ $\displaystyle=\int_{\mathbb{T}^{3}}\nabla{{\bf{v}}}:\nabla({{\bf{v}}}\varphi)$
$\displaystyle=\int_{\mathbb{T}^{3}}|\nabla{{\bf{v}}}|^{2}\varphi+\int_{\mathbb{T}^{3}}\nabla{{\bf{v}}}:({{\bf{v}}}\otimes\nabla\varphi)$
$\displaystyle=\int_{\mathbb{T}^{3}}|\nabla{{\bf{v}}}|^{2}\varphi+\int_{\mathbb{T}^{3}}\nabla\left(\frac{|{{\bf{v}}}|^{2}}{2}\right)\cdot\nabla\varphi$
$\displaystyle=\int_{\mathbb{T}^{3}}|\nabla{{\bf{v}}}|^{2}\varphi-\int_{\mathbb{T}^{3}}\frac{|{{\bf{v}}}|^{2}}{2}\Delta\varphi.$
Now, we can see
$\displaystyle\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla({{\bf{z}}}+{{\bf{v}}})]\cdot{{\bf{v}}}\varphi$
$\displaystyle=\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{z}}}]\cdot{{\bf{v}}}\varphi+\int_{\mathbb{T}^{3}}({{\bf{z}}}+{{\bf{v}}})\cdot\nabla\left(\frac{|{{\bf{v}}}|^{2}}{2}\right)\varphi$
$\displaystyle=\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{z}}}]\cdot{{\bf{v}}}\varphi-\int_{\mathbb{T}^{3}}\frac{|{{\bf{v}}}|^{2}}{2}({{\bf{z}}}+{{\bf{v}}})\cdot\nabla\varphi$
$\displaystyle=-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{v}}}]\cdot{{\bf{z}}}\varphi-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla\varphi]({\bf{v}}\cdot{\bf{z}})-\int_{\mathbb{T}^{3}}\frac{|{{\bf{v}}}|^{2}}{2}({{\bf{z}}}+{{\bf{v}}})\cdot\nabla\varphi$
$\displaystyle=-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{v}}}]\cdot{{\bf{z}}}\varphi-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla\varphi]({\bf{v}}\cdot{\bf{z}})+\frac{|{{\bf{v}}}|^{2}}{2}({{\bf{z}}}+{{\bf{v}}})\cdot\nabla\varphi$
$\displaystyle=-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{v}}}]\cdot{{\bf{z}}}\varphi-\int_{\mathbb{T}^{3}}\left(\frac{|{{\bf{v}}}|^{2}}{2}+({\bf{v}}\cdot{\bf{z}})\right)[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla\varphi].$
Meanwhile, we have
$\int_{\mathbb{T}^{3}}\nabla{\pi}\cdot{{\bf{v}}}\varphi=-\int_{\mathbb{T}^{3}}{\pi}{{\bf{v}}}\cdot\nabla\varphi.$
For the Ericksen stress tensor term, using (3.4) we obtain
$\displaystyle\int_{\mathbb{T}^{3}}-\mathop{\rm div}\nolimits\left(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}|\nabla{{\bf{d}}}|^{2}{\rm I}_{3}-F({{\bf{d}}})\right)\cdot{{\bf{v}}}\varphi$ $\displaystyle=\int_{\mathbb{T}^{3}}[{{\bf{v}}}\cdot\nabla]{{\bf{d}}}\cdot(-\Delta{{\bf{d}}}+f({{\bf{d}}}))\varphi$
(4.2) $\displaystyle=\underbrace{-\int_{\mathbb{T}^{3}}[{{\bf{v}}}\cdot\nabla]{{\bf{d}}}\cdot\Delta{{\bf{d}}}\varphi}_{\mathcal{I}}\underbrace{+\int_{\mathbb{T}^{3}}[{{\bf{v}}}\cdot\nabla]{{\bf{d}}}\cdot f({{\bf{d}}})\varphi}_{\mathcal{II}}.$
Differentiating (${}_{3}$ gives
(4.3) $\partial_{t}\nabla{{\bf{d}}}+\nabla\left[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{d}}}\right]=\Delta(\nabla{{\bf{d}}})-\nabla(f({{\bf{d}}})).$
Taking the inner product of (4.3) with respect to $\nabla{{\bf{d}}}\varphi$ and integrating over $\mathbb{T}^{3}$ we obtain
$\int_{\mathbb{T}^{3}}\partial_{t}(\nabla{{\bf{d}}}):\nabla{{\bf{d}}}\varphi=\frac{\,\mathrm{d}}{{\,\mathrm{d}t}}\int_{\mathbb{T}^{3}}\frac{1}{2}|\nabla{{\bf{d}}}|^{2}\varphi-\int_{\mathbb{T}^{3}}\frac{1}{2}|\nabla{{\bf{d}}}|^{2}\partial_{t}\varphi.$
Meanwhile, we have that
$\displaystyle\int_{\mathbb{T}^{3}}\nabla[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{d}}}]:\nabla{{\bf{d}}}\varphi$ $\displaystyle=\int_{\mathbb{T}^{3}}-[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla]{{\bf{d}}}\cdot\Delta{{\bf{d}}}\varphi-[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla]{{\bf{d}}}\cdot[\nabla\varphi\cdot\nabla]{{\bf{d}}}$
$\displaystyle=-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla]{{\bf{d}}}\cdot\Delta{{\bf{d}}}\varphi-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla]{{\bf{d}}}\cdot[\nabla\varphi\cdot\nabla]{{\bf{d}}}$
(4.4) $\displaystyle=\underbrace{-\int_{\mathbb{T}^{3}}[{{\bf{z}}}\cdot\nabla]{{\bf{d}}}\cdot\Delta{{\bf{d}}}\varphi}_{\mathcal{III}}\underbrace{-\int_{\mathbb{T}^{3}}[{{\bf{v}}}\cdot\nabla]{{\bf{d}}}\cdot\Delta{{\bf{d}}}\varphi}_{\mathcal{IV}}-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla]{{\bf{d}}}\cdot[\nabla\varphi\cdot\nabla]{{\bf{d}}}.$
Then through integration by parts we obtain for the right-hand side
(4.5) $\int_{\mathbb{T}^{3}}\nabla(\Delta{{\bf{d}}}-f({{\bf{d}}})):\nabla{{\bf{d}}}\varphi=$ $-\int_{\mathbb{T}^{3}}|\Delta{{\bf{d}}}|^{2}\varphi-\int_{\mathbb{T}^{3}}\Delta{{\bf{d}}}\cdot(\nabla\varphi\cdot\nabla{{\bf{d}}})-\int_{\mathbb{T}^{3}}\nabla f({{\bf{d}}}):\nabla{{\bf{d}}}\varphi$
$=$ $-\int_{\mathbb{T}^{3}}|\Delta{{\bf{d}}}|^{2}\varphi$
$+\int_{\mathbb{T}^{3}}\frac{1}{2}|\nabla{{\bf{d}}}|^{2}\Delta\varphi+\int_{\mathbb{T}^{3}}\left(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}|\nabla{{\bf{d}}}|^{2}{\rm I}_{3}\right):\nabla^{2}\varphi$
$-\int_{\mathbb{T}^{3}}\nabla f({{\bf{d}}}):\nabla{{\bf{d}}}\varphi.$
Finally, taking the inner product of (${}_{3}$ with $f({\bf{d}})\varphi$ yields
$\displaystyle\int_{\mathbb{T}^{3}}\partial_{t}{{\bf{d}}}\cdot f({{\bf{d}}})\varphi$ $\displaystyle=\frac{\,\mathrm{d}}{{\,\mathrm{d}t}}\int_{\mathbb{T}^{3}}F({{\bf{d}}})\varphi-\int_{\mathbb{T}^{3}}F({{\bf{d}}})\partial_{t}\varphi,$
(4.6) $\displaystyle\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{d}}}]\cdot f({{\bf{d}}})\varphi$ $\displaystyle=\underbrace{\int_{\mathbb{T}^{3}}({\bf{z}}\cdot\nabla{\bf{d}})\cdot f({\bf{d}})\varphi}_{\mathcal{V}}\underbrace{+\int_{\mathbb{T}^{3}}({\bf{v}}\cdot\nabla{\bf{d}})\cdot f({\bf{d}})\varphi}_{\mathcal{VI}},$
(4.7) $\displaystyle\int_{\mathbb{T}^{3}}(\Delta{{\bf{d}}}-f({{\bf{d}}}))\cdot f({{\bf{d}}})\varphi$ $\displaystyle=-\int_{\mathbb{T}^{3}}\nabla{{\bf{d}}}:\nabla f({{\bf{d}}})\varphi+f({{\bf{d}}})\cdot(\nabla\varphi\cdot\nabla{{\bf{d}}})+|f({{\bf{d}}})|^{2}\varphi.$
Combining all the above identities, we obtain (4.1). Particularly, we utilize the following cancellations
$\displaystyle\mathcal{I}-\mathcal{IV}=0,$
$\displaystyle\mathcal{II}-\mathcal{VI}=0.$
Moreover, note that we have
$\mathcal{III}+\mathcal{V}=-\int_{\mathbb{T}^{3}}({\bf{z}}\cdot\nabla){{\bf{d}}}\cdot(\Delta{{\bf{d}}}-f({{\bf{d}}}))\varphi$
which then yields the seventh term on the right-hand-side of (4.1) ∎
We note that the above computations still hold if we choose $\varphi$ to be a constant function and the same calculation will result in the following global energy equality:
Corollary 4.2.
Under the seme assumptions as in Lemma 4.1, it holds that
(4.8) $\frac{{\,\mathrm{d}}}{{\,\mathrm{d}}t}\int_{\mathbb{T}^{3}}\left(\frac{|{{\bf{v}}}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}|^{2}}{2}+F({{\bf{d}}})\right)+\int_{\mathbb{T}^{3}}\left(|\nabla{{\bf{v}}}|^{2}+|\Delta{{\bf{d}}}-f({{\bf{d}}})|^{2}\right)$
$=\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla]{{{\bf{v}}}}\cdot{{{\bf{z}}}}+\int_{\mathbb{T}^{3}}({\bf{z}}\cdot\nabla){{\bf{d}}}\cdot(\Delta{{\bf{d}}}-f({{\bf{d}}})).$
4.2. Local-energy-preserving approximation
Given ${{\bf{z}}}(\omega)$ as in Lemma 3.2, we introduce the following approximated equations to (3.3):
(4.9) $\left\{\begin{array}[]{l}\partial_{t}{{\bf{v}}}^{\sigma}-\Delta{{\bf{v}}}^{\sigma}+({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\cdot\nabla({{\bf{z}}}+{{\bf{v}}}^{\sigma})+\nabla{\pi}^{\sigma}=-\nabla\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})),\\ \nabla\cdot{{\bf{v}}}^{\sigma}=0,\\ \partial_{t}{{\bf{d}}}^{\sigma}-({{\bf{z}}}+{{\bf{v}}}^{\sigma})\cdot\nabla\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]=\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma}),\end{array}\right.$
where $\sigma>0$ is the approximation parameter and $\Phi_{\sigma}[f]$ denotes the smooth mollification of a function $f$. More precisely, suppose that $\psi\in C_{0}^{\infty}(\mathbb{R}^{3})$ is a standard mollifier supported in the $B_{1}(0)$, and let $\psi_{\sigma}:=\sigma^{-3}\psi(x/\sigma)$. Then we define $\Phi_{\sigma}[f]$ via the following convolution
$\Phi_{\sigma}[f]=\int_{\mathbb{R}^{3}}\psi_{\sigma}(x-y)f(y)\,\mathrm{d}y,$
where we extend $f$ periodically to the whole $\mathbb{R}^{3}$. It is easy to show that
$\|\Phi_{\sigma}[f]\|_{L^{p}(\mathbb{T}^{3})}\leq\|f\|_{L^{p}(\mathbb{T}^{3})},\qquad\lim_{\sigma\to 0}\|\Phi_{\sigma}[f]-f\|_{L^{p}(\mathbb{T}^{3})}=0.$
Observe that (4.9) has better regularity in nonlinear terms, and one could obtain the existence of solutions via a similar argument using the Galerkin method and Banach fixed point theorem as in [42]. We remark that this kind of approximation by regularization was introduced by Leray [32] for the Navier–Stokes equations. In [7], Caffarelli, Kohn, and Nirenberg adopt a similar approximation through a “retarded mollification” to obtain the existence of suitable weak solutions for Navier–Stokes equations. It is also applicable to more complicated hydrodynamic equations (see for instance, [13, 15]). The key feature of this type of approximation is that it preserves global and local energy inequalities.
Multiplying (4.9)${}_{1}$ by ${{\bf{v}}}^{\sigma}$, and (4.9)${}_{3}$ by $-\Delta{{\bf{d}}}^{\sigma}+f({{\bf{d}}}^{\sigma})$, integrating over $\mathbb{T}^{3}$ and adding two equations together we obtain (see also Corollary 4.2)
(4.10) $\frac{\,\mathrm{d}}{{\,\mathrm{d}t}}\int_{\mathbb{T}^{3}}\frac{1}{2}(|{{\bf{v}}}^{\sigma}|^{2}+|\nabla{{\bf{d}}}^{\sigma}|^{2})+F({{\bf{d}}}^{\sigma}){\,\mathrm{d}}x+\int_{\mathbb{T}^{3}}(|\nabla{{\bf{v}}}^{\sigma}|^{2}+|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})|^{2}){\,\mathrm{d}}x$
$=\int_{\mathbb{T}^{3}}[({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\cdot\nabla{{\bf{v}}}^{\sigma}]\cdot{{\bf{z}}}+{{\bf{z}}}\cdot\nabla\Phi[{{\bf{d}}}^{\sigma}]\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})){\,\mathrm{d}x}.$
By the Hölder inequality and Young inequality we can estimate
$\displaystyle\left|\int_{\mathbb{T}^{3}}{{\bf{z}}}\cdot\nabla\Phi[{{\bf{d}}}^{\sigma}]\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})){\,\mathrm{d}x}\right|$
$\displaystyle\leq\frac{1}{8}\|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{2}\|\nabla{{\bf{d}}}^{\sigma}\|_{L^{4}(\mathbb{T}^{3})}^{2},$
$\displaystyle\leq\frac{1}{8}\|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{2}\|\nabla^{2}{{\bf{d}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{\frac{3}{2}}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{\frac{1}{2}}$
$\displaystyle\leq\frac{1}{8}\|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})\|_{L^{2}(\mathbb{T}^{3})}^{2}+\frac{1}{4}\|\nabla^{2}{{\bf{d}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2},$
where $C$ is a constant which is independent of $\sigma$, and
$\displaystyle\left|\int_{\mathbb{T}^{3}}[({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\cdot\nabla{{\bf{v}}}^{\sigma}]\cdot{{\bf{z}}}{\,\mathrm{d}x}\right|$
$\displaystyle\leq\int_{\mathbb{T}^{3}}|{{\bf{z}}}||\nabla{{\bf{v}}}^{\sigma}||{{\bf{z}}}|+|\Phi_{\sigma}[{{\bf{v}}}^{\sigma}]||\nabla{{\bf{v}}}^{\sigma}||{{\bf{z}}}|$
$\displaystyle\leq\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}+\|\Phi_{\sigma}[{{\bf{v}}}^{\sigma}]\|_{L^{4}(\mathbb{T}^{3})}\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}$
$\displaystyle\leq\frac{1}{4}\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\left(\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}+\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{2}\|{{\bf{v}}}^{\sigma}\|_{L^{4}(\mathbb{T}^{3})}^{2}\right)$
$\displaystyle\leq\frac{1}{4}\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\left(\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}+\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{2}\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{\frac{3}{2}}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{\frac{1}{2}}\right)$
$\displaystyle\leq\frac{1}{2}\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\left(\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}+\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{4})}^{4}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}\right).$
Meanwhile, by a direct calculation, we can show that
$\displaystyle\int_{\mathbb{T}^{3}}|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})|^{2}{\,\mathrm{d}x}$ $\displaystyle=\int_{\mathbb{T}^{3}}[|\Delta{{\bf{d}}}^{\sigma}|^{2}+|f({{\bf{d}}}^{\sigma})|^{2}-2\Delta{{\bf{d}}}^{\sigma}\cdot f({{\bf{d}}}^{\sigma})]{\,\mathrm{d}x}$
$\displaystyle=\int_{\mathbb{T}^{3}}[|\Delta{{\bf{d}}}^{\sigma}|^{2}+|f({{\bf{d}}}^{\sigma})|^{2}+2\nabla{{\bf{d}}}^{\sigma}:\nabla f({{\bf{d}}}^{\sigma})]{\,\mathrm{d}x}$
$\displaystyle=\int_{\mathbb{T}^{3}}[|\Delta{{\bf{d}}}^{\sigma}|^{2}+|f({{\bf{d}}}^{\sigma})|^{2}-2|\nabla{{\bf{d}}}^{\sigma}|^{2}+2|\nabla{{\bf{d}}}^{\sigma}|^{2}|{{\bf{d}}}^{\sigma}|^{2}+4|(\nabla{{\bf{d}}}^{\sigma})^{\top}{{\bf{d}}}^{\sigma}|^{2}]{\,\mathrm{d}x}.$
By the Gronwall inequality, we can get the uniform energy estimate as follows
(4.11) $\sup_{0<\sigma<1}\sup_{0<t<T}$ $\left(\|{{\bf{v}}}^{\sigma}(t)\|_{H}^{2}+\|{{\bf{d}}}^{\sigma}(t)\|_{H^{1}(\mathbb{T}^{3})}^{2}\right)+\int_{0}^{T}\|\nabla{{\bf{v}}}^{\sigma}(t)\|_{L^{2}(\mathbb{T}^{3})}^{2}+\|\nabla^{2}{{\bf{d}}}^{\sigma}(t)\|_{L^{2}(\mathbb{T}^{3})}^{2}{\,\mathrm{d}t}$
$\leq$ $C\left(\|{\bf{u}}_{0}\|_{H}^{2}+\|{\bf{d}}_{0}\|_{H^{1}(\mathbb{T}^{3})}^{2}+\int_{0}^{T}\|{{\bf{z}}}(t)\|_{L^{4}(\mathbb{T}^{4})}^{4}\,\mathrm{d}t\right)e^{C\int_{0}^{T}\|{\bf{z}}(t)\|_{L^{4}(\mathbb{T}^{3})}^{4}\,\mathrm{d}t}$
=:absent:=:= : $\Psi(T).$
Taking the divergence of (4.9)${}_{1}$ we obtain
(4.12) $-\Delta{\pi}^{\sigma}=\mathop{\rm div}\nolimits^{2}[({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\otimes({{\bf{z}}}+{{\bf{v}}}^{\sigma})]+\mathop{\rm div}\nolimits(\nabla\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma}))).$
By the standard elliptic theory and the Sobolev inequalities, we can estimate
$\displaystyle\|{\pi}^{\sigma}\|_{L^{\frac{5}{3}}(Q_{T})}$ $\displaystyle\leq C\|({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\otimes({{\bf{z}}}+{{\bf{v}}}^{\sigma})\|_{L^{\frac{5}{3}}(Q_{T})}+\|\nabla(\Phi_{\sigma}[{{\bf{d}}}^{\sigma}])\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma}))\|_{L_{t}^{\frac{5}{3}}L_{x}^{\frac{15}{14}}}$
$\displaystyle\leq C\left(\|{{\bf{z}}}\|_{L^{\frac{10}{3}}(Q_{T})}^{2}+\|{{\bf{v}}}^{\sigma}\|^{2}_{L^{\frac{10}{3}}(Q_{T})}+\|\nabla{{\bf{d}}}^{\sigma}\|^{2}_{L_{t}^{10}L_{x}^{\frac{30}{13}}}\|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})\|_{L^{2}(Q_{T})}\right).$
This together with (4.11) implies that
(4.13) $\|{\pi}^{\sigma}\|_{L^{\frac{5}{3}}(Q_{T})}\leq C\Psi(T).$
provided $\int_{\mathbb{T}^{3}}{\pi}^{\sigma}=0.$ Furthermore, we can show with (4.9)${}_{1}$ that
(4.14) $\|\partial_{t}{{\bf{v}}}^{\sigma}\|_{L^{2}(0,T;D(A^{-1}))+L^{2}(0,T;H^{-1}(\mathbb{T}^{3}))+L^{\frac{5}{4}}(Q_{T})}$
$\leq C(\|{{\bf{v}}}^{\sigma}\|_{L^{2}(Q_{T})}+\|{\bf{z}}\|_{L^{4}(Q_{T})}^{2}+\|{{\bf{v}}}^{\sigma}\|_{L^{4}(Q_{T})}^{2}+\|\nabla{\bf{d}}^{\sigma}\|_{L^{\frac{10}{3}}(Q_{T})}\|\Delta{\bf{d}}^{\sigma}-f({{\bf{d}}}^{\sigma})\|_{L^{2}(Q_{T})})$
$\leq C(\sqrt{\Psi(T)}+\Psi(T)).$
And from (4.9)${}_{3}$ we can obtain
(4.15) $\|\partial_{t}{\bf{d}}^{\sigma}\|_{L^{\frac{5}{3}}(Q_{T})}$
$\leq C\left(\|{{\bf{z}}}\|_{L^{\frac{10}{3}}(Q_{T})}\|\nabla{\bf{d}}^{\sigma}\|_{L^{\frac{10}{3}}(Q_{T})}+\|{{\bf{v}}}^{\sigma}\|_{L^{\frac{10}{3}}(Q_{T})}\|\nabla{\bf{d}}^{\sigma}\|_{L^{\frac{10}{3}}(Q_{T})}+T^{\frac{1}{10}}\|\Delta{\bf{d}}^{\sigma}-f({{\bf{d}}}^{\sigma})\|_{L^{2}(Q_{T})}\right)$
$\leq C(\Psi(T)+T^{\frac{1}{10}}\sqrt{\Psi(T)}).$
Let $\nu^{\sigma}$ be the law of the random variable $U_{\sigma}:=({{\bf{v}}}^{\sigma},{\pi}^{\sigma},{{\bf{d}}}^{\sigma},{{\bf{z}}},W)$ with values in $\mathcal{E}$:
$\mathcal{E}=L^{2}(0,T;H)\times L^{\frac{5}{3}}(Q_{T})\times L^{2}(0,T;H^{1}(\mathbb{T}^{3};\mathbb{R}^{3}))\times C([0,T];H)\times C_{0}([0,T];H).$
Let $\mu^{\sigma}$ be the projection of $\nu^{\sigma}$ in the variables $({{\bf{v}}}^{\sigma},{{\bf{d}}}^{\sigma})$, that is, the law of $({{\bf{v}}}^{\sigma},{{\bf{d}}}^{\sigma})$. The key step to establish the probabilistic convergence of this approximation scheme is to show that the family of measures $\mu^{\sigma}$ is tight. Namely, for each $\varepsilon>0$, there exists a compact set $K_{\varepsilon}$ in $\mathcal{E}$ such that
(4.16) $\mu^{\sigma}(K_{\varepsilon})\geq 1-\varepsilon,\text{ for }0<\sigma<1.$
We take
$K_{R}$ :=$:=\Big{\{}({{\bf{v}}},{\bf{d}})\Big{|}\|{{\bf{v}}}\|_{L^{\infty}(0,T;H)\cap L^{2}(0,T;V)}+\|\partial_{t}{{\bf{v}}}\|_{L^{2}(0,T;D(A^{-1}))+L^{2}(0,T;H^{-1}(\mathbb{T}^{3}))+L^{\frac{5}{4}}(Q_{T})}$
$+\|{{\bf{d}}}\|_{L^{\infty}(0,T;H^{1}(\mathbb{T}^{3}))\cap L^{2}(0,T;H^{2})\cap H^{1}(0,T;L^{2}(\mathbb{T}^{3}))}+\|\partial_{t}{\bf{d}}^{\sigma}\|_{L^{\frac{5}{3}}(Q_{T})}\leq R\Big{\}}.$
By the Aubin-Lions lemma, we conclude that $K_{R}$ is compact in $L^{p_{1}}(Q_{T})\times L^{p_{2}}(Q_{T})$ with $1<p_{1}<\frac{10}{3}$, and $10\displaystyle 1<p_{2}<$ 10. Moreover, by the estimates (3.5), (4.11), (4.14) and (4.15), we can follow the same argument as in [19, (381)-(383)] to conclude that there exists a sufficiently large $R>0$ such that $K_{1/R}$ satisfies (4.16). Consequently, the laws of $({{\bf{v}}}^{\sigma},{{\bf{d}}}^{\sigma},{{\bf{z}}},W)$, which will be denoted as $\xi^{\sigma}$ subsequentely, are tight over the phase space $\mathcal{G}$:
$\mathcal{G}=L^{p_{1}}(Q_{T})\times L^{p_{2}}(Q_{T})\times C([0,T];H)\times C_{0}([0,T];H),1<p_{1}<\frac{10}{3},1<p_{2}<10.$
Since the family of measures $\{\xi^{\sigma}\}$ is weakly compact on $\mathcal{G}$, we deduce that $\{\xi^{\sigma}\}$ converges weakly to a probability measure $\xi$ on $\mathcal{G}$ up to a subsequence. Hence, using a version of the Skorokhod embedding theorem (See [42, Lemma 3.3]), we can conclude that there exists a probability tripe $(\widetilde{\Omega},\widetilde{\mathcal{F}},\widetilde{\mathbb{P}})$, a sequence of random variables $\widetilde{U}_{k}=(\widetilde{{{\bf{v}}}}^{\sigma_{k}},\widetilde{\pi}^{\sigma_{k}},\widetilde{{{\bf{d}}}}^{\sigma_{k}},\widetilde{{{\bf{z}}}}^{k},\widetilde{W}^{k})$ with values in $\mathcal{E}$, and an element $\widetilde{U}=(\widetilde{{\bf{v}}},\widetilde{\pi},\widetilde{{\bf{d}}},\widetilde{{\bf{z}}},\widetilde{W})\in\mathcal{E}$ such that for each $k$, $\widetilde{U}_{k}$ has the same law with $U_{\sigma_{k}}$, in particular, $\widetilde{W}^{k}$ is a Brownian motion adapted to the filtration {~$\{\widetilde{\mathcal{F}}^{k}_{t}\}_{t\geq 0}$ which is given by the completion of the $\sigma$-algebra generated by $\{\widetilde{U}_{k}(s);s\leq t\}$, and
(4.17) $(\widetilde{{{\bf{v}}}}^{\sigma_{k}},\widetilde{{{\bf{d}}}}^{\sigma_{k}},\widetilde{{{\bf{z}}}}^{k},\widetilde{W}^{k})\rightarrow(\widetilde{{\bf{v}}},\widetilde{{\bf{d}}},\widetilde{{\bf{z}}},\widetilde{W})\quad\widetilde{\mathbb{P}}-a.s.,\mbox{ as }k\to\infty,$
in the topology of $\mathcal{G}$. Additionally, setting {~$\{\widetilde{\mathcal{F}}_{t}\}_{t\geq 0}$ as the completion of the $\sigma$-algebra generated by $\{\widetilde{U}(s);s\leq t\}$, we obtain a new stochastic basis $\widetilde{\mathfrak{B}}=\left(\widetilde{\Omega},\widetilde{\mathcal{F}},\{\widetilde{\mathcal{F}}_{t}\}_{t\geq 0},\widetilde{\mathbb{P}},(\widetilde{W}(t)\right)_{t\geq 0})$, as in (3.7).
Next we show that for each $k$, the random variable $(\widetilde{{\bf{v}}}^{\sigma_{k}},\widetilde{{\pi}}^{\sigma_{k}},\widetilde{{\bf{d}}}^{\sigma_{k}},\widetilde{{{\bf{z}}}}^{k})$ satisfies the local energy equality (4.1). We adapt a trick of Bensoussan [2]. Given $\varphi\in C_{c}^{\infty}(Q_{T};\mathbb{R}_{+})$, for $({{\bf{v}}}^{\sigma},{\pi}^{\sigma},{{\bf{d}}}^{\sigma},{{{\bf{z}}}})$, the components of ${U}_{\sigma}$, define the random variable $X^{\sigma}$ on $(\Omega,\mathcal{F})$ as
$\displaystyle X^{\sigma}$ $\displaystyle=\Bigg{|}\int_{\mathbb{T}^{3}\times\{T\}}\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}^{\sigma}|^{2}}{2}+F({{\bf{d}}}^{\sigma})\right)\varphi+\int_{Q_{T}}(|\nabla{{\bf{v}}}^{\sigma}|^{2}+|\Delta{{\bf{d}}}^{\sigma}|^{2}+|f({{\bf{d}}}^{\sigma})|^{2})\varphi$
$\displaystyle-\int_{Q_{T}}\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}^{\sigma}|^{2}}{2}+F({{\bf{d}}}^{\sigma})\right)\partial_{t}\varphi-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\cdot\nabla{{\bf{v}}}^{\sigma}]\cdot{{\bf{z}}}\varphi$
$\displaystyle-\int_{Q_{T}}\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}+{{\bf{v}}}^{\sigma}\cdot{\bf{z}}\right)({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\cdot\nabla\varphi-{\pi}^{\sigma}{{\bf{v}}}^{\sigma}\cdot\nabla\varphi-\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}^{\sigma}|^{2}}{2}\right)\Delta\varphi$
$\displaystyle-\int_{Q_{T}}\left(\nabla{{\bf{d}}}^{\sigma}\odot\nabla{{\bf{d}}}^{\sigma}-\frac{1}{2}|\nabla{{\bf{d}}}^{\sigma}|^{2}{\rm I}_{3}\right):\nabla^{2}\varphi$
$\displaystyle-\int_{Q_{T}}({{\bf{z}}}\cdot\nabla\Phi_{\sigma}[{{\bf{d}}}^{\sigma}])\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma}))-\int_{Q_{T}}[({{\bf{z}}}+{{\bf{v}}}^{\sigma})\cdot\nabla{{\bf{d}}}^{\sigma}]\cdot(\nabla\varphi\cdot\nabla){{\bf{d}}}^{\sigma}$
$\displaystyle+\int_{Q_{T}}f({{\bf{d}}}^{\sigma})(\nabla\varphi\cdot\nabla{{\bf{d}}}^{\sigma})+2\nabla f({{\bf{d}}}^{\sigma}):\nabla{{\bf{d}}}^{\sigma}\varphi\Bigg{|}.$
Following a similar proof as for Lemma 4.1, we can conclude that for $1\displaystyle 0<\sigma<$ 1, we have
$\displaystyle\frac{\,\mathrm{d}}{{\,\mathrm{d}t}}\int_{\mathbb{T}^{3}}\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}^{\sigma}|^{2}}{2}+F({{\bf{d}}}^{\sigma})\right)\varphi+\int_{\mathbb{T}^{3}}(|\nabla{{\bf{v}}}^{\sigma}|^{2}+|\Delta{{\bf{d}}}^{\sigma}|^{2}+|f({{\bf{d}}}^{\sigma})|^{2})\varphi$
$\displaystyle\leq\int_{\mathbb{T}^{3}}\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}^{\sigma}|^{2}}{2}+F({{\bf{d}}}^{\sigma})\right)\partial_{t}\varphi+\int_{\mathbb{T}^{3}}[({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\cdot\nabla{{\bf{v}}}^{\sigma}]\cdot{{\bf{z}}}\varphi$
$\displaystyle+\int_{\mathbb{T}^{3}}\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}+{{\bf{v}}}^{\sigma}\cdot{\bf{z}}\right)({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\cdot\nabla\varphi+{\pi}^{\sigma}{{\bf{v}}}^{\sigma}\cdot\nabla\varphi+\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}^{\sigma}|^{2}}{2}\right)\Delta\varphi$
$\displaystyle+\int_{\mathbb{T}^{3}}\left(\nabla{{\bf{d}}}^{\sigma}\odot\nabla{{\bf{d}}}^{\sigma}-\frac{1}{2}|\nabla{{\bf{d}}}^{\sigma}|^{2}{\rm I}_{3}\right):\nabla^{2}\varphi$
$\displaystyle+\int_{\mathbb{T}^{3}}({{\bf{z}}}\cdot{\nabla\Phi_{\sigma}[{{\bf{d}}}^{\sigma}])})\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma}))+[({{\bf{z}}}+{{\bf{v}}}^{\sigma})\cdot\nabla{{\bf{d}}}^{\sigma}]\cdot(\nabla\varphi\cdot\nabla){{\bf{d}}}^{\sigma}$
$\displaystyle-\int_{\mathbb{T}^{3}}f({{\bf{d}}}^{\sigma})\cdot(\nabla\varphi\cdot\nabla{{\bf{d}}}^{\sigma})-2\int_{\mathbb{T}^{3}}\nabla f({{\bf{d}}}^{\sigma}):\nabla{{\bf{d}}}^{\sigma}\varphi.$
Hence $X^{\sigma}=0$ with probability 1. Thus we have $\mathbb{E}(g(U_{\sigma}))=0$, where $g(U_{\sigma}):=X^{\sigma}/(1+X^{\sigma}).$ Let $\widetilde{X}^{k}$ be defined analogously for $(\widetilde{{{\bf{v}}}}^{\sigma_{k}},\widetilde{{\pi}}^{\sigma_{k}},\widetilde{{{\bf{d}}}}^{\sigma_{k}},\widetilde{{{\bf{z}}}}^{k})$, the components of $\widetilde{U}_{k}$, as a random variable on $(\widetilde{\Omega},\widetilde{\mathcal{F}})$. Note that $g(\cdot)$ is a deterministic bounded continuous function on a subset of $\mathcal{E}$, and thus
$\widetilde{\mathbb{E}}\left[\frac{\widetilde{X}^{k}}{1+\widetilde{X}^{k}}\right]=\widetilde{\mathbb{E}}[g(\widetilde{U}_{k})]=\mathbb{E}[g(U_{\sigma_{k}})]=\mathbb{E}\left[\frac{X^{\sigma_{k}}}{1+X^{\sigma_{k}}}\right]=0,$
which implies that
[…]
as desired. With a similar method, we can verify one more thing. That is, for each $\sigma_{k}$, $\widetilde{U}_{k}$ satisfies the equations (3.2) and (4.9) in the sense of distribution $\widetilde{\mathbb{P}}$-a.s.. Particularly, the random variables $(\widetilde{{{\bf{v}}}}^{\sigma_{k}},\widetilde{{\pi}}^{\sigma_{k}},\widetilde{{{\bf{d}}}}^{\sigma_{k}},\widetilde{{{\bf{z}}}}^{k})$ satisfy the approximation equation (4.9) $\widetilde{\mathbb{P}}$-a.s., that is
(4.19) $\left\{\begin{array}[]{l}\partial_{t}\widetilde{{\bf{v}}}^{\sigma_{k}}-\Delta\widetilde{{\bf{v}}}^{\sigma_{k}}+(\widetilde{{\bf{z}}}^{k}+\Phi_{\sigma_{k}}[\widetilde{{\bf{v}}}^{\sigma_{k}}])\cdot\nabla(\widetilde{{\bf{z}}}^{k}+\widetilde{{\bf{v}}}^{\sigma_{k}})+\nabla\widetilde{\pi}^{\sigma_{k}}=-\nabla\Phi_{\sigma_{k}}[\widetilde{{\bf{d}}}^{\sigma_{k}}]\cdot(\Delta\widetilde{{\bf{d}}}^{\sigma_{k}}-f(\widetilde{{\bf{d}}}^{\sigma_{k}})),\\ \nabla\cdot\widetilde{{\bf{v}}}^{\sigma_{k}}=0,\\ \partial_{t}\widetilde{{\bf{d}}}^{\sigma_{k}}-(\widetilde{{\bf{z}}}^{k}+\widetilde{{\bf{v}}}^{\sigma_{k}})\cdot\nabla\Phi_{\sigma_{k}}[\widetilde{{\bf{d}}}^{\sigma_{k}}]=\Delta\widetilde{{\bf{d}}}^{\sigma_{k}}-f(\widetilde{{\bf{d}}}^{\sigma_{k}}).\end{array}\right.$
As a result, passing to the limit on (the weak form of) the equations (3.2) and (4.19) as $\sigma_{k}\to 0$, we conclude that $(\widetilde{{\bf{v}}},\widetilde{{\bf{d}}},\widetilde{{\bf{z}}},\widetilde{W})$ solves (3.2) and (3.3) in the sense of distribution (see Def. 3.1) $\widetilde{\mathbb{P}}$-a.s.. Thus there exists a distribution $\widetilde{\pi}$ defined on the stochastic basis $\widetilde{\mathfrak{B}}$ which plays the role of the pressure field for $\widetilde{{\bf{v}}}$. In fact, let $\widetilde{{\pi}}$ solve the following Poisson equation
$-\Delta\widetilde{{\pi}}=\mathop{\rm div}\nolimits^{2}[(\widetilde{{{\bf{z}}}}+\widetilde{{{\bf{v}}}})\otimes(\widetilde{{{\bf{z}}}}+\widetilde{{{\bf{v}}}})]+\mathop{\rm div}\nolimits(\nabla\widetilde{{\bf{d}}}\cdot(\Delta\widetilde{{\bf{d}}}-f(\widetilde{{\bf{d}}}))),$
with $\int_{{\mathbb{T}}^{3}}\widetilde{\pi}(t)=0$ for every $t$. Thanks to (4.19), for almost sure $\widetilde{\omega}\in\widetilde{\Omega}$, we have the same estimate as in (4.13) for $\widetilde{{\pi}}^{\sigma_{k}}(\widetilde{\omega})$, which implies that the sequence $\widetilde{{\pi}}^{\sigma_{k}}(\widetilde{\omega})$ is bounded uniformly in $L^{5/3}(Q_{T})$ independent of $\sigma$. Hence we have for almost every fixed $\widetilde{\omega}\in\widetilde{\Omega}$, there exists a subsequence that may depend on $\widetilde{\omega}$ such that
(4.20) $\widetilde{{\pi}}^{\sigma_{k}}(\widetilde{\omega})\to\widetilde{\pi}(\widetilde{\omega})\mbox{ weakly in }L^{\frac{5}{3}}(Q_{T})\mbox{ up to the subsequence}.$
Finally, we show that $(\widetilde{{\bf{v}}},\widetilde{\pi},\widetilde{{\bf{d}}})$ satisfies the local energy inequality described in Definition 3.2. For almost sure $\widetilde{\omega}\in\widetilde{\Omega}$, as proven above (see equation (4.18)) we know that the random variables $(\widetilde{{\bf{v}}}^{\sigma_{k}}(\widetilde{\omega}),\widetilde{{\pi}}^{\sigma_{k}}(\widetilde{\omega}),\widetilde{{\bf{d}}}^{\sigma_{k}}(\widetilde{\omega}),\widetilde{{{\bf{z}}}}^{k}(\widetilde{\omega}))$ satisfy the local energy equality (4.1). Thus for each such $\widetilde{\omega}$, passing to the limit upon the corresponding subsequence as in (4.20)(using also (4.17)), we obtain the local energy inequality for $(\widetilde{{\bf{v}}}(\widetilde{\omega}),\widetilde{\pi}(\widetilde{\omega}),\widetilde{{\bf{d}}}(\widetilde{\omega}),\widetilde{{{\bf{z}}}}(\widetilde{\omega}))$ by the lower semicontinuity.
5. Partial regularity
5.1. Blowing up analysis
First, we recall the maximum principle for the $L^{\infty}$ of ${\bf{d}}$ from the transported Ginzburg–Landau heat flow:
(5.1) $\left\{\begin{array}[]{l}\partial_{t}{{\bf{d}}}+{{\bf{u}}}\cdot\nabla{{\bf{d}}}=\Delta{{\bf{d}}}-f({{\bf{d}}}),\\ \nabla\cdot{\bf{u}}=0,\\ {{\bf{d}}}(0)={{\bf{d}}}_{0}.\end{array}\right.$
Lemma 5.1 ([39]).
For $T>0$, assume that ${{\bf{u}}}\in L^{2}(0,T;V)$ and ${{\bf{d}}}_{0}\in H^{1}(\mathbb{T};\mathbb{R}^{3})$ with $|{{\bf{d}}}_{0}(x)|\leq 1$ a.e. $x\in\mathbb{T}^{3}$. Suppose ${{\bf{d}}}\in L^{2}(0,T;H^{2}(\mathbb{T}^{3},\mathbb{R}^{3}))$ is a weak solution to (5.1). Then $|{{\bf{d}}}(x,t)|\leq 1$ for a.e. $(x,t)\in Q_{T}$.
The following lemma will play an essential role in the partial regularity of suitable weak solutions.
Lemma 5.2.
There exist $\varepsilon_{0}>0$, $0<\tau_{0}<\frac{1}{2}$ and $C_{0}>0$ such that if $({{\bf{v}}},{{\bf{d}}},{\pi})$ is a suitable weak solution in $\mathbb{T}^{3}\times(0,\infty)$, then for $z_{0}=(x_{0},t_{0})\in\mathbb{T}^{3}\times(r^{2},\infty)$, and $r>0$, if it holds that
$|{{\bf{z}}}|\leq M,\quad|{{\bf{d}}}|\leq M\text{ in }\mathcal{Q}_{r}(z_{0}),$
where $\mathcal{Q}_{r}(z_{0})=B_{r}(x_{0})\times(t_{0}-r^{2},t)$, and
$\Theta_{z_{0}}(r):=r^{-2}\int_{\mathcal{Q}_{r}(z_{0})}(|{{\bf{v}}}|^{3}+|\nabla{{\bf{d}}}|^{3})+\left(r^{-2}\int_{\mathcal{Q}_{r}(z_{0})}|\pi|^{\frac{3}{2}}\right)^{2}\leq\varepsilon_{0}^{3},$
then
$\Theta_{z_{0}}(\tau_{0}r)=(\tau_{0}r)^{-2}\int_{\mathcal{Q}_{\tau_{0}r}(z_{0})}(|{{\bf{v}}}|^{3}+|\nabla{{\bf{d}}}|^{3})+\left((\tau_{0}r)^{-2}\int_{\mathcal{Q}_{\tau_{0}r}(z_{0})}|\pi|^{\frac{3}{2}}\right)^{2}\leq\frac{1}{2}\max\{\Theta_{z_{0}}(r),C_{0}r^{3}\}.$
Proof.
Proof by contradiction. Suppose that it is false. Then there exists an $M_{0}>0$ such that for any $\tau\in(0,\frac{1}{2})$, there exist sequences $\varepsilon_{i}\to 0$, $C_{i}\to\infty$, $r_{i}>0$, and $z_{i}=(x_{i},t_{i})\in\mathbb{T}^{3}\times(r_{i}^{2},\infty)$ such that
(5.2) $\left\{\begin{array}[]{l}|{{\bf{z}}}|\leq M_{0},,\quad|{{\bf{d}}}|\leq M_{0}\text{ in }\mathcal{Q}_{r_{i}}(z_{i}),\\ \Theta_{z_{i}}(r_{i})=\varepsilon_{i}^{3},\\ \Theta_{z_{i}}(\tau r_{i})>\frac{1}{2}\max\{\varepsilon_{i}^{3},C_{i}r_{i}^{3}\}.\end{array}\right.$
First we notice that $C_{i}r_{i}^{3}\leq 2\Theta_{z_{i}}(\tau r_{i})\leq 2\tau^{-4}\Theta_{z_{i}}(r_{i})=2\tau^{-4}\varepsilon_{i}^{3}$. This implies that
$r_{i}\leq\left(\frac{2\varepsilon_{i}^{3}}{C_{i}\tau^{4}}\right)^{\frac{1}{3}}\to 0.$
Next, we define the rescaling sequence
$({{{\bf{z}}}}^{i},{{{\bf{v}}}}^{i},{{{\bf{d}}}}^{i},\pi^{i})(x,t)=(r_{i}{{\bf{z}}},r_{i}{{\bf{v}}},{{\bf{d}}},r_{i}^{2}{\pi})(x_{i}+r_{i}x,t_{i}+r_{i}^{2}t).$
It is straightforward to verify that $({{\bf{z}}}^{i},{{\bf{v}}}^{i},{{\bf{d}}}^{i},{\pi}^{i})$ satisfies the re-scaled equation
(5.3) $\left\{\begin{array}[]{l}\partial_{t}{{\bf{v}}}^{i}+({{\bf{z}}}^{i}+{{\bf{v}}}^{i})\cdot\nabla({{\bf{z}}}^{i}+{{\bf{v}}}^{i})+\nabla\pi^{i}\\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad=\Delta{\bf{v}}^{i}-\mathop{\rm div}\nolimits\left(\nabla{{\bf{d}}}_{i}\odot\nabla{{\bf{d}}}^{i}-\frac{1}{2}|\nabla{{\bf{d}}}^{i}|^{2}{\rm I}_{3}-r_{i}^{2}F({{\bf{d}}}^{i}){\rm I}_{3}\right),\\ \nabla\cdot{{\bf{v}}}^{i}=0,\\ \partial_{t}{{\bf{d}}}^{i}+({{\bf{z}}}^{i}+{{\bf{v}}}^{i})\cdot\nabla{{\bf{d}}}^{i}=\Delta{{\bf{d}}}^{i}-r^{2}_{i}f({{\bf{d}}}^{i}).\end{array}\right.$
And (5.2) becomes
(5.4) $\left\{\begin{array}[]{l}\displaystyle\int_{\mathcal{Q}_{1}(0)}(|{{\bf{v}}}^{i}|^{3}+|\nabla{{\bf{d}}}^{i}|^{3})+\left(\int_{\mathcal{Q}_{1}(0)}|\pi^{i}|^{\frac{3}{2}}\right)^{2}=\varepsilon_{i}^{3},\\ \displaystyle\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}(|{{\bf{v}}}^{i}|^{3}+|\nabla{{\bf{d}}}^{i}|^{3})+\left(\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}|\pi^{i}|^{\frac{3}{2}}\right)^{3}>\frac{1}{2}\max\{\varepsilon_{i}^{3},C_{i}r_{i}^{3}\}.\end{array}\right.$
Then we define the blowing-up sequence
$(\widehat{{\bf{z}}}^{i},\widehat{{\bf{v}}}^{i},\widehat{{\bf{d}}}^{i},\widehat{\pi}^{i})=\left(\frac{{{\bf{z}}}^{i}}{\varepsilon_{i}},\frac{{{\bf{v}}}^{i}}{\varepsilon_{i}},\frac{{{\bf{d}}}^{i}-\overline{{{\bf{d}}}^{i}}}{\varepsilon_{i}},\frac{{\pi}^{i}}{\varepsilon_{i}}\right),$
where $\overline{{{\bf{d}}}^{i}}=\fint_{\mathcal{Q}_{1}(0)}{{\bf{d}}}^{i}$, which denotes the average of ${{\bf{d}}}^{i}$ over $\mathcal{Q}_{1}(0)$. Then it turns out that $(\widehat{{\bf{z}}}^{i},\widehat{{\bf{v}}}^{i},\widehat{{\bf{d}}}^{i},\widehat{\pi}^{i})$ solves the following equation:
(5.5) $\left\{\begin{array}[]{l}\partial_{t}\widehat{{\bf{v}}}^{i}+\varepsilon_{i}(\widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i})\cdot\nabla(\widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i})+\nabla\widehat{\pi}^{i}=\Delta\widehat{{\bf{v}}}^{i}\\ \quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\mathop{\rm div}\nolimits\left(\varepsilon_{i}\nabla\widehat{{\bf{d}}}^{i}\odot\nabla\widehat{{\bf{d}}}^{i}-\frac{\varepsilon_{i}}{2}|\nabla\widehat{{\bf{d}}}^{i}|^{2}{\rm I}_{3}-\frac{r_{i}^{2}}{\varepsilon_{i}}F({{\bf{d}}}^{i}){\rm I}_{3}\right),\\ \mathop{\rm div}\nolimits\widehat{{\bf{v}}}^{i}=0,\\ \partial_{t}\widehat{{\bf{d}}}^{i}+\varepsilon_{i}(\widehat{{\bf{z}}}^{i}+\widehat{\bf{v}}^{i})\cdot\nabla\widehat{{\bf{d}}}^{i}=\Delta\widehat{{\bf{d}}}^{i}-\frac{r_{i}^{2}}{\varepsilon_{i}}f({\bf{d}}^{i}).\end{array}\right.$
Furthermore, we have
(5.6) $\left\{\begin{array}[]{l}\int_{\mathcal{Q}_{1}(0)}\widehat{{\bf{d}}}^{i}=0,\\ \int_{\mathcal{Q}_{1}(0)}\left(|\widehat{{\bf{v}}}^{i}|^{3}+|\nabla\widehat{{\bf{d}}}^{i}|^{3}\right)+\left(\int_{\mathcal{Q}_{1}(0)}|\widehat{\pi}^{i}|^{\frac{3}{2}}\right)^{2}=1,\\ \tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}\left(|\widehat{{\bf{v}}}^{i}|^{3}+|\nabla\widehat{{\bf{d}}}^{i}|^{3}\right)+\left(\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}|\widehat{\pi}^{i}|^{\frac{3}{2}}\right)^{2}>\frac{1}{2}\max\{1,C_{i}\left(\frac{r_{i}}{\varepsilon_{i}}\right)^{3}\}.\end{array}\right.$
From (5.6), by weak compactness, we can show there exists
$(\widehat{{\bf{v}}},\widehat{{\bf{d}}},\widehat{\pi})\in L^{3}(\mathcal{Q}_{1}(0))\times L_{t}^{3}W_{x}^{1,3}(\mathcal{Q}_{1}(0))\times L^{\frac{3}{2}}(\mathcal{Q}_{1}(0)),$
such that, after passing to a subsequence,
$(\widehat{{\bf{v}}}^{i},\widehat{{\bf{d}}}^{i},\widehat{\pi}^{i})\rightharpoonup(\widehat{{\bf{v}}},\widehat{{\bf{d}}},\widehat{\pi})\text{ in }L^{3}(\mathcal{Q}_{1}(0))\times L_{t}^{3}W_{x}^{1,3}(\mathcal{Q}_{1}(0))\times L^{\frac{3}{2}}(\mathcal{Q}_{1}(0)).$
It follows from (5.6) and the lower semicontinuity that
$\int_{\mathcal{Q}_{1}(0)}\left(|\widehat{{\bf{v}}}|^{3}+|\nabla\widehat{\bf{d}}|^{3}\right)+\left(\int_{\mathcal{Q}_{1}(0)}|\widehat{\pi}|^{\frac{3}{2}}\right)\leq 1.$
We claim that
$\sup_{i}\left(\|\widehat{{\bf{v}}}^{i}\|_{L_{t}^{\infty}L_{x}^{2}\cap L_{t}^{2}H_{x}^{1}(\mathcal{Q}_{1}(0))}+\|\widehat{{\bf{d}}}^{i}\|_{L_{t}^{\infty}H_{x}^{1}\cap L_{t}^{2}H_{x}^{2}(\mathcal{Q}_{1}(0))}\right)<\infty.$
We choose a cut-off function $\varphi\in C_{0}^{\infty}(\mathcal{Q}_{1}(0))$ such that
$0\leq\varphi\leq 1,\quad\varphi\equiv 1\text{ on }\mathcal{Q}_{\frac{1}{2}}(0),\quad\text{ and }|\partial_{t}\varphi|+|\nabla\varphi|+|\nabla^{2}\varphi|\leq C.$
Define $\varphi_{i}(x,t)=\varphi\left(\frac{x-x_{i}}{r_{i}},\frac{t-t_{i}}{r_{i}^{2}}\right)$, $\forall(x,t)\in\mathbb{T}^{3}\times(0,\infty).$ Replacing $\varphi$ by $\varphi_{i}^{2}$ in the local energy inequality (4.1) we can show that
$\displaystyle\sup_{t_{i}-\frac{r_{i}^{2}}{4}\leq t\leq t_{i}}\int_{B_{r_{i}}(x_{i})}\left(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}+F({{\bf{d}}})\right)\varphi_{i}^{2}{\,\mathrm{d}x}+\int_{\mathcal{Q}_{r_{i}/2}(z_{i})}\left(|\nabla{{\bf{v}}}|^{2}+|\Delta{{\bf{d}}}|^{2}+|f({{\bf{d}}})|^{2}\right)\varphi_{i}^{2}$
$\displaystyle\leq C\int_{\mathcal{Q}_{r_{i}}(z_{i})}\left(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}+F({{\bf{d}}})\right)|\partial_{t}\left(\varphi_{i}^{2}\right)|$
$\displaystyle+C\int_{\mathcal{Q}_{r_{i}}(z_{i})}|({{\bf{z}}}+{{\bf{v}}})||{{\bf{z}}}||\nabla{{\bf{v}}}||\nabla\left(\varphi_{i}^{2}\right)|+|{\pi}||{{\bf{v}}}||\nabla\left(\varphi_{i}^{2}\right)|+\left(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}\right)\Delta\left(\varphi_{i}^{2}\right)$
$\displaystyle+C\int_{\mathcal{Q}_{r_{i}}(z_{i})}|\nabla{{\bf{d}}}|^{2}|\nabla^{2}\left(\varphi_{i}^{2}\right)|+|{{\bf{z}}}||\nabla{{\bf{d}}}||\Delta{{\bf{d}}}-f({{\bf{d}}})|^{2}|\varphi_{i}^{2}|+|{{\bf{z}}}+{{\bf{v}}}||\nabla{{\bf{d}}}|^{2}|\nabla\left(\varphi_{i}^{2}\right)|$
$\displaystyle+C\int_{\mathcal{Q}_{r_{i}}(z_{i})}|f({{\bf{d}}})||\nabla{{\bf{d}}}||\nabla\left(\varphi_{i}^{2}\right)|+|\nabla f({{\bf{d}}})||\nabla{{\bf{d}}}|\varphi_{i}^{2}.$
By rescaling and Young’s inequality, we can show that
$\displaystyle\sup_{-\frac{1}{4}\leq t\leq 0}\int_{B_{\frac{1}{2}(0)}}(|\widehat{{\bf{v}}}^{i}|^{2}+|\nabla\widehat{{\bf{d}}}_{i}^{2}|^{2})+\int_{\mathcal{Q}_{\frac{1}{2}}(0)}(|\nabla\widehat{{\bf{v}}}^{i}|+|\nabla^{2}\widehat{{\bf{d}}}^{i}|^{2})\leq C,\qquad\forall i.$
Hence the claim holds. From Aubin–Lions’s lemma, we infer that
$(\widehat{{\bf{v}}}^{i},\widehat{{\bf{d}}}^{i})\to(\widehat{{\bf{v}}},\widehat{{\bf{d}}})\text{ strongly in }L^{3}(\mathcal{Q}_{\frac{1}{2}}(0))\times L_{t}^{3}W_{x}^{1,3}(\mathcal{Q}_{\frac{1}{2}}(0)).$
We observe that $(\widehat{{\bf{v}}},\widehat{{\bf{d}}},\widehat{\pi})$ solves the liner system
(5.7) $\left\{\begin{array}[]{l}\partial_{t}\widehat{{\bf{v}}}+\nabla\widehat{\pi}=\Delta\widehat{{\bf{v}}},\\ \nabla\cdot\widehat{{\bf{v}}}=0,\\ \partial_{t}\widehat{{\bf{d}}}=\Delta\widehat{{\bf{d}}}.\end{array}\right.$
By the standard interior regularity estimate we have that $(\widehat{\textbf{u}},\widehat{\textbf{d}})\in C^{\infty}(\mathcal{Q}_{\frac{1}{4}}(0))$, $\widehat{\pi}\in L^{\infty}([-\frac{1}{16},0],C^{\infty}(B_{\frac{1}{4}}(0)))$ with the following estimate:
(5.8) $\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}(|\widehat{\textbf{u}}|^{3}+|\nabla\widehat{\textbf{d}}|^{3})+\left(\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}|\widehat{\pi}|^{\frac{3}{2}}\right)^{2}$
$\leq C\tau^{3}\left[\int_{\mathcal{Q}_{\frac{1}{2}}(0)}(|\widehat{\textbf{u}}|^{3}+|\nabla\widehat{\textbf{d}}|^{3})+\left(\int_{\mathcal{Q}_{\frac{1}{2}}(0)}|\widehat{\pi}|^{\frac{3}{2}}\right)^{2}\right]$
$\leq C\tau^{3},\forall\tau\in\left(0,\frac{1}{8}\right).$
From the strong convergence of $(\widehat{{\bf{v}}}^{i},\widehat{{\bf{d}}}^{i})$ we infer that
$\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}(|\widehat{\textbf{u}}^{i}|^{3}+|\nabla\widehat{{\bf{d}}}^{i}|^{3})=\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}(|\widehat{{\bf{v}}}|^{3}+|\nabla\widehat{{\bf{d}}}|^{3})+\tau^{-2}o(i),$
where $o(i)\to 0$ as $i\to\infty$. It remains to estimate the pressure term $\widehat{\pi}^{i}$. Taking the divergence in (5.5)${}_{1}$, we see that
(5.9) $-\Delta\widehat{\pi}^{i}=\varepsilon_{i}\mathop{\rm div}\nolimits^{2}\left[\widehat{\textbf{u}}^{i}\otimes\widehat{\textbf{u}}^{i}+\nabla\widehat{{\bf{d}}}^{i}\odot\nabla\widehat{{\bf{d}}}^{i}-\frac{1}{2}|\nabla\widehat{{\bf{d}}}^{i}|^{2}{\rm I}_{3}-\frac{r_{i}^{2}}{\varepsilon_{i}^{2}}F({{\bf{d}}}^{i}){\rm I}_{3}\right].$
We need to show that
(5.10) $\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}|\widehat{\pi}^{i}|^{\frac{3}{2}}\leq C\tau^{-2}(\varepsilon_{i}+o(i))+C\tau.$
This will lead to a contradiction of (5.6)${}_{2}$ if we chose sufficient small $\tau_{0}\in(0,\frac{1}{4})$ and sufficient large $i_{0}$ such that for $i\geq i_{0}$, it holds that
(5.11) $\tau_{0}^{-2}\int_{\mathcal{Q}_{\tau_{0}}(0)}\left(|\widehat{{\bf{v}}}^{i}|^{3}+|\nabla\widehat{{\bf{d}}}^{i}|^{3}\right)+\left(\tau_{0}^{-2}\int_{\mathcal{Q}_{\tau_{0}}(0)}|\widehat{\pi}_{i}|^{\frac{3}{2}}\right)^{2}\leq\frac{1}{4}.$
To prove (5.10), let $\eta\in C_{0}^{\infty}(B_{1}(0))$ be a cut-off function of $B_{\frac{3}{8}}(0)$. For any $t$, $-(\frac{3}{8})^{2}\leq t\leq 0$, we define
$\widehat{\pi}_{(1)}^{i}(x,t)=\int_{\mathbb{T}^{3}}\nabla_{x}^{2}G(x-y)\eta(y)\varepsilon_{i}\left[\widehat{\textbf{u}}^{i}\otimes\widehat{\textbf{u}}^{i}+\nabla\widehat{{\bf{d}}}^{i}\odot\nabla\widehat{{\bf{d}}}^{i}-\frac{1}{2}|\nabla\widehat{{\bf{d}}}^{i}|^{2}{\rm I}_{3}-\frac{r_{i}^{2}}{\varepsilon_{i}^{2}}F({{\bf{d}}}^{i}){\rm I}_{3}\right](y,t){\,\mathrm{d}y},$
where $G$ is the Green function of $-\Delta$ for $\mathbb{T}^{3}$. Then it is straightforward to see that $\widehat{\pi}_{(2)}^{i}(\cdot,t):=(\widehat{\pi}^{i}-\widehat{\pi}_{(1)}^{i})(\cdot,t)$ satisfies
(5.12) $-\Delta\widehat{\pi}_{(2)}^{i}=0\text{ in }B_{\frac{3}{8}}(0).$
Applying the Calderon–Zygmund theory we can show that
(5.13) $\left\|\widehat{\pi}_{(1)}^{i}\right\|_{L^{\frac{3}{2}}(\mathbb{T}^{3})}\leq C\varepsilon_{i}\left[\left\|\widehat{\textbf{u}}^{i}\right\|_{L^{3}(B_{1}(0))}^{2}+\left\|\nabla\widehat{{\bf{d}}}^{i}\right\|_{L^{3}(B_{1}(0))}^{2}+\frac{r_{i}^{2}}{\varepsilon_{i}^{2}}\|F({\bf{d}}^{i})\|_{L^{\frac{3}{2}}(B_{1}(0))}\right]\leq C(\varepsilon_{i}+o(i)).$
From the mean value property of harmonic functions, we infer that for $0<\tau<\frac{1}{4}$,
(5.14) $\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}|\widehat{\pi}_{(2)}^{i}|^{\frac{3}{2}}\leq C\tau\int_{\mathcal{Q}_{\frac{1}{3}}(0)}|\widehat{\pi}_{(2)}^{i}|^{\frac{3}{2}}\leq C\tau\int_{\mathcal{Q}_{\frac{1}{3}}(0)}(|\widehat{\pi}^{i}|^{\frac{3}{2}}+|\widehat{\pi}_{(1)}^{i}|^{\frac{3}{2}})\leq C\tau(1+\varepsilon_{i}+o(i)).$
Combining (5.13) and (5.14) yields (5.10), and the contradiction (5.11) is achieved. ∎
5.2. An almost boundedness result
By iterating Lemma 5.2 and utilizing the Riesz potential in Morrey spaces (see [1, 27]) we can show the following local $L^{p}$ estimate for all $p>1$:
Lemma 5.3.
For any $M>0$, there exists $\varepsilon_{0}>0$ depending on $M$, such that for a suitable weak solution $({{\bf{v}}},{\pi},{{\bf{d}}})$ of (3.3) in ${\mathbb{T}^{3}}\times(0,\infty)$, if for $z_{0}=(x_{0},t_{0})\in{\mathbb{T}^{3}}\times(r_{0}^{2},\infty)$, and $r_{0}>0$, it holds that $|{\bf{z}}|\leq M,|{{\bf{d}}}|\leq M$ in $\mathcal{Q}_{r_{0}}(z_{0})$, and $\Theta_{z_{0}}(r_{0})\leq\varepsilon_{0}^{3}$, then for any $1<p<\infty$, it holds true that
$\|({\bf{v}},\pi,\nabla{{\bf{d}}})\|_{L^{p}(\mathcal{Q}_{\frac{r_{0}}{4}}(z_{0}))}\leq C(p,\varepsilon_{0},M).$
Proof.
From the assumption, we have that $\Phi_{z}(r_{0}/2)\leq 8\varepsilon_{0}^{3}$ holds for any $z\in\mathcal{Q}_{\frac{r_{0}}{2}}(z_{0})$. Then we apply Lemma 5.2 repeatedly on $\mathcal{Q}_{\frac{r_{0}}{2}}(z)$, and we have that for any $k\geq 1$ and $\tau_{0}\in(0,1/4)$,
$\Theta_{z}(\tau_{0}^{k}r_{0})\leq 2^{-k}\max\{\Theta_{z}\left(\frac{r_{0}}{2}\right),\frac{C_{0}r_{0}^{3}}{1-2\tau_{0}^{3}}\}.$
Hence, it holds for $0<s<\frac{r_{0}}{2}$, and $z\in\mathcal{Q}_{\frac{r_{0}}{2}}(z_{0})$ that
(5.15) $s^{-2}\int_{\mathcal{Q}_{s}(z)}(|{\bf{v}}|^{2}+|\nabla{{\bf{d}}}|^{3}+|\pi|^{\frac{3}{2}})\leq C(1+\varepsilon_{0}^{3})\left(\frac{s}{r_{0}}\right)^{3\theta_{0}},$
for $\theta_{0}=\frac{\ln 2}{3|\ln\tau_{0}|}\leq(0,\frac{1}{3})$. Now we apply the local energy inequality and boundedness of $({\bf{z}},{{\bf{d}}})$ on $\mathcal{Q}_{\frac{r_{0}}{2}}(z_{0})$, for $z\in\mathcal{Q}_{\frac{r_{0}}{2}}(z_{0})$ to show that
(5.16) $s^{-1}\int_{\mathcal{Q}_{s}(z)}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})$
$\leq$ $C\left[(2s)^{-3}\int_{\mathcal{Q}_{2s}(z)}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})+(2s)^{-2}\int_{\mathcal{Q}_{2s}(z)}(|{{\bf{v}}}|^{3}+|\nabla{{\bf{d}}}|^{2}+|P|^{\frac{3}{2}})\right]$
$\leq$ $C(1+\varepsilon_{0}^{3})\left(\frac{s}{r_{0}}\right)^{2\theta_{0}}.$
For any open set $U\subset\mathbb{T}^{3}\times\mathbb{R}_{+}$, $1\leq p<\infty$, $0\leq\lambda\leq 5$, the Morrey space $M^{p,\lambda}$ is defined by
$M^{p,\lambda}(U):=\left\{f\in L^{p}_{\rm loc}(U):\|f\|^{p}_{M^{p,\lambda}(U)}=\sup_{z\in U,r>0}r^{\lambda-5}\int_{\mathcal{Q}_{r}(z)}|f|^{p}\,\mathrm{d}x\,\mathrm{d}t<\infty\right\}.$
Then (5.15) and (5.16) read
[…]
Now we view (1.1)${}_{3}$ as
$\partial_{t}{{\bf{d}}}-\Delta{{\bf{d}}}=-{{\bf{u}}}\cdot\nabla{{\bf{d}}}-f({{\bf{d}}})\in M^{\frac{3}{2},3(1-\alpha)}\left(\mathcal{Q}_{\frac{r_{0}}{2}}(z_{0})\right).$
Let $\eta$ be the cut off function of $\mathcal{Q}_{\frac{r_{0}}{2}}(z_{0})$ with $|\partial_{t}\eta|+|\nabla^{2}\eta|\leq Cr_{0}^{-2}$, let $\widecheck{\bf d}=\eta^{2}({{\bf{d}}}-({{\bf{d}}})_{z_{0},r_{0}})$, where $({{\bf{d}}})_{z_{0},r_{0}}$ is the average of ${{\bf{d}}}$ over $\mathcal{Q}_{\frac{r_{0}}{2}}(z_{0})$. Then we have
$\partial_{t}\widecheck{{\bf{d}}}-\Delta\widecheck{\bf d}=\mathsf{F},\quad\mathsf{F}:=\eta^{2}(-{{\bf{u}}}\cdot\nabla{{\bf{d}}}-f({{\bf{d}}}))+(\partial_{t}\eta^{2}-\Delta\eta^{2})({{\bf{d}}}-({{\bf{d}}})_{z_{0},r_{0}})-\nabla\eta^{2}\cdot\nabla{{\bf{d}}},$
where $\mathbb{F}$ satisfies
$\|\mathsf{F}\|_{M^{\frac{3}{2},3(1-\alpha)}(\mathbb{R}^{4})}\leq C(1+\varepsilon_{0}).$
By the Duhamel formula, we have that
[…]
where $\Gamma$ is the heat kernel in $\mathbb{R}^{3}$, and $\mathcal{I}_{\beta}$ is the Riesz potential of order $\beta$ on $\mathbb{R}^{4}$, $\beta\in[0,4]$, defined by
[…]
Apply the Riesz potential estimates, we conclude that
(5.17) $\|\nabla{\widecheck{\bf d}}\|_{M^{\frac{3(1-\alpha)}{1-2\alpha},3(1-\alpha)}(\mathbb{R}^{4})}\leq C\|\mathsf{F}\|_{M^{\frac{3}{2},3(1-\alpha)}(\mathbb{R}^{4})}\leq C(1+\varepsilon_{0}).$
Since ${{\bf{d}}}-{\widecheck{\bf d}}$ is caloric in $\mathcal{Q}_{\frac{r_{0}}{2}}(z_{0})$,
(5.18) $\|\nabla({{\bf{d}}}-{\widecheck{\bf d}})\|_{L^{p}(\mathcal{Q}_{\frac{r_{0}}{2}(z_{0})})}\leq C(p,r_{0},\varepsilon_{0}).$
On the other hand, let $\widecheck{{\bf{v}}}$ solve the following Stokes equations
(5.19) $\left\{\begin{array}[]{l}\partial_{t}\widecheck{{\bf{v}}}-\Delta\widecheck{{\bf{v}}}+\nabla\widecheck{\pi}\\ =-\nabla\cdot[\eta^{2}({{\bf{u}}}\otimes{{\bf{u}}})+(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}|\nabla{{\bf{d}}}|^{2}{\rm I}_{3})-(F({{\bf{d}}})-(F({{\bf{d}}}))_{z_{0},r_{0}}){\rm I}_{3}]=:-\mathop{\rm div}\nolimits\mathsf{X},\\ \nabla\cdot\widecheck{{\bf{v}}}=0,\\ \widecheck{{\bf{v}}}(\cdot,0)=0.\end{array}\right.$
By using the Oseen kernel, we can estimate $\widecheck{{\bf{v}}}$ by
$|\widecheck{{\bf{v}}}(x,t)|\leq C\mathcal{I}_{1}(|\mathsf{X}|)(x,t).$
Note that
$\displaystyle\|\mathsf{X}\|_{M^{\frac{3}{2},3(1-\alpha)}}\leq C(1+\varepsilon_{0}).$
This yields that
(5.20) $\|\widecheck{{\bf{v}}}\|_{M^{\frac{3(1-\alpha)}{1-2\alpha},3(1-\alpha)}(\mathbb{R}^{4})}\leq C\|\mathsf{X}\|_{M^{\frac{3}{2},3(1-\alpha)}(\mathbb{R}^{4})}\leq C(1+\varepsilon_{0}).$
Meanwhile, we have $({{\bf{v}}}-\widecheck{{\bf{v}}})$ solves the linear homogeneous Stokes equation in $\mathcal{Q}_{\frac{r_{0}}{2}}(z_{0})$ and
(5.21) $\|{{\bf{v}}}-\widecheck{{\bf{v}}}\|_{L^{p}(\mathcal{Q}_{\frac{r_{0}}{4}}(z_{0}))}\leq C(p,r_{0},\varepsilon_{0}).$
Bootstrapping $\alpha$ in the previous estimates (5.17), (5.18), (5.20) and (5.21) to get $\alpha\uparrow\frac{1}{2}$, and then by the embedding theorem for the Morrey spaces, we obtain that
$\|({{\bf{v}}},\nabla{{\bf{d}}})\|_{L^{p}(\mathcal{Q}_{\frac{r_{0}}{4}}(z_{0}))}\leq C(p,r_{0},\varepsilon_{0}).$
The estimate for ${\pi}$ comes directly from the standard $L^{p}$ estimate for the Poisson equation:
$-\Delta{\pi}=\mathop{\rm div}\nolimits^{2}[{{\bf{u}}}\otimes{{\bf{u}}}+(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}|\nabla{{\bf{d}}}|^{2}{\rm I}_{3})-(F({{\bf{d}}})-(F({{\bf{d}}}))_{z_{0},r_{0}})].$
5.3. The A-B-C-D Lemma and Criteria for the regular points
Now we are ready to verify the first criteria for the regular points:
Proposition 5.4.
If $z_{0}$ satisfies the assumption in Lemma 5.2, then $z_{0}$ is a regular point.
Proof.
By Lemma 5.3, we have that $({{\bf{u}}},{\pi},\nabla{{\bf{d}}})\in L^{p}(\mathcal{Q}_{r_{0}/4}(z_{0}))$ for $1<p<\infty$. Applying the regularity estimate for the generalized Stokes system (c.f. [28, Lemma 2.2]) to the ${\bf{v}}$ equation in (3.3)${}_{1}$, we get that ${\bf{v}}\in C^{\alpha}(\mathcal{Q}_{\frac{r_{0}}{8}}(z_{0}))$ for some $\alpha\in(0,1)$. On the other hand, the ${{\bf{d}}}$ equation reads
$\partial_{t}{{\bf{d}}}-\Delta{{\bf{d}}}=-({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{d}}}-f({{\bf{d}}})\in L^{p}(\mathcal{Q}_{\frac{r_{0}}{4}}(z_{0})).$
By the standard $W^{2,p}$ estimate we can get that $\|(\partial_{t}{{\bf{d}}},\nabla^{2}{{\bf{d}}})\|_{L^{p}(\mathcal{Q}_{\frac{r_{0}}{8}}(z_{0}))}<C(p,r_{0},\varepsilon_{0})$. And the Sobolev embedding implies $\nabla{{\bf{d}}}\in C^{\alpha}(\mathcal{Q}_{\frac{r_{0}}{8}}(z_{0})),$ thus $z_{0}$ is a regular point. ∎
In fact we can improve Proposition 5.4 into the following theorem:
Theorem 5.5.
There exists $\varepsilon_{1}>0$ such that if $({{\bf{u}}},P,{{\bf{d}}})$ is a suitable weak solution to (1.1), which satisfies, for $z_{0}\in\mathbb{T}^{3}\times(0,\infty)$,
(5.22) $\limsup_{r\to 0}\frac{1}{r}\int_{\mathcal{Q}_{r}(z_{0})}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})\leq\varepsilon_{1}^{2},$
then $({{\bf{u}}},\nabla{{\bf{d}}})$ is bounded near $z_{0}$.
For simplicity, we assume that $z_{0}=0$. We generalize the dimensionless estimates in [7, 35] with the following quantities:
$\displaystyle A(r)$ $\displaystyle:=\sup_{-r^{2}\leq t\leq 0}r^{-1}\int_{B_{r}(0)\times\{t\}}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}),$
$\displaystyle B(r)$ $\displaystyle:=r^{-1}\int_{\mathcal{Q}_{r}(0)}\left(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2}\right),$
$\displaystyle C(r)$ $\displaystyle:=r^{-2}\int_{Q_{r}(0)}(|{{\bf{v}}}|^{3}+|\nabla{{\bf{d}}}|^{3}),$
$\displaystyle D(r)$ $\displaystyle:=r^{-2}\int_{\mathcal{Q}_{r}(0)}|{\pi}|^{\frac{3}{2}},$
To prove Theorem 5.5, we will need some preparations. The first step is to note that we have the following interpolation lemma.
Lemma 5.6.
For $v\in H^{1}(\mathbb{T}^{3})$, we have
(5.23) $\int_{B_{r}(0)}|v|^{q}(x,t){\,\mathrm{d}x}\lesssim\left(\int_{B_{r}(0)}|\nabla v|^{2}(x,t){\,\mathrm{d}x}\right)^{\frac{q}{2}-a}\left(\int_{B_{r}(0)}|v|^{2}(x,t)\right)^{a}$
$+r^{3(1-\frac{q}{2})}\left(\int_{B_{r}(0)}|v|^{2}(x,t){\,\mathrm{d}x}\right)^{\frac{q}{2}},$
for every $B_{r}(0)\subset\mathbb{T}^{3}$, $2\leq q\leq 6$, $a=\frac{3}{2}(1-\frac{q}{6}).$
With application of Lemma 5.6 we can establish the following result.
Lemma 5.7.
For any ${{\bf{u}}}\in L^{\infty}([-\rho^{2},0];L^{2}(B_{\rho}(0)))\cap L^{2}([-\rho^{2},0];H^{1}(B_{\rho}(0)))$, ${{\bf{d}}}\in L^{\infty}([-\rho^{2},0];H^{1}(B_{\rho}(0)))\cap L^{2}([-\rho^{2},0];H^{2}(B_{\rho}(0)))$, it holds that for any $0<r\leq\rho$,
(5.24) $C(r)\lesssim\left(\frac{r}{\rho}\right)^{3}A^{\frac{3}{2}}(\rho)+\left(\frac{\rho}{r}\right)^{3}A^{\frac{3}{4}}(\rho)B^{\frac{3}{4}}(\rho).$
Proof.
By Poincaré’s inequality, we can show that for $0<r\leq\rho$,
$\displaystyle\int_{B_{r}(0)}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})$
$\displaystyle\lesssim\int_{B_{r}(0)}\left(\left||{{\bf{v}}}|^{2}-(|{{\bf{v}}}|^{2})\rho\right|+\left||\nabla{{\bf{d}}}|^{2}-(|\nabla{{\bf{d}}}|^{2})_{\rho}\right|\right){\,\mathrm{d}x}+\left(\frac{r}{\rho}\right)^{3}\int_{B_{\rho}(0)}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}){\,\mathrm{d}x}$
$\displaystyle\lesssim\rho\int_{B_{\rho}(0)}(|{{\bf{v}}}||\nabla{{\bf{v}}}|+|\nabla{{\bf{d}}}||\nabla^{2}{{\bf{d}}}|){\,\mathrm{d}x}+\left(\frac{r}{\rho}\right)^{3}\int_{B_{\rho}(0)}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}){\,\mathrm{d}x}$
$\displaystyle\lesssim\rho^{\frac{3}{2}}\left(\rho^{-1}\int_{B_{\rho}(0)}|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}{\,\mathrm{d}x}\right)^{\frac{1}{2}}\left(\int_{B_{\rho}(0)}|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2}\right)^{\frac{1}{2}}+\left(\frac{r}{\rho}\right)^{3}\int_{B_{\rho}(0)}\left(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}\right){\,\mathrm{d}x}$
$\displaystyle\lesssim\rho^{\frac{3}{2}}A^{\frac{1}{2}}(\rho)\left(\int_{B_{\rho}(0)}|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2}{\,\mathrm{d}x}\right)^{\frac{1}{2}}+\left(\frac{r}{\rho}\right)^{3}A(\rho).$
Applying Lemma 5.6 with $q=3$, $a=\frac{3}{4}$ we can get
$\displaystyle\int_{B_{r}(0)}(|{{\bf{v}}}|^{3}+|\nabla{{\bf{d}}}|^{3}){\,\mathrm{d}x}$
$\displaystyle\lesssim\rho^{\frac{4}{3}}\left(\int_{B_{r}(0)}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2}){\,\mathrm{d}x}\right)^{\frac{3}{4}}\left(\rho^{-1}\int_{B_{r}(0)}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}){\,\mathrm{d}x}\right)^{\frac{3}{4}}+r^{-\frac{3}{2}}\left(\int_{B_{r}(0)}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}){\,\mathrm{d}x}\right)^{\frac{3}{2}}$
$\displaystyle\lesssim\rho^{\frac{3}{4}}A^{\frac{3}{4}}(\rho)\left(\int_{B_{r}(0)}|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2}\right)^{\frac{3}{4}}+r^{-\frac{3}{2}}\left(\int_{B_{r}(0)}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})\right)^{\frac{3}{2}}$
$\displaystyle\lesssim\left(\rho^{\frac{3}{4}}+\frac{\rho^{\frac{9}{4}}}{r^{\frac{3}{2}}}\right)\left(\int_{B_{r}(0)}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})\right)^{\frac{3}{4}}A^{\frac{3}{4}}(\rho)+\left(\frac{r}{\rho}\right)^{3}A^{\frac{3}{2}}(\rho),$
where in the last we apply the previous inequality. Integrating the inequality over $[-r^{2},0]$ and by Hölder’s inequality, we have
$\displaystyle C(r)$ $\displaystyle=\frac{1}{r^{2}}\int_{\mathcal{Q}_{r}(0)}(|{{\bf{v}}}|^{3}+|\nabla{{\bf{d}}}|^{3})$
$\displaystyle\lesssim\left(\frac{r}{\rho}\right)^{3}A^{\frac{3}{2}}(\rho)+\left(\rho^{\frac{3}{4}}+\frac{\rho^{\frac{9}{4}}}{r^{\frac{3}{2}}}\right)\int_{-r^{2}}^{0}\left(\int_{B_{r}(0)}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})\right)^{\frac{3}{4}}A^{\frac{3}{4}}(\rho)$
$\displaystyle\lesssim\left(\frac{r}{\rho}\right)^{3}A^{\frac{3}{2}}(\rho)+r^{-\frac{3}{2}}\rho^{\frac{3}{4}}\left(\rho^{\frac{3}{4}}+\frac{\rho^{\frac{9}{4}}}{r^{\frac{3}{2}}}\right)A^{\frac{3}{4}}(\rho)B^{\frac{3}{4}}(\rho)$
$\displaystyle=\left(\frac{r}{\rho}\right)^{3}A^{\frac{3}{2}}(\rho)+\left[\left(\frac{\rho}{r}\right)^{\frac{3}{2}}+\left(\frac{\rho}{r}\right)^{3}\right]A^{\frac{3}{4}}(\rho)B^{\frac{3}{4}}(\rho)$
$\displaystyle\lesssim\left(\frac{r}{\rho}\right)^{3}A^{\frac{3}{2}}(\rho)+\left(\frac{\rho}{r}\right)^{3}A^{\frac{3}{4}}(\rho)B^{\frac{3}{4}}(\rho).$
Next, we estimate the pressure function to get
Lemma 5.8.
Under the same assumption of Lemma 5.7, it holds that for any $2\displaystyle 0<r<\rho/$ 2,
(5.25) $D(r)\lesssim\frac{r}{\rho}D(\rho)+\left(\frac{\rho}{r}\right)^{2}\left[A^{\frac{3}{4}}(\rho)B^{\frac{3}{4}}(\rho)+\rho^{\frac{3}{2}}B^{\frac{3}{4}}(\rho)+\rho^{3}\right].$
Proof.
Taking the divergence of (3.3)${}_{1}$ yields
(5.26) $-\Delta{\pi}$ $=\mathop{\rm div}\nolimits^{2}\Big{[}{{\bf{v}}}\otimes{{\bf{v}}}+\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}(\nabla{{\bf{d}}}:\nabla{{\bf{d}}}){\rm I}_{3}+{{\bf{z}}}\otimes{{\bf{v}}}+{{\bf{v}}}\otimes{{\bf{z}}}+{{\bf{z}}}\otimes{{\bf{z}}}-F({{\bf{d}}}){\rm I}_{3}\Big{]}$
$=\mathop{\rm div}\nolimits^{2}\Big{[}({{\bf{v}}}-({{\bf{v}}})_{\rho})\otimes({{\bf{v}}}-({{\bf{v}}})_{\rho})+(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho})\odot(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho})$
$-\frac{1}{2}(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}):(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}){\rm I}_{3}$
$+(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho})\odot(\nabla{{\bf{d}}})_{\rho}+(\nabla{{\bf{d}}})_{\rho}\odot(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho})-\frac{1}{2}(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}):(\nabla{{\bf{d}}})_{\rho}{\rm I}_{3}$
$-\frac{1}{2}(\nabla{{\bf{d}}})_{\rho}\odot(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}){\rm I}_{3}$
$+{{\bf{z}}}\otimes({{\bf{v}}}-({{\bf{v}}})_{\rho})+({{\bf{v}}}-({{\bf{v}}})_{\rho})\otimes{{\bf{z}}}+{{\bf{z}}}\otimes{{\bf{z}}}-F({{\bf{d}}}){\rm I}_{3}\Big{]}$
$:=\mathop{\rm div}\nolimits^{2}\mathcal{G}\text{ in }B_{\rho}.$
Here $({{\bf{v}}})_{\rho}$ denotes the average of ${{\bf{v}}}$ over $B_{\rho}$. Let $\eta\in C_{0}^{\infty}(\mathbb{T}^{3})$ be a cut-off function of $B_{\rho/2}$ such that $|\nabla\eta|\lesssim\rho^{-1}$. Define an auxiliary function by
$\displaystyle\widecheck{\pi}$ $\displaystyle(x)=-\int_{\mathbb{T}^{3}}\nabla_{y}^{2}G(x-y):\eta^{2}(y)\mathcal{G}(y){\,\mathrm{d}y},$
where $G$ is the corresponding kernel of Laplacian on $\mathbb{T}^{3}$. Then we have
$-\Delta\widecheck{\pi}=\mathop{\rm div}\nolimits^{2}\mathcal{G}\text{ in }B_{\rho/2},$
and
$-\Delta({\pi}-\widecheck{\pi})=0\text{ in }B_{\rho/2}.$
From the singular integral theory (Calderon–Zygmund theory) and the $L^{\infty}$ bound on $({{\bf{z}}},{{\bf{d}}})$ we can infer that
$\displaystyle\int_{\mathbb{T}^{3}}|\widecheck{\pi}|^{\frac{3}{2}}{\,\mathrm{d}x}$ $\displaystyle\lesssim\int_{\mathbb{T}^{3}}\eta^{3}\left(|{{\bf{v}}}-({{\bf{v}}})_{\rho}|^{3}+|\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}|^{3}\right)$
$\displaystyle+|(\nabla{{\bf{d}}})_{\rho}|^{\frac{3}{2}}\int_{\mathbb{T}^{3}}\eta^{3}|\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}|^{\frac{3}{2}}+\int_{\mathbb{T}^{3}}\eta^{3}|{{\bf{v}}}-({{\bf{v}}})_{\rho}|^{\frac{3}{2}}+\int_{\mathbb{T}^{3}}\eta^{3}$
$\displaystyle\lesssim\int_{B_{\rho}}(|{{\bf{v}}}-({{\bf{v}}})_{\rho}|^{3}+|\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}|^{3})+|(\nabla{{\bf{d}}})_{\rho}|^{\frac{3}{2}}\int_{B_{\rho}}|\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}|^{\frac{3}{2}}+\int_{B_{\rho}}(|{{\bf{v}}}-({{\bf{v}}})_{\rho}|^{\frac{3}{2}}+1)$
Notice that ${\pi}-\widecheck{\pi}$ is harmonic in $B_{\rho/2}$, and we have that for $2\displaystyle 0<r\leq\rho/$ 2,
$\displaystyle\frac{1}{r^{2}}\int_{B_{r}}|{\pi}-\widecheck{\pi}|^{\frac{3}{2}}$ $\displaystyle\lesssim\left(\frac{r}{\rho}\right)\frac{1}{\rho^{2}}\int_{B_{\rho/2}}|{\pi}-\widecheck{\pi}|^{\frac{3}{2}}$
$\displaystyle\lesssim\left(\frac{r}{\rho}\right)\left[\frac{1}{\rho^{2}}\int_{B_{\rho/2}}|{\pi}|^{\frac{3}{2}}+\frac{1}{\rho^{2}}\int_{B_{\rho/2}}|\widecheck{\pi}|^{\frac{3}{2}}\right].$
Integrating the expression above over $[-r^{2},0]$, we can show that
$\displaystyle\frac{1}{r^{2}}\int_{\mathcal{Q}_{r}(0)}|{\pi}|^{\frac{3}{2}}$ $\displaystyle\lesssim\left(\frac{r}{\rho}\right)\frac{1}{\rho^{2}}\int_{\mathcal{Q}_{\rho}}|{\pi}|^{\frac{3}{2}}+\left(\frac{\rho}{r}\right)^{2}\frac{1}{\rho^{2}}\int_{\mathcal{Q}_{\rho}}|\widecheck{\pi}|^{\frac{3}{2}}$
$\displaystyle\lesssim\left(\frac{r}{\rho}\right)D(\rho)+\left(\frac{\rho}{r}\right)^{2}\Big{[}\frac{1}{\rho^{2}}\int_{\mathcal{Q}_{\rho}}|{{\bf{v}}}-({{\bf{v}}})_{\rho}|^{3}+|\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}|^{3}$
$\displaystyle+\frac{1}{\rho^{2}}(\sup_{-\rho^{2}\leq t\leq 0}|(\nabla{{\bf{d}}}(t))_{\rho}|^{\frac{3}{2}})\int_{\mathcal{Q}_{\rho}}|\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}|^{\frac{3}{2}}+\frac{1}{\rho^{2}}\int_{\mathcal{Q}_{\rho}}|{{\bf{v}}}-({{\bf{v}}})_{\rho}|^{\frac{3}{2}}+\rho^{3}\Big{]}.$
Now we apply the interpolation inequality
$\displaystyle\frac{1}{\rho^{2}}\int_{\mathcal{Q}_{\rho}}(|{{\bf{v}}}-({{\bf{v}}})_{\rho}|^{3}+|\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}|^{3})$
$\displaystyle\lesssim\left(\sup_{-\rho^{2}\leq t\leq 0}\frac{1}{\rho}\int_{B_{\rho}}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})\right)^{\frac{3}{4}}\left(\frac{1}{\rho}\int_{\mathcal{Q}_{\rho}}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})\right)^{\frac{3}{4}}=A^{\frac{3}{4}}(\rho)B^{\frac{3}{4}}(\rho),$
and by Hölder’s inequality and Poincaré’s inequality, we can show
$\displaystyle\frac{1}{\rho^{2}}(\sup_{-\rho^{2}\leq t\leq 0}|(\nabla{{\bf{d}}})_{\rho}|^{\frac{3}{2}})\int_{\mathcal{Q}_{\rho}}|\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}|^{\frac{3}{2}}$
$\displaystyle\lesssim\rho^{2}\left(\sup_{-\rho^{2}\leq t\leq 0}\int_{B_{\rho}}|\nabla{{\bf{d}}}|^{2}\right)^{\frac{3}{4}}\left(\frac{1}{\rho}\int_{Q_{\rho}}|\nabla^{2}{{\bf{d}}}|^{2}\right)^{\frac{3}{4}}\leq A^{\frac{3}{4}}(\rho)B^{\frac{3}{4}}(\rho),$
$\displaystyle\frac{1}{\rho^{2}}\int_{\mathcal{Q}_{\rho}}|{{\bf{v}}}-({{\bf{v}}})_{\rho}|^{\frac{3}{2}}\lesssim\frac{1}{\rho^{2}}\left(\int_{\mathcal{Q}_{\rho}}|{{\bf{v}}}-({{\bf{v}}})_{\rho}|^{2}\right)^{\frac{3}{4}}\lesssim\rho^{\frac{3}{2}}B^{\frac{3}{4}}(\rho).$
Now we have the interpolation estimate (5.24) and the pressure estimate (5.25), the proof of Theorem 5.5 is based on a dichotomy argument:
Proof of Theorem 5.5.
. For $\theta\in(0,\frac{1}{2})$ and $\rho\in(0,1)$, let $\varphi\in C_{0}^{\infty}(\mathcal{Q}_{\theta\rho}(0))$ be a function such that
$\varphi=1\text{ in }\mathcal{Q}_{\theta\rho/2}(0),\qquad|\nabla\varphi|\lesssim(\theta\rho)^{-1},\qquad|\nabla^{2}\varphi|+|\partial_{t}\varphi|\lesssim(\theta\rho)^{-2}.$
Applying the local energy inequality in Lemma 4.1 with $\varphi^{2}$, the $L^{\infty}$-bounded of $({{\bf{z}}},{{\bf{d}}})$, and integrating by parts, we obtain that
$\displaystyle\sup_{-(\theta\rho)^{2}\leq t\leq 0}\int_{\mathbb{T}^{3}}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})\varphi^{2}+\int_{\mathbb{T}^{3}\times[-(\theta\rho)^{2},0]}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})\varphi^{2}$
$\displaystyle\lesssim\int_{\mathbb{T}^{3}\times[-(\theta\rho)^{2},0]}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})(|\partial_{t}\varphi|+|\nabla\varphi|^{2}+|\nabla^{2}\varphi|)$
$\displaystyle+\int_{\mathbb{T}^{3}\times[-(\theta\rho)^{2},0]}[(|{{\bf{v}}}|^{2}-(|{{\bf{v}}}|^{2})_{\rho})+|{\pi}||{{\bf{v}}}|]|\nabla\varphi|$
$\displaystyle+\int_{\mathbb{T}^{3}\times[-(\theta\rho)^{2},0]}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})\varphi^{2}+|\nabla{\bf{d}}|^{2}|{{\bf{v}}}||\nabla\varphi|+\int_{\mathbb{T}^{3}\times[-(\theta\rho)^{2},0]}|{{\bf{z}}}|^{2}|\nabla\varphi|$
$\displaystyle+\int_{\mathbb{T}^{3}\times[-(\theta\rho)^{2},0]}\varphi^{2}.$
Hence we obtain that
$\displaystyle A\left(\frac{1}{2}\theta\rho\right)+B\left(\frac{1}{2}\theta\rho\right)$
$\displaystyle=\sup_{-\left(\frac{\theta\rho}{2}\right)^{2}\leq t\leq 0}\frac{2}{\theta\rho}\int_{B_{\frac{\theta\rho}{2}}}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})+\frac{2}{\theta\rho}\int_{\mathcal{Q}_{\frac{\theta\rho}{2}}(0)}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})$
$\displaystyle\lesssim\sup_{-(\theta\rho)^{2}\leq t\leq 0}\frac{1}{\theta\rho}\int_{\mathbb{T}^{3}}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})\varphi^{2}+\frac{1}{\theta\rho}\int_{\mathbb{T}^{3}\times[-(\theta\rho)^{2},0]}(|\nabla{{\bf{u}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})\varphi^{2}$
Putting together all the estimates, we have
$\displaystyle A(\frac{1}{2}\theta\rho)+B(\frac{1}{2}\theta\rho)$ $\displaystyle\lesssim\left[C^{\frac{2}{3}}(\theta\rho)+A^{\frac{1}{2}}(\theta\rho)B^{\frac{1}{2}}(\theta\rho)C^{\frac{1}{3}}(\theta\rho)+C^{\frac{1}{3}}(\theta\rho)D^{\frac{2}{3}}(\theta\rho)+(\theta\rho)^{3}\right]$
$\displaystyle\lesssim C^{\frac{2}{3}}(\theta\rho)+A(\theta\rho)B(\theta\rho)+D^{\frac{4}{3}}(\theta\rho)+(\theta\rho)^{3}.$
So that
$\displaystyle A^{\frac{3}{2}}(\frac{1}{2}\theta\rho)$ $\displaystyle\lesssim C(\theta\rho)+A^{\frac{3}{2}}(\theta\rho)B^{\frac{3}{2}}(\theta\rho)+D^{2}(\theta\rho)+(\theta\rho)^{\frac{9}{2}}.$
On the other hand, from Lemma 5.7 and 5.8 we can conclude that
$\displaystyle C(\theta\rho)$ $\displaystyle\lesssim\theta^{3}A^{\frac{3}{2}}(\rho)+\theta^{-3}A^{\frac{3}{4}}(\rho)B^{\frac{3}{4}}(\rho),$
$\displaystyle D^{2}(\theta\rho)$ $\displaystyle\lesssim\theta^{2}D^{2}(\rho)+\theta^{-4}(A^{\frac{3}{2}}(\rho)B^{\frac{3}{2}}(\rho)+\rho^{3}B^{\frac{3}{2}}(\rho)+\rho^{6}),$
and it is easy to see that
$A^{\frac{3}{2}}(\theta\rho)B^{\frac{3}{2}}(\theta\rho)\lesssim\theta^{-3}A^{\frac{3}{2}}(\rho)B^{\frac{3}{2}}(\rho).$
Therefore we conclude that for $0<\theta<\frac{1}{2}$,
$\displaystyle A^{\frac{3}{2}}(\frac{1}{2}\theta\rho)+D^{2}(\frac{1}{2}\theta\rho)$
$\displaystyle\leq c\left[\theta^{2}D^{2}(\rho)+\theta^{-4}(A^{\frac{3}{2}}(\rho)B^{\frac{3}{2}}(\rho)+\rho^{3}B^{\frac{3}{2}}(\rho)+\rho^{6})+\theta^{3}A^{\frac{3}{2}}(\rho)+\theta^{-8}A^{\frac{3}{2}}(\rho)B^{\frac{3}{2}}(\rho)+\theta^{2}+(\theta\rho)^{\frac{9}{2}}\right]$
$\displaystyle\leq c(\theta^{2}+\theta^{-8}B^{\frac{3}{2}}(\rho))(A^{\frac{3}{2}}(\rho)+D^{2}(\rho))+c(\theta^{-4}\rho^{3}B^{\frac{3}{2}}(\rho)+\theta^{-4}\rho^{6}+\theta^{2}+(\theta\rho)^{\frac{9}{2}}).$
For $\varepsilon_{0}$ given by Lemma 5.2, let $\theta_{0}\in(0,\frac{1}{2})$ be such that
$c\theta_{0}=\min\{\frac{1}{4},\frac{1}{8}\varepsilon_{1}^{2}\}.$
Since
$\limsup_{r\to 0}\frac{1}{r}\int_{\mathcal{Q}_{r}}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})<\varepsilon_{1}^{2}$
we can choose $\rho_{0}>0$ such that
$c\theta_{0}^{-8}B^{\frac{3}{2}}(\rho)\leq\frac{1}{4}\qquad\forall 0<\rho<\rho_{0},$
and
$c(\theta_{0}^{-4}\rho^{3}B^{\frac{3}{2}}(\rho)+\theta_{0}^{-4}\rho^{6}+\theta_{0}^{2}+(\theta_{0}\rho)^{\frac{9}{2}})\leq\frac{1}{2}\varepsilon_{1}^{2}\qquad\forall 0<\rho<\rho_{0}.$
Therefore we obtained that
$A^{\frac{3}{2}}(\frac{1}{2}\theta_{0}\rho)+D^{2}(\frac{1}{2}\theta_{0}\rho)\leq\frac{1}{2}(A^{\frac{3}{2}}(\rho)+D^{2}(\rho))+\frac{1}{2}\varepsilon_{1}^{2},\qquad\forall 0<\rho<\rho_{0}.$
Iterating this inequality yields that
$A^{\frac{3}{2}}((\frac{1}{2}\theta_{0})^{k}\rho)+D^{2}((\frac{1}{2}\theta_{0})^{k}\rho)\leq\frac{1}{2^{k}}(A^{\frac{3}{2}}(\rho)+D^{2}(\rho))+\varepsilon_{1}^{2}$
holds for all $0<\rho<\rho_{0}$ and $k\geq 1$. Furthermore, we can obtain from (5.24) that
$\displaystyle C((\frac{1}{2}\theta_{0})^{k}\rho)$ $\displaystyle\leq c[(\frac{1}{2}\theta_{0})^{3}A^{\frac{3}{2}}((\frac{1}{2}\theta_{0})^{k-1}\rho)+(\frac{1}{2}\theta_{0})^{-3}A^{\frac{3}{4}}((\frac{1}{2}\theta_{0})^{k-1}\rho)B^{\frac{3}{4}}((\frac{1}{2}\theta_{0})^{k-1}\rho)]$
$\displaystyle\leq c[(\frac{1}{2}\theta_{0})^{3}+(\frac{1}{2}\theta_{0})^{-3}\varepsilon_{1}^{\frac{3}{2}}][(\frac{1}{2})^{k-1}(A^{\frac{3}{2}}(\rho)+D^{2}(\rho))+\varepsilon_{1}^{2}]$
holds for all $0<\rho<\rho_{0}$ and $k\geq 1$. Then we can arrive at
$\limsup_{k\to\infty}\left[C((\frac{1}{2}\theta_{0})^{k}\rho)+D^{2}((\frac{1}{2}\theta_{0})^{k}\rho)\right]\leq c[1+(\frac{1}{2}\theta_{0})^{3}+(\frac{1}{2}\theta_{0})^{-3}\varepsilon_{1}^{\frac{3}{2}}]\varepsilon_{1}^{2}\leq\frac{1}{2}\varepsilon_{0}^{3}$
holds for all $\rho\in(0,\rho_{0})$ provided $\varepsilon_{1}=\varepsilon_{1}(\theta_{0},\varepsilon_{0})>0$ is chosen sufficiently small. Hence there exists a $k_{0}>0$ such that
$C((\frac{1}{2}\theta_{0})^{k_{0}}\rho)+D((\frac{1}{2}\theta_{0})^{k_{0}}\rho)\leq\varepsilon_{0}^{3}$
which satisfies the assumption of Lemma 5.2, and as a consequence, $z_{0}=0$ is a regular point. ∎
5.4. The Hausdorff measure estimate
Now we have all the ingredients to prove Theorem 3.3.
Proof of Theorem 3.3.
If $z\in\Sigma$, then by Theorem 5.5,
$\limsup_{r\to 0}\frac{1}{r}\int_{\mathcal{Q}_{r}(z)}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})\geq\varepsilon_{1}.$
Let $V$ be a neighborhood of $\Sigma$ and $\eta>0$ such that for all $z\in\Sigma$, then there exists a $r<\eta$ such that $\mathcal{Q}_{\eta}(z)\subset V$ and
$\frac{1}{r}\int_{\mathcal{Q}_{r}(z)}(|\nabla{{\bf{u}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})\geq\varepsilon_{1}.$
By Vitali’s covering lemma, we can find a pairwisely disjoint collection of $\mathcal{Q}_{r_{i}}(z_{i})$ such that
$\Sigma\subset\bigcup_{i=1}^{\infty}\mathcal{Q}_{5r_{i}}(z_{i}).$
Hence,
(5.27) $\mathcal{P}_{5\eta}^{1}(\Sigma)$ $\leq\sum_{i=1}^{\infty}5r_{i}\leq\frac{5}{\varepsilon_{1}}\sum_{i=1}^{\infty}\int_{\mathcal{Q}_{r_{i}}(z_{i})}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})$
$\leq\frac{5}{\varepsilon_{1}}\int_{\bigcup_{i}\mathcal{Q}_{r_{i}}(z_{i})}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})$
$\leq\frac{5}{\varepsilon_{1}}\int_{V}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})<\infty.$
Then we can conclude that $\Sigma$ is of zero Lesbegue measure $\,\mathrm{d}x\,\mathrm{d}t$. Furthermore, we can choose $V$ with arbitrarily small measure, then by the absolute continuity of integral for $|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2}$, we can conclude from (5.27) that $\mathcal{P}^{1}(\Sigma)=0$. ∎