### v2

# Partial Regularity for the Three-Dimensional Stochastic Ericksen–Leslie Equations

Abstract
In this article, we investigate the global existence of martingale suitable weak solutions to stochastic Ericksen–Leslie equations with additive noise in a 3D torus. The notion of suitable weak solutions has been introduced to address possible emergence of finite-time singularities, which remains a notably challenging question in the field of fluid dynamics. Weak solutions offer an approach to account for these potential singularities. A restricted class of weak solutions that exhibit a higher level of regularity which are therefore more likely to be physically meaningful, is naturally called for. Consequently, suitable weak solutions, i.e., weak solutions that satisfy a local energy inequality, become a focus of research, including investigations about how regular these solutions can be. In this article, we prove that, despite the presence of white noise, the paths of martingale suitable weak solutions of 3D stochastic Ericksen–Leslie equations exhibit singular points of one-dimensional parabolic Hausdorff measure zero. To establish this result, we have utilized two techniques, which can potentially be generalized to handle other stochastically forced complex fluid dynamics equations with a similar structure. Firstly, a local energy-preserving approximation is constructed which markedly facilitates the proof of the global existence of martingale suitable weak solutions; secondly, to demonstrate partial regularity of these solutions, a blow-up argument is formulated, which efficiently yields the desired key estimate.
keywords:
Stochastic PDEs, Partial regularity, Complex fluid dynamics, Hydrodynamics of liquid crystal
pacs:
MSC Classification]60H15, 35R60, 60H30, 76D06, 76D03
1 Introduction
Liquid crystal, an anisotropic phase situated between the isotropic liquid phase and solid phase, has drawn considerable attention from researchers from various disciplines. Among all types of liquid crystal states, the nematic phase is particularly interesting for its characteristic that rod-shaped molecules display the orientational order (tending to point in the same direction) without positional order. The Ericksen–Leslie equations (see for instance [22, 23, 61]) serve as a fundamental mathematical model for describing the hydrodynamics of nematic liquid crystal flows, a subject of extensive research by itself [68, 91, 92]. To characterize orientations of nematic liquid crystal molecules, ‘order parameters’ are commonly adopted; the so-called ‘director theory’, which covers the Ericksen–Leslie theory [62] as a special case, employs unit vector field to represent alignment of the molecules.
The Ericksen–Leslie equation system typically combines momentum balance with the evolution of molecular orientation, captured by the Navier–Stokes equations [27, 86, 87] and the heat flow of harmonic maps [66], respectively. In this article we focus on the simplified version of the full Ericksen–Leslie equation system, which was initially studied in a series of works including [64, 65]:
$\left\{\begin{array}[]{l}\,\mathrm{d}{{\bf{u}}}^{\epsilon}-\Delta{{\bf{u}}}^{\epsilon}\,\mathrm{d}t+{{\bf{u}}}^{\epsilon}\cdot\nabla{{\bf{u}}}^{\epsilon}\,\mathrm{d}t+\nabla P^{\epsilon}\,\mathrm{d}t=-\mathop{\rm div}\nolimits(\nabla{{\bf{d}}}^{\epsilon}\odot\nabla{{\bf{d}}}^{\epsilon})\,\mathrm{d}t,\\ \mathop{\rm div}\nolimits{{\bf{u}}}^{\epsilon}=0,\\ \,\mathrm{d}{{\bf{d}}}^{\epsilon}-\Delta{{\bf{d}}}^{\epsilon}\,\mathrm{d}t+{{\bf{u}}}^{\epsilon}\cdot\nabla{{\bf{d}}}^{\epsilon}\,\mathrm{d}t=-f_{\epsilon}({{\bf{d}}}^{\epsilon})\,\mathrm{d}t.\end{array}\right$. (1)
Here with every fixed $\epsilon>0$, the unknown triple $({{\bf{u}}}^{\epsilon},P^{\epsilon},{{\bf{d}}}^{\epsilon})$ represents the fluid velocity field, the pressure function, and the director field which characterizes the mean orientation of molecules, respectively. Moreover
$f_{\epsilon}({{\bf{d}}}^{\epsilon})=\nabla_{{\bf{d}}}F_{\epsilon}({{\bf{d}}}^{\epsilon}),$ (2)
where
$F_{\epsilon}({{\bf{d}}}^{\epsilon})=\frac{1}{\epsilon^{2}}(|{{\bf{d}}}^{\epsilon}|^{2}-1)^{2}.$ (3)
It is evident that (1)1,2 is a forced Stokes system with the so-called Ericksen stress tensor
$(\nabla{{\bf{d}}}^{\epsilon}\odot\nabla{{\bf{d}}}^{\epsilon})_{ij}=\partial_{x_{i}}{{\bf{d}}^{\epsilon}}\cdot\partial_{x_{j}}{{\bf{d}}}^{\epsilon},$
while (1)3 is a transported semi-linear heat flow associated to the Ginzburg–Landau potential function $F_{\epsilon}({{\bf{d}}}^{\epsilon})$.
Subsequently we will only consider (1) with a fixed $\epsilon$; without loss of generality we will fix $\epsilon=1$ and omit the corresponding subscripts and superscripts. Indeed exploration of the singular limit as $\epsilon\to 0$ in (1)3 is an ultimate goal, which leads to the transported heat flow of harmonic maps for the director field ${\bf{d}}$ as follows:
$\mathrm{d}{{\bf{d}}}-\Delta{{\bf{d}}}\,\mathrm{d}t+{{\bf{u}}}\cdot\nabla{{\bf{d}}}\,\mathrm{d}t=|\nabla{{\bf{d}}}|^{2}{{\bf{d}}}\,\mathrm{d}t.$
It is well-established that the critical spatial dimension for heat flow of harmonic maps is two [12, 13]. Consequently, in space dimension two, it has been confirmed that solutions to the (limited) Ericksen–Leslie equation are partially regular [67], and singularities could occur [46, 55]. Moreover, in space dimension two, the singular limit as $\epsilon\to 0$ has been justified through concentration-cancellation compactness [20, 54]. However, in space dimension three, the global well-posedness for the (limited) Ericksen–Leslie equation remains a prominent challenge, with only a handful of partial results available, such as those presented in [36, 37, 38, 39, 40, 69, 89].
While the study of deterministic Ericksen–Leslie equation system has attracted considerable attention, it is worth noting that the deterministic framework could possibly overlook complexities introduced by real-world scenarios, including environmental disturbances [14], measurement uncertainties, and inherent thermal fluctuations [6, 57]. To address these intricacies, a typical way is to introduce a stochastic component into the system. Stochastic forcing in the nematic liquid crystal flows may potentially lead to emergence of singularities, which can cause turbulence in fluid flow [10, 31, 44, 85] and defects in molecular alignments [18, 56, 88]. These phenomena bear substantial implications, impacting our understanding of both theory and the practical applications of liquid crystal materials.
To this end, in this article, we study the stochastically forced Ericksen–Leslie equations (SEL) on the three-dimensional tori $\mathbb{T}^{3}$ driven by an additive white noise $W$ as follows:
$\left\{\begin{array}[]{l}\,\mathrm{d}{{\bf{u}}}-\Delta{{\bf{u}}}\,\mathrm{d}t+{{\bf{u}}}\cdot\nabla{{\bf{u}}}\,\mathrm{d}t+\nabla P\,\mathrm{d}t=-\mathop{\rm div}\nolimits(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}})\,\mathrm{d}t+\,\mathrm{d}W,\\ \mathop{\rm div}\nolimits{{\bf{u}}}=0,\\ \,\mathrm{d}{{\bf{d}}}-\Delta{{\bf{d}}}\,\mathrm{d}t+{{\bf{u}}}\cdot\nabla{{\bf{d}}}\,\mathrm{d}t=-f({{\bf{d}}})\,\mathrm{d}t.\end{array}\right$. (4)
Here we note again that $f({{\bf{d}}})$ ccorresponds to $f_{\epsilon}({{\bf{d}}}^{\epsilon})$ when $\epsilon$ is fixed as 1 (see also Equations (2) and (3)). Our main objectives here are two-fold. Firstly, we introduce the concept of martingale suitable weak solutions to (4) and establish their global existence. Secondly, we demonstrate that under certain relatively reasonable assumptions on $W$ (see Section 2.2), the set of singular points, where solutions become unbounded, cannot have a positive one-dimensional parabolic Hausdorff measure. We remark that the concept of martingale solutions that we formulate here is equivalent to that of statistical solutions, a topic extensively studied in fluid dynamics (see for instance [28, 29, 30]).
It is worthwhile to discuss the physical relevance of our choice of modeling the stochastic nematic liquid crystals dynamics by introducing noise into the fluid velocity, as in (4). For example, dynamic scattering, a well-documented and extensively studied physical phenomenon observed in nematics (see, e.g., [34, 35, 52, 72, 43, 48]), typically arises in experiments where the cellular flow of the liquid intensifies to the point of becoming turbulent. Here the liquid velocity plays a central role in the system dynamics, heavily affecting the alignment of the optical director (see e.g. [18]). As a result, it is reasonable to treat liquid flow as the main driver of the system dynamics and assume velocity fluctuations as the dominant source of randomness. Moreover, there are other electrohydrodynamic instabilities identified in previous studies (see e.g. [3, 76, 70, 71, 73, 50]) that also support the class of SEL models in which liquid flow plays a central role. A classical example is the Williams domain instability. Visually resembling the Rayleigh-Bénard instability in isotropic fluids, it can produce a variety of flow geometries depending on the angle between the wave vector and the director field, including normal rolls, oblique rolls, and more complex structures [41, 42, 77]. Additionally, some recent works on stochastic nematic liquid crystal systems also employ this SEL modeling framework where noise is introduced in the velocity component; see, for example, [33, 75].
The question of whether fluid dynamics equations such as 3D Navier-Stokes equations develop finite-time singularities has been a well-known open problem for a long time; to take into account these singularities, it becomes necessary to consider weak solutions, which live in a larger function space than that of continuous or differentiable functions. However, issues may arise with weak solutions, such as non-uniqueness and non-physical properties (see e.g. [2, 10]), due to insufficient regularity. With this motivation, the concept of suitable weak solutions which satisfy a local energy inequality is introduced (for the first time in [60]). Subsequently, a natural question arises: how regular can these solutions be? Various results have been found for the Navier–Stokes equation, as in [80, 81], where suitable weak solutions are shown to be smooth except for a singular set of parabolic Hausdorff dimension $5/3$. Later on in [11], this result was improved; the dimension of this singular set is demonstrated to be no greater than 1. Since then, numerous works have been carried out to explore partial regularity in the context of general fluid dynamics equations, see e.g. [15, 32, 49]. Notably, in [65], a partial regularity result for the deterministic Ericksen–Leslie equations (4) in space-dimension three was achieved.
In contrast, there are relatively limited findings regarding the partial regularity for stochastic PDEs. To the best of our knowledge, such investigations have primarily focused on the stochastic Navier-Stokes equations; notably, in [26, 79], it was established that in 3D the same partial regulairy result as in [11] holds for martingale suitable weak solutions for almost surely every random path. Building upon the insights of [26, 79], here in this article we establish the same type of partial regularity result for SEL in space dimension three. As far as we know, there has been few reported results about the regularity of the SEL equations in space dimension three, even with $\epsilon$ being fixed ([7, 8]), despite a growing interest in this topic recently. In space dimension two, the well-posedness results concerning the SEL equations have been founded in [17] (with additive noise) and [90] (with multiplicative noise), and for further details see e.g. [9, 84].
The article is organized as follows. In Section 2, we specify notations and stochastic framework needed to set up the SEL system. Then in Section 3 we introduce the notion of suitable weak solutions, drawing inspirations from a split-up scheme which is introduced in [26, 79] in the context of the stochastic Navier–Stokes equation. Specifically, this approach involves splitting solutions to the SEL equations into two components. The first part is governed by linear stochastic Stokes equations, and the second part is subject to the remaining nonlinear coupled system. For the first component of the solution, we directly obtain a desired bound, whenever a the noise term is bounded in a suitable fractional Sobolev space almost surely (Section 3.2). Subsequently over each random path that satisfies this assumption, we define deterministic suitable weak solutions which satisfy two local energy inequalities (Section 3.2). Thus we come up with the notion of a martingale suitable weak solution as a martingale solution each trajectory of which is almost surely a deterministic suitable weak solution. Next in Section 3.3, the main result of the article is stated, namely, the global existence and partial regularity of martingale suitable weak solutions. These two claims are proved in Sections 4 and 5, respectively. For the proof of the global existence (Section 4), it is worth noting that the main challenge is to construct an approximation such that the local energy inequality is preserved after passage to the limit. Here we employ a specific type of approximation by regularization, which was introduced in [60] for the Navier–Stokes equations. In [11], a similar approximation is adopted through a ’retarded mollification’ to establish the existence of suitable weak solutions for the Navier–Stokes equations.
2 Mathematical settings and notations
2.1 Functional spaces
We introduce two fundamental functional spaces that take into account the incompressibility condition (4)2 (cf. [87]):
$H:=\{{\bf{u}}\in L^{2}(\mathbb{T}^{3};\mathbb{R}^{3})|\mathop{\rm div}\nolimits{\bf{u}}=0\},$
and
$V:=\{{\bf{u}}\in H^{1}(\mathbb{T}^{3};\mathbb{R}^{3})|\mathop{\rm div}\nolimits{\bf{u}}=0\}.$
The operator $A:D(A)\subset H\to H$ is defined as $Au=-P_{L}\Delta{\bf{u}}$, where $P_{L}$ is the Helmholtz–Leray projection from $L^{2}(\mathbb{T}^{3};\mathbb{R}^{3})$ onto $H$ and $D(A)=H^{2}(\mathbb{T}^{3};\mathbb{R}^{3})\cap V$. It is well-known that $A^{-1}$ is well-defined from $H$ to $D(A)$. It is also self-adjoint. By the spectral theory of compact self-adjoint operators in a Hilbert space, there exists a sequence of eigenvalues $0<\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{m}\leq\cdots$, $\lambda_{m}\to\infty$ as $m\to\infty$ and an orthonormal basis $\{e_{i}\}_{i=1}^{\infty}$ such that $Ae_{i}=\lambda_{i}e_{i}$. The fractional power $A^{\alpha}$ of $A$, $\alpha\geq 0$, is simply defined by
$A^{\alpha}{\bf{u}}=\sum_{i=1}^{\infty}\lambda_{i}^{\alpha}e_{i}({\bf{u}},e_{i})_{L^{2}(\mathbb{T}^{3})},$
with the domain
$D(A^{\alpha})=\left\{{\bf{u}}\in H\Big|\|{\bf{u}}\|_{D(A^{\alpha})}^{2}=\sum_{i=1}^{\infty}\lambda^{2\alpha}_{i}({\bf{u}},e_{i})_{L^{2}(\mathbb{T}^{3})}^{2}=\|A^{\alpha}{\bf{u}}\|_{L^{2}(\mathbb{T}^{3})}^{2}<\infty\right\}.$
We also note that $D(A^{\alpha})\subset H^{2\alpha}(\mathbb{T}^{3};\mathbb{R}^{3})$. For the definition of $H^{2\alpha}(\mathbb{T}^{3};\mathbb{R}^{3})$, we refer to [78, Definition 1.10].
2.2 Stochastic framework
In order to define the noise term, we first recall some basic concepts and notations of probability and stochastic processes. For more details see for instance [16, 24, 25, 26, 74, 79].
We fix a probability space $(\Omega,\mathcal{F},\mathbb{P})$, where $\Omega$ is a sample space equipped with a $\sigma$-algebra $\mathcal{F}$, and a probability measure $\mathbb{P}$. Then we define a stochastic basis
$\mathfrak{B}:=\left(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P},\left(W(t)\right)_{t\geq 0}\right),$ (5)
which is a filtered probability space with a Brownian motion $\left(W(t)\right)_{t\geq 0}$ adapted to the filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$. In order to avoid unnecessary complications below we may assume that $\{\mathcal{F}_{t}\}_{t\geq 0}$ is complete and right continuous (see [74]). We then assume that
$\left(W(t)\right)_{t\geq 0}$ takes value in $D(A^{\delta})$ for some $\delta>0$ with a covariance operator $\mathcal{L}$. (6)
Here $\mathcal{L}\in\mathcal{B}(D(A^{\delta}))$ and is non-negative and of trace class. Let $\{b_{i}\}_{i\geq 1}$ be the orthonormal basis of $D(A^{\delta})$ consisting of eigenvectors of $\mathcal{L}$ subject to non-negative eigenvalues $\{\gamma_{i}\}_{i\geq 1}$. Then we have the following expansion (cf. [74, Theorem 4.20])
$W(t)=\sum_{i}W_{i}(t)b_{i},$
where the real-valued Brownian motions
$W_{i}(t)=(W(t),b_{i})_{D(A^{\alpha})}$
are independent and their covariances are
$\mathbb{E}[W_{i}(t)W_{i}(s)]=(t\wedge s)\gamma_{i}.$
Moreover, by Kolmogorov’s continuity criterion (see for instance [83]), we have that for every $\beta\in(0,1/2)$, it holds that
$W(t)\in C^{\beta}\left([0,T];D(A^{\delta})\right)\quad\mathbb{P}-a.s..$ (7)
For the purpose of obtaining a key preliminary estimation later on (see Lemma 1 in Section 3.2), we will need to further assume that $\delta\geq\frac{3}{4}$. So this assumption together with (6) yields the assumption that we require on the noise:
$\left(W(t)\right)_{t\geq 0}$ takes value in $D(A^{\delta})$ for some $\delta>3/4$ with a covariance operator $\mathcal{L}$. (8)
3 Deterministic and martingale suitable weak solutions
In order to define suitable weak solutions that are deterministic and are martingale, we follow the approach of [26, 79], and split the SEL into the sum of a linear stochastic Stokes equation and a nonlinear modified Ericksen-Leslie equation.
3.1 A split-up scheme of the SEL
We will assume throughout the paper the following assumptions on the initial data
$({{\bf{u}}}_{0},{{\bf{d}}}_{0})\in H\times H^{1}(\mathbb{T}^{3};\mathbb{R}^{3}),\quad|{{\bf{d}}}_{0}(x)|=1\text{ for all }x\in\mathbb{T}^{3}.$ (9)
As a preparation to introduce notions of deterministic and stochastic suitable weak martingale solutions, we will first split the solutions ${{\bf{u}}}$ and $P$ properly. We introduce new variables ${{\bf{u}}}={{\bf{z}}}+{{\bf{v}}}$, $P={Q}+{\pi}$, where $({{\bf{z}}},{Q})$ solves the linear stochastic Stokes system as follows
$\left\{\begin{array}[]{l}{\,\mathrm{d}}{{\bf{z}}}-\Delta{{\bf{z}}}{\,\mathrm{d}}t+\nabla{Q}{\,\mathrm{d}}t={\,\mathrm{d}}W,\\ \mathop{\rm div}\nolimits{{\bf{z}}}=0,\\ {{\bf{z}}}(0)=0.\end{array}\right$. (10)
The triple $({{\bf{v}}},{\pi},{{\bf{d}}})$ solves the modified Ericksen–Leslie equations as follows
$\begin{cases}&\partial_{t}{{\bf{v}}}-\Delta{{\bf{v}}}+({{\bf{z}}}+{{\bf{v}}})\cdot\nabla({{\bf{z}}}+{{\bf{v}}})+\nabla{\pi}\\ &\quad\quad\quad\quad\quad=-\mathop{\rm div}\nolimits\left(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}|\nabla{{\bf{d}}}|^{2}{\rm I}_{3}-F({{\bf{d}}}){\rm I}_{3}\right),\\ &\mathop{\rm div}\nolimits{{\bf{v}}}=0,\\ &\partial_{t}{{\bf{d}}}-\Delta{{\bf{d}}}+({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{d}}}=-f({{\bf{d}}}),\\ &{{\bf{v}}}(0)={{\bf{u}}}_{0},{{\bf{d}}}(0)={{\bf{d}}}_{0}.\end{cases}$ (11)
Remark 1.
Thanks to the incompressibility condition (11)2, the right-hand side of (11)1 is is equivalent to the deterministic term on the right-hand side of (4)1. Specifically, for the extra term, we observe:
$\mathop{\rm div}\nolimits\left(\frac{1}{2}|\nabla{{\bf{d}}}|^{2}{\rm I}_{3}+F({{\bf{d}}}){\rm I}_{3}\right)=\nabla\left(\frac{1}{2}|\nabla{{\bf{d}}}|^{2}+F({{\bf{d}}})\right),$
which can be treated as part of the pressure gradient. Moreover, this observation, together with direct computations, yields
$\mathop{\rm div}\nolimits\left(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}|\nabla{{\bf{d}}}|^{2}{\rm I}_{3}-F({{\bf{d}}}){\rm I}_{3}\right)=-\nabla{\bf{d}}\cdot(\Delta{\bf{d}}-f({\bf{d}})).$ (12)
As a result, for computational convenience, we shall use the right-hand-side of the above equation to replace the right-hand side of of (11)1.
3.2 Definitions of suitable weak solutions
In order to define what it means for the modified Ericksen-Leslie equation (11) to have a weak solution, we need a preliminary estimation of the solution to the linear stochastic Stokes system (10). We first derive a result that ensures the boundedness of ${{\bf{z}}}(t)$, the stochastic process that solves (10). To achieve this goal, we need to utilize the following result, which is also in [26] (see [26, Lemma 2.2]).
Lemma 1 (An almost-sure bound of the linear stochastic Stokes system).
With the assumption on the noise as in (8), we have
${{\bf{z}}}(\omega)\in L_{\rm loc}^{\infty}(\mathbb{T}^{3}\times[0,\infty))\mbox{ for $\displaystyle\mathbb{P}$- a.s. $\displaystyle\omega$}.$ (13)
The proof of Lemma 1 is provided in Appendix A.
With Lemma 1 in hand, now we are ready to introduce the notions of deterministic and martingale suitable weak solutions to SEL.
Definition 1.
Assume that $T>0$ and fix $\omega\in\Omega$ such that ${\bf{z}}(\omega)$ satisfies the bound specified in (13). Then we say $({{\bf{v}}},{{\bf{d}}})$ is a weak solution to (11) in the sense of distribution on
$Q_{T}:=\mathbb{T}^{3}\times(0,T],$
subject to ${{\bf{v}}}(0)={{\bf{u}}}_{0},{{\bf{d}}}(0)={{\bf{d}}}_{0}$, if for any $\phi_{1}\in C^{\infty}(\mathbb{T}^{3},\mathbb{R}^{3})$, $\mathop{\rm div}\nolimits\phi_{1}=0$, $\phi_{2}\in C^{\infty}(\mathbb{T}^{3},\mathbb{R}^{3})$, and $t>0$, the following weak formulation of (11) is satisfied:
$\int_{\mathbb{T}^{3}\times{\{t\}}}{{\bf{v}}}\cdot\phi_{1}$- $\int_{0}^{t}\int_{\mathbb{T}^{3}}{{\bf{v}}}\cdot\Delta\phi_{1}+({{\bf{z}}}+{{\bf{v}}})\otimes({{\bf{z}}}+{{\bf{v}}}):\nabla\phi_{1}$
$=\int_{\mathbb{T}^{3}}{\bf{u}}_{0}\cdot\phi_{1}+\int_{0}^{t}\int_{\mathbb{T}^{3}}\left(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}|\nabla{{\bf{d}}}|^{2}{\rm I}_{3}-F({{\bf{d}}}){\rm I}_{3}\right):\nabla\phi_{1},$
$\int_{\mathbb{T}^{3}\times{\{t\}}}{{\bf{d}}}\cdot\phi_{2}$- $\int_{0}^{t}\int_{\mathbb{T}^{3}}{{\bf{d}}}\cdot\Delta\phi_{2}+({{\bf{z}}}+{{\bf{v}}})\otimes{\bf{d}}:\nabla\phi_{2}=\int_{\mathbb{T}^{3}}{\bf{d}}_{0}\cdot\phi_{2}-\int_{0}^{t}\int_{\mathbb{T}^{3}}f({\bf{d}})\cdot\phi_{2}.$
Definition 2.
Assume that $T>0$ and fix $\omega\in\Omega$ such that ${\bf{z}}(\omega)$ satisfies the bound specified in (13). Then a suitable weak solution to (4) is a triple $({{\bf{u}}},P,{{\bf{d}}})$ such that if ${{\bf{v}}}={{\bf{u}}}-{{\bf{z}}}$ and ${\pi}=P-{Q}$ where $({{\bf{z}}},{Q})$ is the solution to (10), then
  1. 1.
    $({{\bf{v}}},{{\bf{d}}})$ is weakly continuous with respect to time,
  2. 2.
    ${{\bf{v}}}\in L^{\infty}(0,T;H)\cap L^{2}(0,T;V)$, ${{\bf{d}}}\in L^{\infty}(0,T;H^{1}(\mathbb{T}^{3};\mathbb{R}^{3}))\cap L^{2}(0,T;H^{1}(\mathbb{T}^{3};\mathbb{R}^{3}))$ and ${\pi}\in L^{5/3}(0,T;L^{5/3}(\mathbb{T}^{3};\mathbb{R})),$
  3. 3.
    $({{\bf{v}}},{{\bf{d}}})$ is a weak solution to (11) in the sense of distribution on $Q_{T}$ as specified in Def. 1,
  4. 4.
    for all $t\leq T$ and almost every $s<t$, we have
    $\int_{\mathbb{T}^{3}\times\{t\}}\left(\frac{1}{2}|{{\bf{v}}}|^{2}+\frac{1}{2}|\nabla{{\bf{d}}}|^{2}+F({{\bf{d}}})\right)+\int_{s}^{t}\int_{\mathbb{T}^{3}}|\nabla{{\bf{v}}}|^{2}+|\Delta{{\bf{d}}}-f({{\bf{d}}})|^{2}$ (14)
    $\leq$ $\int_{\mathbb{T}^{3}\times\{s\}}\left(\frac{1}{2}|{{\bf{v}}}|^{2}+\frac{1}{2}|\nabla{{\bf{d}}}|^{2}+F({{\bf{d}}})\right)$
    $+\int_{s}^{t}\int_{\mathbb{T}^{3}}{{\bf{z}}}\cdot[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{v}}}]+({{\bf{z}}}\cdot\nabla){{\bf{d}}}\cdot(\Delta{{\bf{d}}}-f({{\bf{d}}})).$
  5. 5.
    for any $\varphi\in C_{0}^{\infty}(Q_{T})$ (we note that the subscript “0” refers to compact support in the corresponding domain), $\varphi\geq 0$, we have
    $\int_{\mathbb{T}^{3}\times\{T\}}\left(\frac{|{{\bf{v}}}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}|^{2}}{2}+F({{\bf{d}}})\right)+\int_{Q_{T}}(|\nabla{{\bf{v}}}|^{2}+|\Delta{{\bf{d}}}|^{2}+|f({{\bf{d}}})|^{2})\varphi$ (15)
    $\leq$ $\int_{Q_{T}}\left(\frac{|{{\bf{v}}}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}|^{2}}{2}+F({{\bf{d}}})\right)\partial_{t}\varphi+\int_{Q_{T}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{\bf{v}}]\cdot{\bf{z}}\varphi$
    $+\int_{Q_{T}}\left(\frac{|{{\bf{v}}}|^{2}}{2}+{\bf{v}}\cdot{\bf{z}}\right)[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla\varphi]+{\pi}{{\bf{v}}}\cdot\nabla\varphi+\left(\frac{|{{\bf{v}}}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}|^{2}}{2}\right)\Delta\varphi$
    $+\int_{Q_{T}}\left(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}|\nabla{\bf{d}}|^{2}{\rm I}_{3}\right):\nabla^{2}\varphi$
    $+\int_{Q_{T}}({{\bf{z}}}\cdot\nabla{{\bf{d}}})\cdot(\Delta{{\bf{d}}}-f({{\bf{d}}}))\varphi+[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{d}}}]\cdot(\nabla\varphi\cdot\nabla){{\bf{d}}}$
    $-\int_{Q_{T}}f({{\bf{d}}})\cdot(\nabla\varphi\cdot\nabla){{\bf{d}}}-2\int_{Q_{T}}\nabla f({{\bf{d}}}):\nabla{{\bf{d}}}\varphi.$
Definition 3.
[…]
$\widetilde{\mathfrak{B}}=(\widetilde{\Omega},\widetilde{\mathcal{F}},\{\widetilde{\mathcal{F}}_{t}\}_{t\geq 0},\widetilde{\mathbb{P}},(\widetilde{W}(t))_{t\geq 0}),$ (16)
and $(\widetilde{W}(t))_{t\geq 0}$ is a Brownian motion adapted to the filtration $\{\widetilde{\mathcal{F}}_{t}\}_{t\geq 0}$ with values in $D(A^{\delta})$ and covariance operator $\mathcal{L}$, such that
$(\widetilde{{\bf{u}}}(\cdot),\widetilde{P}(\cdot),\widetilde{{\bf{d}}}(\cdot)):\widetilde{\Omega}\times[0,\infty)\rightarrow H\times\mathbb{R}\times H^{1}(\mathbb{T}^{3};\mathbb{R}^{3})$
is a $\widetilde{\mathcal{F}}_{t}$ adapted process, and
$\widetilde{\omega}\in\widetilde{\Omega}\to(\widetilde{{\bf{u}}}(\widetilde{\omega}),\widetilde{P}(\widetilde{\omega}),\widetilde{{\bf{d}}}(\widetilde{\omega}))\in L^{2}(0,T;H)\times L^{5/3}(Q_{T})\times L^{2}(0,T;H^{1}(\mathbb{T}^{3};\mathbb{R}^{3}))$
is a measurable mapping, and there exists a set $\widetilde{\Omega}_{0}\subset\widetilde{\Omega}$ of full probability such that on this set the triple $)\displaystyle(\widetilde{{\bf{u}}}(\cdot,\widetilde{\omega}),\widetilde{P}(\cdot,\widetilde{\omega}),\widetilde{{\bf{d}}}(\cdot,\widetilde{\omega}$)) is a suitable weak solution in the sense of Definition 2.
3.3 Main results: existence and partial regularity
To state the main result of our article we introduce the notion of singular points for a suitable weak solution.
Definition 4.
We call a point $z=(x,t)\in\mathbb{T}^{3}\times(0,\infty)$ regular if there exists a neighborhood of $z$ where $({{\bf{v}}},\nabla{{\bf{d}}})$ is essentially bounded. The other points will be called singular. The set of singular points will be denoted by $\Sigma$.
Our main result is as follows.
Theorem 2.
With the assumptions (9), suppose furthermore that the Brownian motion $\left(W(t)\right)_{t\geq 0}$ takes value in $D(A^{\frac{1}{4}+\delta})$ for some $\delta>0$. Then there exists a martingale suitable weak solution to (4) globally in time such that with full probability,
$\mathcal{P}^{1}(\Sigma)=0,$
where $\mathcal{P}^{k}$, $0\leq k\leq 4$, denotes the $k$-dimensional Hausdorff measure on $\mathbb{T}^{3}\times\mathbb{R}_{+}$ with respect to the parabolic distance $\delta_{p}((x,t),(y,s)):=\max\{|x-y|,\sqrt{|t-s|}\},\forall(x,t),(y,s)\in\mathbb{T}^{3}\times\mathbb{R}_{+}.$
4 Existence of martingale suitable weak solutions
In this section, we will prove the first part of the conclusion of Theorem 2, that is, existence of martingale suitable weak solutions to (4) in the sense of Definition 3.
4.1 Energy-inequality-preserving approximation
We begin the proof of existence of martingale suitable weak solutions by introducing a system that converges to the SEL while preserving relevant energy inequalities.
Given ${{\bf{z}}}(\omega)$ as in Lemma 1, we introduce the following approximating equations to (11):
$\begin{cases}&\partial_{t}{{\bf{v}}}^{\sigma}-\Delta{{\bf{v}}}^{\sigma}+({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\cdot\nabla({{\bf{z}}}+{{\bf{v}}}^{\sigma})+\nabla{\pi}^{\sigma}\\ &=-\nabla\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})),\\ &\nabla\cdot{{\bf{v}}}^{\sigma}=0,\\ &\partial_{t}{{\bf{d}}}^{\sigma}-({{\bf{z}}}+{{\bf{v}}}^{\sigma})\cdot\nabla\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]=\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma}),\end{cases}$ (17)
where $\sigma>0$ is the approximation parameter and $\Phi_{\sigma}[f]$ denotes the smooth mollification of a function $f$. More precisely, suppose that $\psi\in C_{0}^{\infty}(\mathbb{R}^{3})$ is a standard mollifier supported in $B_{1}(0)$, and let $\psi_{\sigma}:=\sigma^{-3}\psi(x/\sigma)$. Then we define $\Phi_{\sigma}[f]$ via the following convolution
$\Phi_{\sigma}[f]=\int_{\mathbb{R}^{3}}\psi_{\sigma}(x-y)f(y)\,\mathrm{d}y,$
where we extend $f$ periodically to the whole $\mathbb{R}^{3}$. It is easy to show that
$\|\Phi_{\sigma}[f]\|_{L^{p}(\mathbb{T}^{3})}\leq\|f\|_{L^{p}(\mathbb{T}^{3})},\qquad\lim_{\sigma\to 0}\|\Phi_{\sigma}[f]-f\|_{L^{p}(\mathbb{T}^{3})}=0.$
As the key feature of this type of approximations is that global and local energy inequalities are preserved, it is also applicable to more complicated hydrodynamic equations (see, for instance, [19, 21]). Moreover, we would like to point out that, for the Ericksen stress tensor term, we have employed the same observation as discussed in Remark 1.
We note that (17) has better regularity in nonlinear terms, and one could obtain the existence of solutions globally strong in the PDE sense via a similar argument using the Galerkin method and Banach fixed point theorem as in [26] and [79]. More precisely, for each fixed $\sigma$, we have
$\displaystyle\begin{cases}&{{\bf{v}}}^{\sigma}\in L^{2}(0,T;H^{2}(\mathbb{T}^{3}))\cap L^{\infty}(0,T;V),\quad\partial_{t}{{\bf{v}}}^{\sigma}\in L^{2}(0,T;H),\\ &{{{\bf{d}}}}^{\sigma}\in L^{2}(0,T;H^{2}(\mathbb{T}^{3}))\cap L^{\infty}(0,T;V),\quad\partial_{t}{{{\bf{d}}}}^{\sigma}\in L^{2}(0,T;H),\\ &{\pi}^{\sigma}\in L^{2}(0,T;H^{1}(\mathbb{T}^{3})).\end{cases}$ (18)
Moreover, we note that subtracting any constant from $\pi^{\sigma}$ yields again a solution to (17). So by subtracting ${\pi}^{\sigma}$ with its mean, $\int_{\mathbb{T}^{3}}{\pi}^{\sigma}$, we will obtain a function with zero mean; we will choose this function as the solution. Thus, without loss of generality, from now on we will assume that
$\int_{\mathbb{T}^{3}}{\pi}^{\sigma}=0.$ (19)
4.1.1 Pathwise estimates independent of the approximation parameter
We can derive the following estimates on $({{\bf{v}}}^{\sigma},{\pi}^{\sigma},{{\bf{d}}}^{\sigma})$ independent of $\sigma$, with a fixed arbitrary $\omega\in\Omega$ such that ${\bf{z}}(\omega)$ satisfies the bound specified in (13); these estimates will be used to achieve the compactness argument later on (see Section 4.2.1).
Lemma 3.
[Energy estimates independent of $\sigma$] Fix $\omega\in\Omega$ such that ${{\bf{z}}}(\omega)\in L_{\rm loc}^{\infty}(\mathbb{T}^{3}\times[0,\infty))$. Consider the triple $({{\bf{v}}}^{\sigma}(\omega),{\pi}^{\sigma}(\omega),{{\bf{d}}}^{\sigma}(\omega))$, a classical strong (in the PDE sense) solution to (17), then we have the following estimates:
$\begin{cases}&{\|{{\bf{v}}}^{\sigma}\|_{L^{\infty}(0,T;H)}^{2}+\|{{\bf{d}}}^{\sigma}\|_{L^{\infty}(0,T;H^{1}(\mathbb{T}^{3}))}^{2}+\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(Q_{T})}^{2}+\|\nabla^{2}{{\bf{d}}}^{\sigma}\|_{L^{2}(Q_{T})}^{2}}\\ &\leq\Psi(T),\\ &\|{\pi}^{\sigma}\|_{L^{\frac{5}{3}}(Q_{T})}\leq C\Psi(T)+{C\|{{\bf{z}}}\|_{L^{\infty}(Q_{T})}^{2}},\\ &\|\partial_{t}{{\bf{v}}}^{\sigma}\|_{L^{2}(0,T;D(A^{-1}))+L^{2}(0,T;H^{-1}(\mathbb{T}^{3}))+L^{\frac{5}{4}}(Q_{T})}\leq C(\sqrt{\Psi(T)}+\Psi(T)),\\ &\|\partial_{t}{\bf{d}}^{\sigma}\|_{L^{\frac{5}{3}}(Q_{T})}\leq C(\Psi(T)+T^{\frac{1}{10}}\sqrt{\Psi(T)}),\\ &|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})|^{2}_{L^{2}(Q_{T})}\leq C\Psi(T),\end{cases}$ (20)
where $C$ is a constant that is independent of $\sigma$, and
$\Psi(T)=$ $C\left(\|{\bf{u}}_{0}\|_{H}^{2}+\|{\bf{d}}_{0}\|_{H^{1}(\mathbb{T}^{3})}^{2}+\|{{\bf{z}}}(t)\|_{L^{4}(Q_{T})}^{4}\right)e^{C\int_{0}^{T}{\frac{1}{2}}+\|{\bf{z}}(t)\|_{L^{4}(\mathbb{T}^{3})}^{4}\,\mathrm{d}t}.$ (21)
Proof.
The proof of this lemma is based on deriving suitable energy estimates. Due to the length of the derivation of the global energy equality, we have placed it in Appendix B. In particular, all of the estimates stated above will be derived from equation (86) in Corollary 14.
Firstly, we deal with the two terms on the right-hand-side (RHS) of (86) one by one. For the first term we have the following estimation:
$\displaystyle\left|\int_{\mathbb{T}^{3}}[({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\cdot\nabla{{\bf{v}}}^{\sigma}]\cdot{{\bf{z}}}{\,\mathrm{d}x}\right|$
$\displaystyle\leq$ $\displaystyle\int_{\mathbb{T}^{3}}|{{\bf{z}}}||\nabla{{\bf{v}}}^{\sigma}||{{\bf{z}}}|+|\Phi_{\sigma}[{{\bf{v}}}^{\sigma}]||\nabla{{\bf{v}}}^{\sigma}||{{\bf{z}}}|$
$\displaystyle\leq$ $\displaystyle\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}+\|\Phi_{\sigma}[{{\bf{v}}}^{\sigma}]\|_{L^{4}(\mathbb{T}^{3})}\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}$
$\displaystyle\leq$ $\displaystyle\frac{1}{4}\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\left(\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}+\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{2}\|{{\bf{v}}}^{\sigma}\|_{L^{4}(\mathbb{T}^{3})}^{2}\right)$
$\displaystyle\leq$ $\displaystyle\frac{1}{4}\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\left(\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}+\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{2}\|{{\bf{v}}}^{\sigma}\|_{{H^{1}(\mathbb{T}^{3})}}^{\frac{3}{2}}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{\frac{1}{2}}\right)$
$\displaystyle\leq$ $\displaystyle\frac{1}{4}\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\left(\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}+\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{2}\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{\frac{3}{2}}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{\frac{1}{2}}+{\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{2}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}}\right)$
$\displaystyle\leq$ $\displaystyle\frac{1}{2}\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\left(\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}+\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{4})}^{4}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+{\left(\dfrac{1}{2}\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}+\dfrac{1}{2}\right)\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}}\right)$
$\displaystyle\leq$ $\displaystyle\frac{1}{2}\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\left(\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}+{\dfrac{3}{2}}\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{4})}^{4}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+{\dfrac{1}{2}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}}\right),$
where $C$ is a constant [^1] Note that C may be different at each occurrence below but all of them are independent of $\sigma$. which is independent of $\sigma$. For the second term, by the Hölder inequality and Young inequality, we obtain the following estimation:
$\displaystyle\left|\int_{\mathbb{T}^{3}}{{\bf{z}}}\cdot\nabla{\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]}\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})){\,\mathrm{d}x}\right|$
$\displaystyle\leq$ $\displaystyle\frac{1}{8}\|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{2}\|\nabla{{\bf{d}}}^{\sigma}\|_{L^{4}(\mathbb{T}^{3})}^{2},$
$\displaystyle\leq$ $\displaystyle\frac{1}{8}\|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{2}\|\nabla^{2}{{\bf{d}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{\frac{3}{2}}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{\frac{1}{2}}$
$\displaystyle\leq$ $\displaystyle\frac{1}{8}\|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})\|_{L^{2}(\mathbb{T}^{3})}^{2}+\frac{1}{4}\|\nabla^{2}{{\bf{d}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}.$
Combining these two estimations, together with (86), we obtain
$\frac{\,\mathrm{d}}{{\,\mathrm{d}t}}\int_{\mathbb{T}^{3}}\frac{1}{2}(|{{\bf{v}}}^{\sigma}|^{2}+|\nabla{{\bf{d}}}^{\sigma}|^{2})+F({{\bf{d}}}^{\sigma}){\,\mathrm{d}}x+\dfrac{1}{2}\int_{\mathbb{T}^{3}}|\nabla{{\bf{v}}}^{\sigma}|^{2}{\,\mathrm{d}}x$ (22)
$+\dfrac{7}{8}\int_{\mathbb{T}^{3}}|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})|^{2}{\,\mathrm{d}}x$
$\leq$ $C\left(\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}+{\dfrac{3}{2}}\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{4})}^{4}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+{\dfrac{1}{2}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}}\right)$
$+\frac{1}{4}\|\nabla^{2}{{\bf{d}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}$
$\leq$ $C\left(\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}+{\dfrac{1}{2}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}}\right)+\frac{1}{4}\|\nabla^{2}{{\bf{d}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}.$
Here we note again that the constant C may be different at each occurrence. Now we deal with the last term on the left-hand side (LHS) of (22) as follows
$\displaystyle\int_{\mathbb{T}^{3}}|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})|^{2}{\,\mathrm{d}x}$
$\displaystyle=$ $\displaystyle\int_{\mathbb{T}^{3}}[|\Delta{{\bf{d}}}^{\sigma}|^{2}+|f({{\bf{d}}}^{\sigma})|^{2}-2\Delta{{\bf{d}}}^{\sigma}\cdot f({{\bf{d}}}^{\sigma})]{\,\mathrm{d}x}$
$\displaystyle=$ $\displaystyle\int_{\mathbb{T}^{3}}[|\Delta{{\bf{d}}}^{\sigma}|^{2}+|f({{\bf{d}}}^{\sigma})|^{2}+2\nabla{{\bf{d}}}^{\sigma}:\nabla f({{\bf{d}}}^{\sigma})]{\,\mathrm{d}x}$
$\displaystyle=$ $\displaystyle\int_{\mathbb{T}^{3}}[|\Delta{{\bf{d}}}^{\sigma}|^{2}+|f({{\bf{d}}}^{\sigma})|^{2}-2|\nabla{{\bf{d}}}^{\sigma}|^{2}+2|\nabla{{\bf{d}}}^{\sigma}|^{2}|{{\bf{d}}}^{\sigma}|^{2}+4|(\nabla{{\bf{d}}}^{\sigma})^{\top}{{\bf{d}}}^{\sigma}|^{2}]{\,\mathrm{d}x}.$
By slightly rearranging the terms on the RHS, we arrive at
$\int_{\mathbb{T}^{3}}|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})|^{2}{\,\mathrm{d}x}$ (23)
$=$ $\int_{\mathbb{T}^{3}}[|\Delta{{\bf{d}}}^{\sigma}|^{2}+|f({{\bf{d}}}^{\sigma})|^{2}+2|\nabla{{\bf{d}}}^{\sigma}|^{2}|{{\bf{d}}}^{\sigma}|^{2}+4|(\nabla{{\bf{d}}}^{\sigma})^{\top}{{\bf{d}}}^{\sigma}|^{2}]{\,\mathrm{d}x}-2\int_{\mathbb{T}^{3}}|\nabla{{\bf{d}}}^{\sigma}|^{2}{\,\mathrm{d}x}.$
Combining (23) and (22), we obtain, after some slight rearrangements:
$\frac{\,\mathrm{d}}{{\,\mathrm{d}t}}\int_{\mathbb{T}^{3}}\frac{1}{2}(|{{\bf{v}}}^{\sigma}|^{2}+|\nabla{{\bf{d}}}^{\sigma}|^{2})+F({{\bf{d}}}^{\sigma}){\,\mathrm{d}}x+|\nabla{{\bf{v}}}^{\sigma}|^{2}_{L^{2}(\mathbb{T}^{3})}+\dfrac{1}{8}\int_{\mathbb{T}^{3}}|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})|^{2}{\,\mathrm{d}x}$ (24)
$+\dfrac{6}{8}\int_{\mathbb{T}^{3}}[|\Delta{{\bf{d}}}^{\sigma}|^{2}+|f({{\bf{d}}}^{\sigma})|^{2}+2|\nabla{{\bf{d}}}^{\sigma}|^{2}|{{\bf{d}}}^{\sigma}|^{2}+4|(\nabla{{\bf{d}}}^{\sigma})^{\top}{{\bf{d}}}^{\sigma}|^{2}]{\,\mathrm{d}x}$
$\leq C\left(\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}+{\dfrac{1}{2}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}}\right)+\frac{1}{4}\|\nabla^{2}{{\bf{d}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}$
$+\dfrac{6}{4}\int_{\mathbb{T}^{3}}|\nabla{{\bf{d}}}^{\sigma}|^{2}{\,\mathrm{d}x}.$
By employing techniques including an application of the Gronwall inequality to (24) (for a detailed derivation, we refer the reader to Section C of the Appendix), we obtain
$\sup_{0<t<T}\left(\|{{\bf{v}}}^{\sigma}(t)\|_{H}^{2}+\|{{\bf{d}}}^{\sigma}(t)\|_{H^{1}(\mathbb{T}^{3})}^{2}\right)$ (25)
$+\int_{0}^{T}\|\nabla{{\bf{v}}}^{\sigma}(t)\|_{L^{2}(\mathbb{T}^{3})}^{2}+\|\nabla^{2}{{\bf{d}}}^{\sigma}(t)\|_{L^{2}(\mathbb{T}^{3})}^{2}{\,\mathrm{d}t}$
$\leq$ $C\left(\|{\bf{u}}_{0}\|_{H}^{2}+\|{\bf{d}}_{0}\|_{H^{1}(\mathbb{T}^{3})}^{2}+\int_{0}^{T}\|{{\bf{z}}}(t)\|_{L^{4}(\mathbb{T}^{4})}^{4}\,\mathrm{d}t\right)e^{C\int_{0}^{T}{\frac{1}{2}}+\|{\bf{z}}(t)\|_{L^{4}(\mathbb{T}^{3})}^{4}\,\mathrm{d}t}$
$=$ $C\left(\|{\bf{u}}_{0}\|_{H}^{2}+\|{\bf{d}}_{0}\|_{H^{1}(\mathbb{T}^{3})}^{2}+\|{{\bf{z}}}(t)\|_{L^{4}(Q_{T})}^{4}\right)e^{C\int_{0}^{T}{\frac{1}{2}}+\|{\bf{z}}(t)\|_{L^{4}(\mathbb{T}^{3})}^{4}\,\mathrm{d}t}$
$=:$ $\Psi(T).$
Then from (25) we infer (21), and also we obtain
$\sup_{0<t<T}\left(\|{{\bf{v}}}^{\sigma}(t)\|_{H}^{2}+\|{{\bf{d}}}^{\sigma}(t)\|_{H^{1}(\mathbb{T}^{3})}^{2}\right)+\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(Q_{T})}^{2}+\|\nabla^{2}{{\bf{d}}}^{\sigma}\|_{L^{2}(Q_{T})}^{2}\leq\Psi(T),$
which implies (20)1. Also from (24) and (25) we obtain (20)5.
Next, we aim to derive (20)2, a uniform bound for ${\pi}^{\sigma}$ in $\sigma$. Taking the divergence of (17)1 we obtain
$\Delta{\pi}^{\sigma}=\mathop{\rm div}\nolimits^{2}[({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\otimes({{\bf{z}}}+{{\bf{v}}}^{\sigma})]+\mathop{\rm div}\nolimits(\nabla\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma}))).$ (26)
Applying the elliptic theory as in Lemma 5.1 of [78], and utilizing Hölder’s inequality we obtain:
$\|{\pi}^{\sigma}\|_{L^{\frac{5}{3}}(Q_{T})}\leq$ $C\|({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\otimes({{\bf{z}}}+{{\bf{v}}}^{\sigma})\|_{L^{\frac{5}{3}}(Q_{T})}$ (27)
$+\|\nabla(\Phi_{\sigma}[{{\bf{d}}}^{\sigma}])\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma}))\|_{L^{\frac{5}{3}}(0,T;W^{-1,\frac{5}{3}}(\mathbb{T}^{3}))}$
[…]
$+\|\nabla(\Phi_{\sigma}[{{\bf{d}}}^{\sigma}])\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma}))\|_{L^{\frac{5}{3}}(0,T;W^{-1,\frac{5}{3}}(\mathbb{T}^{3}))}.$
Now we first deal with the first term on the RHS. We easily observe that
$\|{{\bf{z}}}\|_{L^{\frac{10}{3}}(Q_{T})}\leq\|{{\bf{z}}}\|_{L^{\infty}(Q_{T})}.$ (28)
By the Riesz–Thorin theorem and the Sobolev embedding theorem in 3D we observe that
$\|{{\bf{v}}}^{\sigma}\|_{L^{\frac{10}{3}}(Q_{T})}$ $\leq C\|{{\bf{v}}}^{\sigma}\|_{L^{2}(0,T;L^{6}(\mathbb{T}^{3}))}^{\frac{2}{5}}\|{{\bf{v}}}^{\sigma}\|_{L^{\infty}(0,T;L^{2}(\mathbb{T}^{3}))}^{\frac{3}{5}}$ (29)
$\leq C\|{{\bf{v}}}^{\sigma}\|_{L^{2}(0,T;H^{1}(\mathbb{T}^{3}))}^{\frac{2}{5}}\|{{\bf{v}}}^{\sigma}\|_{L^{\infty}(0,T;L^{2}(\mathbb{T}^{3}))}^{\frac{3}{5}}$
Then this together with (28) and (27) imply that
$\|{\pi}^{\sigma}\|_{L^{\frac{5}{3}}(Q_{T})}\leq$ $C\left(\|{{\bf{z}}}\|_{L^{\infty}(Q_{T})}^{2}+C\|{{\bf{v}}}^{\sigma}\|_{L^{2}(0,T;H^{1}(\mathbb{T}^{3}))}^{\frac{4}{5}}\|{{\bf{v}}}^{\sigma}\|_{L^{\infty}(0,T;L^{2}(\mathbb{T}^{3}))}^{\frac{6}{5}}\right)$ (30)
$+\|\nabla(\Phi_{\sigma}[{{\bf{d}}}^{\sigma}])\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma}))\|_{L^{\frac{5}{3}}(0,T;W^{-1,\frac{5}{3}}(\mathbb{T}^{3}))}.$
Next we deal with the second term on the RHS of (30). We first observe that in 3D, the space $L^{\frac{15}{14}}(\mathbb{T}^{3})$ is embedded in $W^{-1,\frac{5}{3}}(\mathbb{T}^{3})$ thanks to the Sobolev inequalities , so applying this observation to the second term on the RHS we obtain
$\|{\pi}^{\sigma}\|_{L^{\frac{5}{3}}(Q_{T})}\leq$ $C\left(\|{{\bf{z}}}\|_{L^{\infty}(Q_{T})}^{2}+C\|{{\bf{v}}}^{\sigma}\|_{L^{2}(0,T;H^{1}(\mathbb{T}^{3}))}^{\frac{4}{5}}\|{{\bf{v}}}^{\sigma}\|_{L^{\infty}(0,T;L^{2}(\mathbb{T}^{3}))}^{\frac{6}{5}}\right)$ (31)
$+\|\nabla(\Phi_{\sigma}[{{\bf{d}}}^{\sigma}])\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma}))\|_{L^{\frac{5}{3}}(0,T;L^{\frac{15}{14}}(\mathbb{T}^{3}))}.$
Now applying Hölder’s inequality to the second term on the RHS we obtain
$\|{\pi}^{\sigma}\|_{L^{\frac{5}{3}}(Q_{T})}\leq$ $C\left(\|{{\bf{z}}}\|_{L^{\infty}(Q_{T})}^{2}+C\|{{\bf{v}}}^{\sigma}\|_{L^{2}(0,T;H^{1}(\mathbb{T}^{3}))}^{\frac{4}{5}}\|{{\bf{v}}}^{\sigma}\|_{L^{\infty}(0,T;L^{2}(\mathbb{T}^{3}))}^{\frac{6}{5}}\right)$ (32)
$+C\|\nabla{{\bf{d}}}^{\sigma}\|^{2}_{L^{10}(0,T;L^{\frac{30}{13}}(\mathbb{T}^{3})}\|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})\|_{L^{2}(Q_{T})}.$
Now applying Riesz–Thorin theorem to the second term on the RHS of (32), and utilizing the Sobolev embedding theorem in 3D, we observe that
$\|\nabla{{\bf{d}}}^{\sigma}\|_{L^{10}(0,T;L^{\frac{30}{13}}(\mathbb{T}^{3}))}\leq C\|\nabla{{\bf{d}}}^{\sigma}\|_{L^{\infty}(0,T;L^{2}(\mathbb{T}^{3}))}^{\frac{1}{5}}\times\|\nabla{{\bf{d}}}^{\sigma}\|_{L^{2}(0,T;L^{6}(\mathbb{T}^{3}))}^{\frac{4}{5}}$ (33)
$\leq C\|\nabla{{\bf{d}}}^{\sigma}\|_{L^{\infty}(0,T;L^{2}(\mathbb{T}^{3}))}^{\frac{1}{5}}\times\|\nabla{{\bf{d}}}^{\sigma}\|_{L^{2}(0,T;H^{1}(\mathbb{T}^{3}))}^{\frac{4}{5}}$
This together with (32) implies that
$\|{\pi}^{\sigma}\|_{L^{\frac{5}{3}}(Q_{T})}$ (34)
$\leq$ $C\left(\|{{\bf{z}}}\|_{L^{\infty}(Q_{T})}^{2}+C\|{{\bf{v}}}^{\sigma}\|_{L^{2}(0,T;H^{1}(\mathbb{T}^{3}))}^{\frac{4}{5}}\|{{\bf{v}}}^{\sigma}\|_{L^{\infty}(0,T;L^{2}(\mathbb{T}^{3}))}^{\frac{6}{5}}\right)$
$+C\|\nabla{{\bf{d}}}^{\sigma}\|_{L^{\infty}(0,T;L^{2}(\mathbb{T}^{3}))}^{\frac{2}{5}}\|\nabla{{\bf{d}}}^{\sigma}\|_{L^{2}(0,T;H^{1}(\mathbb{T}^{3}))}^{\frac{8}{5}}\|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})\|_{L^{2}(Q_{T})}$
Finally from (34), (20)1, and (20)5 we obtain
$\|{\pi}^{\sigma}\|_{L^{\frac{5}{3}}(Q_{T})}\leq C\Psi(T)+{C\|{{\bf{z}}}\|_{L^{\infty}(Q_{T})}^{2}},$ (35)
which is (20)4.
Next, to prove (20)3, we can utilize (17)1 to show that
$\|\partial_{t}{{\bf{v}}}^{\sigma}\|_{L^{2}(0,T;D(A^{-1}))+L^{2}(0,T;H^{-1}(\mathbb{T}^{3}))+L^{\frac{5}{4}}(Q_{T})}$ (36)
$\leq$ $C(\|{{\bf{v}}}^{\sigma}\|_{L^{2}(Q_{T})}+\|{\bf{z}}\|_{L^{4}(Q_{T})}^{2}+\|{{\bf{v}}}^{\sigma}\|_{L^{4}(Q_{T})}^{2}$
$+\|\nabla{\bf{d}}^{\sigma}\|_{L^{\frac{10}{3}}(Q_{T})}\|\Delta{\bf{d}}^{\sigma}-f({{\bf{d}}}^{\sigma})\|_{L^{2}(Q_{T})})$
$\leq$ $C(\sqrt{\Psi(T)}+\Psi(T)).$
Finally, to prove (20)4, we can utilize (17)3 to derive that
$\|\partial_{t}{\bf{d}}^{\sigma}\|_{L^{\frac{5}{3}}(Q_{T})}$ (37)
$\leq$ $C\|{{\bf{z}}}\|_{L^{\frac{10}{3}}(Q_{T})}\|\nabla{\bf{d}}^{\sigma}\|_{L^{\frac{10}{3}}(Q_{T})}+C\|{{\bf{v}}}^{\sigma}\|_{L^{\frac{10}{3}}(Q_{T})}\|\nabla{\bf{d}}^{\sigma}\|_{L^{\frac{10}{3}}(Q_{T})}$
$+CT^{\frac{1}{10}}\|\Delta{\bf{d}}^{\sigma}-f({{\bf{d}}}^{\sigma})\|_{L^{2}(Q_{T})}$
$\leq$ $C(\Psi(T)+CT^{\frac{1}{10}}\sqrt{\Psi(T)}).$
4.2 Approaching the existence of martingale suitable solutions
Based on the pathwise estimates derived above, we can infer a compactness argument from tightness properties of the approximation sequence. Next we will use the Skorokhod embedding theorem (see Theorem 2.4 in [16]) to construct an approximation sequence with a possible shift of the stochastic basis. It is then necessary to verify that the key energy equality (81) is preserved by the constructed sequence. Finally passing to the limit in (17) we obtain the global existence of martingale suitable weak solutions satisfying the local energy inequalities as desired.
4.2.1 Compactness of the approximation sequence
Let $\nu^{\sigma}$ be the law of the random variable $U_{\sigma}:=({{\bf{v}}}^{\sigma},{\pi}^{\sigma},{{\bf{d}}}^{\sigma},{{\bf{z}}},W)$ with values in $\mathcal{E}$:
$\mathcal{E}=L^{2}(0,T;H)\times L^{\frac{5}{3}}(Q_{T})\times L^{2}(0,T;H^{1}(\mathbb{T}^{3};\mathbb{R}^{3}))\times C([0,T];H)\times C_{0}([0,T];H).$
Let $\mu^{\sigma}$ be the projection of $\nu^{\sigma}$ in the variables $({{\bf{v}}}^{\sigma},{{\bf{d}}}^{\sigma})$, that is, the law of $({{\bf{v}}}^{\sigma},{{\bf{d}}}^{\sigma})$. The key step to establish the probabilistic convergence of this approximation scheme is to show that the family of measures $\mu^{\sigma}$ is tight. Namely, for each $\varepsilon>0$, there exists a compact set $K_{\varepsilon}$ in $\mathcal{E}$ such that
$\mu^{\sigma}(K_{\varepsilon})\geq 1-\varepsilon,\text{ for }0<\sigma<1.$ (38)
We take
$K_{R}$ $=\Big\{({{\bf{v}}},{\bf{d}})\Big|\|{{\bf{v}}}\|_{L^{\infty}(0,T;H)\cap L^{2}(0,T;V)}+\|\partial_{t}{{\bf{v}}}\|_{L^{2}(0,T;D(A^{-1}))+L^{2}(0,T;H^{-1}(\mathbb{T}^{3}))+L^{\frac{5}{4}}(Q_{T})}$
$+\|{{\bf{d}}}\|_{L^{\infty}(0,T;H^{1}(\mathbb{T}^{3}))\cap L^{2}(0,T;H^{2})\cap H^{1}(0,T;L^{2}(\mathbb{T}^{3}))}+\|\partial_{t}{\bf{d}}^{\sigma}\|_{L^{\frac{5}{3}}(Q_{T})}\leq R\Big\}.$
By the Aubin-Lions lemma, we conclude that $K_{R}$ is compact in $L^{p_{1}}(Q_{T})\times L^{p_{2}}(Q_{T})$ with $1<p_{1}<\frac{10}{3}$, and $1<p_{2}<10$. Moreover, by the estimates (78), (25), (36) and (37), we can follow the same argument as in [25, (381)-(383)] to conclude that there exists a sufficiently large $R>0$ such that $K_{1/R}$ satisfies (38). Consequently, the laws of $({{\bf{v}}}^{\sigma},{{\bf{d}}}^{\sigma},{{\bf{z}}},W)$, which will be denoted as $\xi^{\sigma}$ subsequently, are tight over the phase space $\mathcal{G}$:
$\mathcal{G}=L^{p_{1}}(Q_{T})\times L^{p_{2}}(Q_{T})\times C([0,T];H)\times C_{0}([0,T];H),1<p_{1}<\frac{10}{3},1<p_{2}<10.$
[…]
in the topology of $\mathcal{G}$. Additionally, setting $\{\widetilde{\mathcal{F}}_{t}\}_{t\geq 0}$ as the completion of the $\sigma$-algebra generated by $\{\widetilde{U}(s);s\leq t\}$, we obtain a new stochastic basis $\widetilde{\mathfrak{B}}=\left(\widetilde{\Omega},\widetilde{\mathcal{F}},\{\widetilde{\mathcal{F}}_{t}\}_{t\geq 0},\widetilde{\mathbb{P}},\left(\widetilde{W}(t)\right)_{t\geq 0}\right)$, as in (16).
4.2.2 Preservation of the key energy equality
Next we show that for each $k$, the random variable $(\widetilde{{\bf{v}}}^{\sigma_{k}},\widetilde{{\pi}}^{\sigma_{k}},\widetilde{{\bf{d}}}^{\sigma_{k}},\widetilde{{{\bf{z}}}}^{k})$ satisfies the local energy equality (81). We adapt a technique due to Bensoussan [4]. Given $\varphi\in C_{0}^{\infty}(Q_{T};\mathbb{R}_{+})$, for $({{\bf{v}}}^{\sigma},{\pi}^{\sigma},{{\bf{d}}}^{\sigma},{{{\bf{z}}}})$, the components of ${U}_{\sigma}$, define the random variable $X^{\sigma}$ on $(\Omega,\mathcal{F})$ as
$\displaystyle X^{\sigma}$ $\displaystyle:=\left|X_{1}^{\sigma}-X_{2}^{\sigma}\right|,$ (40)
where
$\displaystyle X_{1}^{\sigma}$ $\displaystyle:=\int_{\mathbb{T}^{3}\times\{T\}}\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}^{\sigma}|^{2}}{2}+F({{\bf{d}}}^{\sigma})\right)\varphi+\int_{Q_{T}}(|\nabla{{\bf{v}}}^{\sigma}|^{2}+|\Delta{{\bf{d}}}^{\sigma}|^{2}+|f({{\bf{d}}}^{\sigma})|^{2})\varphi$
and
$\displaystyle X_{2}^{\sigma}$ $\displaystyle:=\int_{Q_{T}}\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}^{\sigma}|^{2}}{2}+F({{\bf{d}}}^{\sigma})\right)\partial_{t}\varphi+\int_{\mathbb{T}^{3}}[({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\cdot\nabla{{\bf{v}}}^{\sigma}]\cdot{{\bf{z}}}\varphi$
$\displaystyle+\int_{Q_{T}}\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}+{{\bf{v}}}^{\sigma}\cdot{\bf{z}}\right)({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\cdot\nabla\varphi+{\pi}^{\sigma}{{\bf{v}}}^{\sigma}\cdot\nabla\varphi+\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}^{\sigma}|^{2}}{2}\right)\Delta\varphi$
$\displaystyle+\int_{Q_{T}}\left(\nabla{{\bf{d}}}^{\sigma}\odot\nabla{{\bf{d}}}^{\sigma}-\frac{1}{2}|\nabla{{\bf{d}}}^{\sigma}|^{2}{\rm I}_{3}\right):\nabla^{2}\varphi$
$\displaystyle+\int_{Q_{T}}({{\bf{z}}}\cdot\nabla\Phi_{\sigma}[{{\bf{d}}}^{\sigma}])\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma}))+\int_{Q_{T}}[({{\bf{z}}}+{{\bf{v}}}^{\sigma})\cdot\nabla{{\bf{d}}}^{\sigma}]\cdot(\nabla\varphi\cdot\nabla){{\bf{d}}}^{\sigma}$
$\displaystyle-\int_{Q_{T}}f({{\bf{d}}}^{\sigma})(\nabla\varphi\cdot\nabla{{\bf{d}}}^{\sigma})-2\nabla f({{\bf{d}}}^{\sigma}):\nabla{{\bf{d}}}^{\sigma}\varphi.$
From (81) in Lemma 13, we infer that
$\displaystyle X_{1}^{\sigma}=X_{2}^{\sigma}\mbox{ with Probability 1}.$ (41)
[…]
$\widetilde{\mathbb{E}}\left[\frac{\widetilde{X}^{k}}{1+\widetilde{X}^{k}}\right]=\widetilde{\mathbb{E}}[g(\widetilde{U}_{k})]=\mathbb{E}[g(U_{\sigma_{k}})]=\mathbb{E}\left[\frac{X^{\sigma_{k}}}{1+X^{\sigma_{k}}}\right]=0,$
which implies that
[…]
as required.
4.2.3 Passage to the limit
[…]
$\begin{cases}&\partial_{t}\widetilde{{\bf{v}}}^{\sigma_{k}}-\Delta\widetilde{{\bf{v}}}^{\sigma_{k}}+(\widetilde{{\bf{z}}}^{k}+\Phi_{\sigma_{k}}[\widetilde{{\bf{v}}}^{\sigma_{k}}])\cdot\nabla(\widetilde{{\bf{z}}}^{k}+\widetilde{{\bf{v}}}^{\sigma_{k}})+\nabla\widetilde{\pi}^{\sigma_{k}}\\ &\quad\quad\quad\quad\quad\quad=-\nabla\Phi_{\sigma_{k}}[\widetilde{{\bf{d}}}^{\sigma_{k}}]\cdot(\Delta\widetilde{{\bf{d}}}^{\sigma_{k}}-f(\widetilde{{\bf{d}}}^{\sigma_{k}})),\\ &\nabla\cdot\widetilde{{\bf{v}}}^{\sigma_{k}}=0,\\ &\partial_{t}\widetilde{{\bf{d}}}^{\sigma_{k}}-(\widetilde{{\bf{z}}}^{k}+\widetilde{{\bf{v}}}^{\sigma_{k}})\cdot\nabla\Phi_{\sigma_{k}}[\widetilde{{\bf{d}}}^{\sigma_{k}}]=\Delta\widetilde{{\bf{d}}}^{\sigma_{k}}-f(\widetilde{{\bf{d}}}^{\sigma_{k}}).\end{cases}$ (43)
[…]
$\widetilde{{\pi}}^{\sigma_{k}}(\widetilde{\omega})\to\widetilde{\pi}(\widetilde{\omega})\mbox{ weakly in }L^{\frac{5}{3}}(Q_{T})\mbox{ up to the subsequence}.$ (44)
[…]
5 Partial regularity of the martingale suitable weak solutions
With the global existence of martingale suitable weak solutions in hand (Section 4), we now focus on the second result of Theorem 2, that is, to establish the partial regularity of these martingale suitable weak solutions.
5.1 Blowing up analysis
First, we recall the maximum principle [69, Lemma 2.2] for the $L^{\infty}$ of ${\bf{d}}$ from the transported Ginzburg–Landau heat flow :
$\left\{\begin{array}[]{l}\partial_{t}{{\bf{d}}}+{{\bf{u}}}\cdot\nabla{{\bf{d}}}=\Delta{{\bf{d}}}-f({{\bf{d}}}),\\ \nabla\cdot{\bf{u}}=0,\\ {{\bf{d}}}(0)={{\bf{d}}}_{0}.\end{array}\right$. (45)
Lemma 4.
For $T>0$, assume that ${{\bf{u}}}\in L^{2}(0,T;V)$ and ${{\bf{d}}}_{0}\in H^{1}(\mathbb{T};\mathbb{R}^{3})$ with $|{{\bf{d}}}_{0}(x)|\leq 1$ a.e. $x\in\mathbb{T}^{3}$. Suppose ${{\bf{d}}}\in L^{2}(0,T;H^{2}(\mathbb{T}^{3},\mathbb{R}^{3}))$ is a weak solution to (45). Then $|{{\bf{d}}}(x,t)|\leq 1$ for a.e. $(x,t)\in Q_{T}$.
The following lemma will play an essential role in the partial regularity of suitable weak solutions.
Lemma 5.
There exist $\varepsilon_{0}>0$, $0<\tau_{0}<\frac{1}{2}$ and $C_{0}>0$ such that if $({{\bf{v}}},{{\bf{d}}},{\pi})$ is a suitable weak solution in $\mathbb{T}^{3}\times(0,\infty)$, then for $z_{0}=(x_{0},t_{0})\in\mathbb{T}^{3}\times(r^{2},\infty)$, and $r>0$, if it holds that
$|{{\bf{z}}}|\leq M,\quad|{{\bf{d}}}|\leq M\text{ in }\mathcal{Q}_{r}(z_{0}),$
where $\mathcal{Q}_{r}(z_{0})=B_{r}(x_{0})\times(t_{0}-r^{2},t)$, and
$\Theta_{z_{0}}(r):=r^{-2}\int_{\mathcal{Q}_{r}(z_{0})}(|{{\bf{v}}}|^{3}+|\nabla{{\bf{d}}}|^{3})+\left(r^{-2}\int_{\mathcal{Q}_{r}(z_{0})}|\pi|^{\frac{3}{2}}\right)^{2}\leq\varepsilon_{0}^{3},$
then
$\displaystyle\Theta_{z_{0}}(\tau_{0}r)$ $\displaystyle=(\tau_{0}r)^{-2}\int_{\mathcal{Q}_{\tau_{0}r}(z_{0})}(|{{\bf{v}}}|^{3}+|\nabla{{\bf{d}}}|^{3})+\left((\tau_{0}r)^{-2}\int_{\mathcal{Q}_{\tau_{0}r}(z_{0})}|\pi|^{\frac{3}{2}}\right)^{2}$
$\displaystyle\leq\frac{1}{2}\max\{\Theta_{z_{0}}(r),C_{0}r^{3}\}.$
Proof.
Proof by contradiction. Suppose that it is false. Then there exists an $M_{0}>0$ such that for any $\tau\in(0,\frac{1}{2})$, there exist sequences $\varepsilon_{i}\to 0$, $C_{i}\to\infty$, $r_{i}>0$, and $z_{i}=(x_{i},t_{i})\in\mathbb{T}^{3}\times(r_{i}^{2},\infty)$ such that
$\left\{\begin{array}[]{l}|{{\bf{z}}}|\leq M_{0},,\quad|{{\bf{d}}}|\leq M_{0}\text{ in }\mathcal{Q}_{r_{i}}(z_{i}),\\ \Theta_{z_{i}}(r_{i})=\varepsilon_{i}^{3},\\ \Theta_{z_{i}}(\tau r_{i})>\frac{1}{2}\max\{\varepsilon_{i}^{3},C_{i}r_{i}^{3}\}.\end{array}\right$. (46)
First we notice that $C_{i}r_{i}^{3}\leq 2\Theta_{z_{i}}(\tau r_{i})\leq 2\tau^{-4}\Theta_{z_{i}}(r_{i})=2\tau^{-4}\varepsilon_{i}^{3}$. This implies that
$r_{i}\leq\left(\frac{2\varepsilon_{i}^{3}}{C_{i}\tau^{4}}\right)^{\frac{1}{3}}\to 0.$
Next, we define the rescaling sequence
$({{{\bf{z}}}}^{i},{{{\bf{v}}}}^{i},{{{\bf{d}}}}^{i},\pi^{i})(x,t)=(r_{i}{{\bf{z}}},r_{i}{{\bf{v}}},{{\bf{d}}},r_{i}^{2}{\pi})(x_{i}+r_{i}x,t_{i}+r_{i}^{2}t).$
It is straightforward to verify that $({{\bf{z}}}^{i},{{\bf{v}}}^{i},{{\bf{d}}}^{i},{\pi}^{i})$ satisfies the rescaled equation
$\left\{\begin{array}[]{l}\partial_{t}{{\bf{v}}}^{i}+({{\bf{z}}}^{i}+{{\bf{v}}}^{i})\cdot\nabla({{\bf{z}}}^{i}+{{\bf{v}}}^{i})+\nabla\pi^{i}\\ \qquad\qquad\qquad=\Delta{\bf{v}}^{i}-\mathop{\rm div}\nolimits\left(\nabla{{\bf{d}}}_{i}\odot\nabla{{\bf{d}}}^{i}-\frac{1}{2}|\nabla{{\bf{d}}}^{i}|^{2}{\rm I}_{3}-r_{i}^{2}F({{\bf{d}}}^{i}){\rm I}_{3}\right),\\ \nabla\cdot{{\bf{v}}}^{i}=0,\\ \partial_{t}{{\bf{d}}}^{i}+({{\bf{z}}}^{i}+{{\bf{v}}}^{i})\cdot\nabla{{\bf{d}}}^{i}=\Delta{{\bf{d}}}^{i}-r^{2}_{i}f({{\bf{d}}}^{i}).\end{array}\right$. (47)
And (46) becomes
$\left\{\begin{array}[]{l}\displaystyle\int_{\mathcal{Q}_{1}(0)}(|{{\bf{v}}}^{i}|^{3}+|\nabla{{\bf{d}}}^{i}|^{3})+\left(\int_{\mathcal{Q}_{1}(0)}|\pi^{i}|^{\frac{3}{2}}\right)^{2}=\varepsilon_{i}^{3},\\ \displaystyle\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}(|{{\bf{v}}}^{i}|^{3}+|\nabla{{\bf{d}}}^{i}|^{3})+\left(\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}|\pi^{i}|^{\frac{3}{2}}\right)^{3}>\frac{1}{2}\max\{\varepsilon_{i}^{3},C_{i}r_{i}^{3}\}.\end{array}\right$. (48)
Then we define the blow-up sequence
$(\widehat{{\bf{z}}}^{i},\widehat{{\bf{v}}}^{i},\widehat{{\bf{d}}}^{i},\widehat{\pi}^{i})=\left(\frac{{{\bf{z}}}^{i}}{\varepsilon_{i}},\frac{{{\bf{v}}}^{i}}{\varepsilon_{i}},\frac{{{\bf{d}}}^{i}-\overline{{{\bf{d}}}^{i}}}{\varepsilon_{i}},\frac{{\pi}^{i}}{\varepsilon_{i}}\right),$
where $\overline{{{\bf{d}}}^{i}}={\fint}_{\mathcal{Q}_{1}(0)}{{\bf{d}}}^{i}$, which denotes the average of ${{\bf{d}}}^{i}$ over $\mathcal{Q}_{1}(0)$. Then it turns out that $(\widehat{{\bf{z}}}^{i},\widehat{{\bf{v}}}^{i},\widehat{{\bf{d}}}^{i},\widehat{\pi}^{i})$ solves the following equation:
$\left\{\begin{array}[]{l}\partial_{t}\widehat{{\bf{v}}}^{i}+\varepsilon_{i}(\widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i})\cdot\nabla(\widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i})+\nabla\widehat{\pi}^{i}\\ \quad\quad\quad=\Delta\widehat{{\bf{v}}}^{i}-\mathop{\rm div}\nolimits\left(\varepsilon_{i}\nabla\widehat{{\bf{d}}}^{i}\odot\nabla\widehat{{\bf{d}}}^{i}-\frac{\varepsilon_{i}}{2}|\nabla\widehat{{\bf{d}}}^{i}|^{2}{\rm I}_{3}-\frac{r_{i}^{2}}{\varepsilon_{i}}F({{\bf{d}}}^{i}){\rm I}_{3}\right),\\ \mathop{\rm div}\nolimits\widehat{{\bf{v}}}^{i}=0,\\ \partial_{t}\widehat{{\bf{d}}}^{i}+\varepsilon_{i}(\widehat{{\bf{z}}}^{i}+\widehat{\bf{v}}^{i})\cdot\nabla\widehat{{\bf{d}}}^{i}=\Delta\widehat{{\bf{d}}}^{i}-\frac{r_{i}^{2}}{\varepsilon_{i}}f({\bf{d}}^{i}).\end{array}\right$. (49)
Furthermore, we have
$\left\{\begin{array}[]{l}\int_{\mathcal{Q}_{1}(0)}\widehat{{\bf{d}}}^{i}=0,\\ \int_{\mathcal{Q}_{1}(0)}\left(|\widehat{{\bf{v}}}^{i}|^{3}+|\nabla\widehat{{\bf{d}}}^{i}|^{3}\right)+\left(\int_{\mathcal{Q}_{1}(0)}|\widehat{\pi}^{i}|^{\frac{3}{2}}\right)^{2}=1,\\ \tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}\left(|\widehat{{\bf{v}}}^{i}|^{3}+|\nabla\widehat{{\bf{d}}}^{i}|^{3}\right)+\left(\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}|\widehat{\pi}^{i}|^{\frac{3}{2}}\right)^{2}\\ \quad\quad>{\frac{1}{2}\max\left\{1,C_{i}\left(\frac{r_{i}}{\varepsilon_{i}}\right)^{3}\right\}.}\end{array}\right$. (50)
From (50), by weak compactness, we can show there exists
$(\widehat{{\bf{v}}},\widehat{{\bf{d}}},\widehat{\pi})\in L^{3}(\mathcal{Q}_{1}(0))\times L_{t}^{3}W_{x}^{1,3}(\mathcal{Q}_{1}(0))\times L^{\frac{3}{2}}(\mathcal{Q}_{1}(0)),$
such that, after passing to the limit upon a subsequence, we have
[…]
It follows from (50) and the lower semicontinuity that
$\int_{\mathcal{Q}_{1}(0)}\left(|\widehat{{\bf{v}}}|^{3}+|\nabla\widehat{\bf{d}}|^{3}\right)+\left(\int_{\mathcal{Q}_{1}(0)}|\widehat{\pi}|^{\frac{3}{2}}\right)\leq 1.$
We claim that
$\sup_{i}\left(\|\widehat{{\bf{v}}}^{i}\|_{L_{t}^{\infty}L_{x}^{2}\cap L_{t}^{2}H_{x}^{1}({\mathcal{Q}_{1/2}(0)})}+\|\widehat{{\bf{d}}}^{i}\|_{L_{t}^{\infty}H_{x}^{1}\cap L_{t}^{2}H_{x}^{2}({\mathcal{Q}_{1/2}(0)})}\right)<\infty.$
We choose a cut-off function $\varphi\in C_{0}^{\infty}(\mathcal{Q}_{1}(0))$ such that
$0\leq\varphi\leq 1,\quad\varphi\equiv 1\text{ on }\mathcal{Q}_{\frac{1}{2}}(0),\quad\text{ and }|\partial_{t}\varphi|+|\nabla\varphi|+|\nabla^{2}\varphi|\leq C.$
Define $\varphi_{i}(x,t)=\varphi\left(\frac{x-x_{i}}{r_{i}},\frac{t-t_{i}}{r_{i}^{2}}\right)$, $\forall(x,t)\in\mathbb{T}^{3}\times(0,\infty).$ Replacing $\varphi$ by $\varphi_{i}^{2}$ in the local energy inequality (81) we can show that
$\displaystyle\sup_{t_{i}-\frac{r_{i}^{2}}{4}\leq t\leq t_{i}}\int_{B_{r_{i}}(x_{i})}\left(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}+F({{\bf{d}}})\right)\varphi_{i}^{2}{\,\mathrm{d}x}$
$\displaystyle+\int_{\mathcal{Q}_{r_{i}/2}(z_{i})}\left(|\nabla{{\bf{v}}}|^{2}+|\Delta{{\bf{d}}}|^{2}+|f({{\bf{d}}})|^{2}\right)\varphi_{i}^{2}$
$\displaystyle\leq$ $\displaystyle C\int_{\mathcal{Q}_{r_{i}}(z_{i})}\left(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}+F({{\bf{d}}})\right)|\partial_{t}\left(\varphi_{i}^{2}\right)|$
$\displaystyle+C\int_{\mathcal{Q}_{r_{i}}(z_{i})}|({{\bf{z}}}+{{\bf{v}}})||{{\bf{z}}}||\nabla{{\bf{v}}}||\nabla\left(\varphi_{i}^{2}\right)|+|{\pi}||{{\bf{v}}}||\nabla\left(\varphi_{i}^{2}\right)|+\left(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}\right)\Delta\left(\varphi_{i}^{2}\right)$
$\displaystyle+C\int_{\mathcal{Q}_{r_{i}}(z_{i})}|\nabla{{\bf{d}}}|^{2}|\nabla^{2}\left(\varphi_{i}^{2}\right)|+|{{\bf{z}}}||\nabla{{\bf{d}}}||\Delta{{\bf{d}}}-f({{\bf{d}}})|^{2}|\varphi_{i}^{2}|+|{{\bf{z}}}+{{\bf{v}}}||\nabla{{\bf{d}}}|^{2}|\nabla\left(\varphi_{i}^{2}\right)|$
$\displaystyle+C\int_{\mathcal{Q}_{r_{i}}(z_{i})}|f({{\bf{d}}})||\nabla{{\bf{d}}}||\nabla\left(\varphi_{i}^{2}\right)|+|\nabla f({{\bf{d}}})||\nabla{{\bf{d}}}|\varphi_{i}^{2}.$
By rescaling and Young’s inequality, we can show that for every $i$,
$\displaystyle\sup_{-\frac{1}{4}\leq t\leq 0}\int_{B_{\frac{1}{2}(0)}}(|\widehat{{\bf{v}}}^{i}|^{2}+|\nabla\widehat{{\bf{d}}}^{i}|^{2})+\int_{\mathcal{Q}_{\frac{1}{2}}(0)}(|\nabla\widehat{{\bf{v}}}^{i}|+|\nabla^{2}\widehat{{\bf{d}}}^{i}|^{2})\leq C.$ (51)
Hence the claim holds. Now we intend to apply the Aubin–Lions lemma; up to now what we still need is to obtain estimates on $\partial_{t}\widehat{{\bf{v}}}^{i}$ and $\partial_{t}\widehat{{\bf{d}}}^{i}$. To achieve this goal, firstly, we observe that from the equations for $\widehat{{\bf{v}}}^{i}$ in (49)1,2, we can show that, for any test functions $\zeta\in C_{c,\rm{div}}^{\infty}(B_{\frac{1}{2}(0)};\mathbb{R}^{3})$ we have
$\left|\int_{Q_{\frac{1}{2}}(0)}\partial_{t}\widehat{{\bf{v}}}^{i}\cdot\zeta\right|$ $\leq\varepsilon_{i}\left|\int_{Q_{\frac{1}{2}}(0)}(\widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i})\cdot((\widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i})\cdot\nabla\zeta)\right|+\left|\int_{Q_{\frac{1}{2}}(0)}\nabla\widehat{{\bf{v}}}^{i}:\nabla\zeta\right|$ (52)
$+\left|\int_{Q_{\frac{1}{2}(0)}}\varepsilon_{i}\left(\nabla\widehat{{\bf{d}}}^{i}\odot\nabla\widehat{{\bf{d}}}^{i}-\frac{\varepsilon_{i}}{2}|\nabla\widehat{{\bf{d}}}^{i}{\rm I_{3}}-\frac{r_{i}^{2}}{\varepsilon_{i}}F({{\bf{d}}}^{i}){\rm I}_{3}\right):\nabla\zeta\right|$
$\leq C\|\widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i}\|_{L^{\infty}_{t}L_{x}^{2}(Q_{\frac{1}{2}}(0))}^{\frac{1}{2}}\|\ \widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i}\|^{\frac{3}{2}}_{L_{t}^{2}L_{x}^{6}(Q_{\frac{1}{2}}(0))}\|\nabla\zeta\|_{L_{t}^{4}{L_{x}^{2}}(Q_{\frac{1}{2}}(0))}$
$+C\|\widehat{{\bf{v}}}^{i}\|_{L_{t}^{2}H_{x}^{1}(Q_{\frac{1}{2}}(0))}\|\nabla\zeta\|_{L_{t}^{2}L_{x}^{2}(Q_{\frac{1}{2}}(0))}$
$+C\|\nabla\widehat{{\bf{d}}}^{i}\|_{L_{t}^{\infty}L_{x}^{2}(Q_{\frac{1}{2}}(0))}^{\frac{1}{2}}\|\nabla\widehat{{\bf{d}}}^{i}\|_{L_{t}^{2}L_{x}^{6}(Q_{\frac{1}{2}}(0))}^{\frac{3}{2}}\|\nabla\zeta\|_{L_{t}^{4}{L_{x}^{2}}(Q_{\frac{1}{2}}(0))}$
$+C\|F({{\bf{d}}}^{i})\|_{L_{t}^{2}L_{x}^{2}(Q_{\frac{1}{2}}(0))}\|\nabla\zeta\|_{L_{t}^{2}L_{x}^{2}(Q_{\frac{1}{2}}(0))}$
$\leq C\|\zeta\|_{L_{t}^{4}{\bf V}(Q_{\frac{1}{2}}(0))},$
where we have used the uniform estimates already obtained in (51), together with the uniform $L^{\infty}$ bounds for ${{\bf{d}}}$ and ${{\bf{z}}}$ provided in (46). Then from (52) we obtain the following uniform estimate:
$\|\partial_{t}\widehat{{\bf{v}}}^{i}\|_{L_{t}^{\frac{4}{3}}{\bf V}^{\prime}(Q_{\frac{1}{2}}(0))}\leq C,\quad\forall i=1,2,,\cdots.$ (53)
Now for the equations for $\widehat{{\bf{d}}}^{i}$ in (49)3, similarly, we can conclude that for any test function $\phi\in C_{0}^{\infty}(B_{\frac{1}{2}}(0))$ we have
$\left|\int_{Q_{\frac{1}{2}}(0)}\partial_{t}\widehat{{\bf{d}}}^{i}\cdot\phi\right|$ $\leq\varepsilon_{i}\left|\int_{Q_{\frac{1}{2}}(0)}((\widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i})\otimes\widehat{{\bf{d}}}^{i}):\nabla\phi\right|+\left|\int_{Q_{\frac{1}{2}}(0)}\nabla\widehat{{\bf{d}}}^{i}:\nabla\phi\right|$
$+\frac{r_{i}^{2}}{\varepsilon_{i}}\left|\int_{Q_{\frac{1}{2}}(0)}f({{\bf{d}}}^{i})\cdot\phi\right|$
$\leq C\|\widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i}\|_{L_{t}^{2}L_{x}^{2}(Q_{\frac{1}{2}}(0))}\|\widehat{{\bf{d}}}^{i}\|_{L_{t}^{\infty}L_{x}^{\infty}(Q_{\frac{1}{2}}(0))}\|\nabla\phi\|_{L_{t}^{2}L_{x}^{2}(Q_{\frac{1}{2}}(0))}$
$+C\|\nabla\widehat{{\bf{d}}}^{i}\|_{L_{t}^{2}L_{x}^{2}(Q_{\frac{1}{2}}(0))}\|\nabla\phi\|_{L_{t}^{2}L_{x}^{2}(Q_{\frac{1}{2}}(0))}$
$+C\|f({{\bf{d}}}^{i})\|_{L_{t}^{2}L_{x}^{2}(Q_{\frac{1}{2}}(0))}\|\phi\|_{L_{t}^{2}L_{x}^{2}(Q_{\frac{1}{2}}(0))}$
$\leq C\|\phi\|_{L_{t}^{2}H_{x}^{1}(Q_{\frac{1}{2}}(0))}.$
This implies that
$\|\partial_{t}\widehat{{\bf{d}}}^{i}\|_{L_{t}^{2}H_{x}^{-1}(Q_{\frac{1}{2}}(0))}\leq C,\quad\forall i=1,2,\cdots.$ (54)
With the estimates on $\partial_{t}\widehat{{\bf{v}}}^{i}$ and $\partial_{t}\widehat{{\bf{d}}}^{i}$ in hand, now we are ready to apply the Aubin–Lions lemma. We find that
[…]
By the standard interior regularity estimate we can conclude that $({{\widehat{{\bf{v}}}}},\widehat{\textbf{d}})\in C^{\infty}(\mathcal{Q}_{\frac{1}{4}}(0))$ and $\widehat{\pi}\in L^{\infty}([-\frac{1}{16},0],C^{\infty}(B_{\frac{1}{4}}(0)))$ with the following estimate:
$\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}(|{{\widehat{{\bf{v}}}}}|^{3}+|\nabla\widehat{\textbf{d}}|^{3})+\left(\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}|\widehat{\pi}|^{\frac{3}{2}}\right)^{2}$ (56)
$\leq C\tau^{3}\left[\int_{\mathcal{Q}_{\frac{1}{2}}(0)}(|{{\widehat{{\bf{v}}}}}|^{3}+|\nabla\widehat{\textbf{d}}|^{3})+\left(\int_{\mathcal{Q}_{\frac{1}{2}}(0)}|\widehat{\pi}|^{\frac{3}{2}}\right)^{2}\right]$
$\leq C\tau^{3},\forall\tau\in\left(0,\frac{1}{8}\right).$
From the strong convergence of $(\widehat{{\bf{v}}}^{i},\widehat{{\bf{d}}}^{i})$ we infer that
$\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}(|{{\widehat{{\bf{v}}}}}^{i}|^{3}+|\nabla\widehat{{\bf{d}}}^{i}|^{3})=\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}(|\widehat{{\bf{v}}}|^{3}+|\nabla\widehat{{\bf{d}}}|^{3})+\tau^{-2}o(i),$
where $o(i)\to 0$ as $i\to\infty$. It remains to estimate the pressure term $\widehat{\pi}^{i}$. Taking the divergence in (49)1, we see that
$-\Delta\widehat{\pi}^{i}$ (57)
$=\varepsilon_{i}\mathop{\rm div}\nolimits^{2}\left[{\left(\widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i}\right)}\otimes{\left(\widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i}\right)}+\nabla\widehat{{\bf{d}}}^{i}\odot\nabla\widehat{{\bf{d}}}^{i}-\frac{1}{2}|\nabla\widehat{{\bf{d}}}^{i}|^{2}{\rm I}_{3}-\frac{r_{i}^{2}}{\varepsilon_{i}^{2}}F({{\bf{d}}}^{i}){\rm I}_{3}\right].$
We need to show that
$\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}|\widehat{\pi}^{i}|^{\frac{3}{2}}\leq C\tau^{-2}(\varepsilon_{i}+o(i))+C\tau.$ (58)
This will lead to a contradiction of (50)2 if we chose sufficient small $\tau_{0}\in(0,\frac{1}{4})$ and sufficient large $i_{0}$ such that for $i\geq i_{0}$, it holds that
$\tau_{0}^{-2}\int_{\mathcal{Q}_{\tau_{0}}(0)}\left(|\widehat{{\bf{v}}}^{i}|^{3}+|\nabla\widehat{{\bf{d}}}^{i}|^{3}\right)+\left(\tau_{0}^{-2}\int_{\mathcal{Q}_{\tau_{0}}(0)}|\widehat{\pi}_{i}|^{\frac{3}{2}}\right)^{2}\leq\frac{1}{4}.$ (59)
To prove (58), let $\eta\in C_{0}^{\infty}(B_{1}(0))$ be a cut-off function of $B_{\frac{3}{8}}(0)$. For any $t$, $-(\frac{3}{8})^{2}\leq t\leq 0$, we define
$\displaystyle\widehat{\pi}_{(1)}^{i}(x,t)$
$\displaystyle=$ $\displaystyle\int_{\mathbb{T}^{3}}\nabla_{x}^{2}G(x-y)\eta(y)\varepsilon_{i}\left[{\left(\widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i}\right)}\otimes{\left(\widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i}\right)}+\nabla\widehat{{\bf{d}}}^{i}\odot\nabla\widehat{{\bf{d}}}^{i}-\frac{1}{2}|\nabla\widehat{{\bf{d}}}^{i}|^{2}{\rm I}_{3}-\frac{r_{i}^{2}}{\varepsilon_{i}^{2}}F({{\bf{d}}}^{i}){\rm I}_{3}\right](y,t){\,\mathrm{d}y},$
where $G$ is the Green function of $-\Delta$ for $\mathbb{T}^{3}$. Then it is straightforward to see that $\widehat{\pi}_{(2)}^{i}(\cdot,t):=(\widehat{\pi}^{i}-\widehat{\pi}_{(1)}^{i})(\cdot,t)$ satisfies
$\Delta\widehat{\pi}_{(2)}^{i}=0\text{ in }B_{\frac{3}{8}}(0).$ (60)
Applying the Calderon–Zygmund theory we can show that
$\displaystyle\left\|\widehat{\pi}_{(1)}^{i}\right\|_{L^{\frac{3}{2}}(\mathbb{T}^{3})}$ $\displaystyle\leq C\varepsilon_{i}\left[\left\|{\widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i}}\right\|_{L^{3}(B_{1}(0))}^{2}+\left\|\nabla\widehat{{\bf{d}}}^{i}\right\|_{L^{3}(B_{1}(0))}^{2}+\frac{r_{i}^{2}}{\varepsilon_{i}^{2}}\|F({\bf{d}}^{i})\|_{L^{\frac{3}{2}}(B_{1}(0))}\right]$
$\displaystyle\leq C(\varepsilon_{i}+o(i)).$ (61)
From the mean value property of harmonic functions, we infer that for $0<\tau<\frac{1}{4}$,
$\displaystyle\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}|\widehat{\pi}_{(2)}^{i}|^{\frac{3}{2}}$ $\displaystyle\leq C\tau\int_{\mathcal{Q}_{\frac{1}{3}}(0)}|\widehat{\pi}_{(2)}^{i}|^{\frac{3}{2}}$
$\displaystyle\leq C\tau\int_{\mathcal{Q}_{\frac{1}{3}}(0)}(|\widehat{\pi}^{i}|^{\frac{3}{2}}+|\widehat{\pi}_{(1)}^{i}|^{\frac{3}{2}})$
$\displaystyle\leq C\tau(1+\varepsilon_{i}+o(i)).$ (62)
Combining (61) and (62) yields (58), and the contradiction (59) is achieved. ∎
5.2 An almost boundedness result
By iterating Lemma 5 and utilizing the Riesz potential in Morrey spaces (see [1, 47]) we can show the following local $L^{p}$ estimate for all $p>1$:
Lemma 6.
For any $M>0$, there exists $\varepsilon_{0}>0$ depending on $M$, such that for a suitable weak solution $({{\bf{v}}},{\pi},{{\bf{d}}})$ of (11) in ${\mathbb{T}^{3}}\times(0,\infty)$, if, for $z_{0}=(x_{0},t_{0})\in{\mathbb{T}^{3}}\times(r_{0}^{2},\infty)$, and $r_{0}>0$, it holds that $|{\bf{z}}|\leq M,|{{\bf{d}}}|\leq M$ in $\mathcal{Q}_{r_{0}}(z_{0})$, and $\Theta_{z_{0}}(r_{0})\leq\varepsilon_{0}^{3}$, then for any $1<p<\infty$, it holds true that
$\|({\bf{v}},\pi,\nabla{{\bf{d}}})\|_{L^{p}(\mathcal{Q}_{\frac{r_{0}}{4}}(z_{0}))}\leq C(p,\varepsilon_{0},M).$
Proof.
From the assumption, we have that $\Phi_{z}(r_{0}/2)\leq 8\varepsilon_{0}^{3}$ holds for any $z\in\mathcal{Q}_{\frac{r_{0}}{2}}(z_{0})$. Then we apply Lemma 5 repeatedly on $\mathcal{Q}_{\frac{r_{0}}{2}}(z)$, and we have that for any $k\geq 1$ and $\tau_{0}\in(0,1/4)$,
$\Theta_{z}(\tau_{0}^{k}r_{0})\leq 2^{-k}\max\{\Theta_{z}\left(\frac{r_{0}}{2}\right),\frac{C_{0}r_{0}^{3}}{1-2\tau_{0}^{3}}\}.$
Hence, it holds for $0<s<\frac{r_{0}}{2}$, and $z\in\mathcal{Q}_{\frac{r_{0}}{2}}(z_{0})$ that
$s^{-2}\int_{\mathcal{Q}_{s}(z)}(|{\bf{v}}|^{2}+|\nabla{{\bf{d}}}|^{3}+|\pi|^{\frac{3}{2}})\leq C(1+\varepsilon_{0}^{3})\left(\frac{s}{r_{0}}\right)^{3\theta_{0}},$ (63)
for $\theta_{0}=\frac{\ln 2}{3|\ln\tau_{0}|}\leq(0,\frac{1}{3})$. Now we apply the local energy inequality and boundedness of $({\bf{z}},{{\bf{d}}})$ on $\mathcal{Q}_{\frac{r_{0}}{2}}(z_{0})$, for $z\in\mathcal{Q}_{\frac{r_{0}}{2}}(z_{0})$ to show that
$s^{-1}\int_{\mathcal{Q}_{s}(z)}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})$ (64)
$\leq$ $C\left[(2s)^{-3}\int_{\mathcal{Q}_{2s}(z)}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})+(2s)^{-2}\int_{\mathcal{Q}_{2s}(z)}(|{{\bf{v}}}|^{3}+|\nabla{{\bf{d}}}|^{2}+|P|^{\frac{3}{2}})\right]$
$\leq$ $C(1+\varepsilon_{0}^{3})\left(\frac{s}{r_{0}}\right)^{2\theta_{0}}.$
For any open set $U\subset\mathbb{T}^{3}\times\mathbb{R}_{+}$, $1\leq p<\infty$, $0\leq\lambda\leq 5$, the Morrey space $M^{p,\lambda}$ is defined by
$M^{p,\lambda}(U):=\left\{f\in L^{p}_{\rm loc}(U):\|f\|^{p}_{M^{p,\lambda}(U)}=\sup_{z\in U,r>0}r^{\lambda-5}\int_{\mathcal{Q}_{r}(z)}|f|^{p}\,\mathrm{d}x\,\mathrm{d}t<\infty\right\}.$
Then (63) and (64) read
$\left\{\begin{aligned} &({{\bf{v}}},\nabla{{\bf{d}}})\in M^{3,3(1-\alpha)}\left(\mathcal{Q}_{\frac{r_{0}}{2}}(z_{0})\right),\\ &\pi\in M^{\frac{3}{2},3(1-\alpha)}\left(\mathcal{Q}_{\frac{r_{0}}{2}}(z_{0})\right),\\ &(\nabla{{\bf{v}}},\nabla^{2}{{\bf{d}}})\in M^{2,4-2\alpha}(\mathcal{Q}_{\frac{r_{0}}{2}}(z_{0})).\end{aligned}\right$.
Now we view (4)3 as
$\partial_{t}{{\bf{d}}}-\Delta{{\bf{d}}}=-{{\bf{u}}}\cdot\nabla{{\bf{d}}}-f({{\bf{d}}})\in M^{\frac{3}{2},3(1-\alpha)}\left(\mathcal{Q}_{\frac{r_{0}}{2}}(z_{0})\right).$
Let $\eta$ be the cut off function of $\mathcal{Q}_{\frac{r_{0}}{2}}(z_{0})$ with $|\partial_{t}\eta|+|\nabla^{2}\eta|\leq Cr_{0}^{-2}$, let $\widecheck{\bf d}=\eta^{2}({{\bf{d}}}-({{\bf{d}}})_{z_{0},r_{0}})$, where $({{\bf{d}}})_{z_{0},r_{0}}$ is the average of ${{\bf{d}}}$ over $\mathcal{Q}_{\frac{r_{0}}{2}}(z_{0})$. Then we have
$\partial_{t}\widecheck{{\bf{d}}}-\Delta\widecheck{\bf d}=\mathsf{F},\quad\mathsf{F}:=\eta^{2}(-{{\bf{u}}}\cdot\nabla{{\bf{d}}}-f({{\bf{d}}}))+(\partial_{t}\eta^{2}-\Delta\eta^{2})({{\bf{d}}}-({{\bf{d}}})_{z_{0},r_{0}})-\nabla\eta^{2}\cdot\nabla{{\bf{d}}},$
where $\mathsf{F}$ satisfies
$\|\mathsf{F}\|_{M^{\frac{3}{2},3(1-\alpha)}(\mathbb{R}^{4})}\leq C(1+\varepsilon_{0}).$
By the Duhamel formula, we have that
$|\nabla\widecheck{\bf d}(x,t)|\leq\int_{0}^{t}\int_{\mathbb{R}^{3}}|\nabla\Gamma(x-y,t-s)||\mathsf{F}(y,s)|dyds\leq C\mathcal{I}_{1}(|\mathsf{F}|)(x,t),$ (65)
where $\Gamma$ is the heat kernel in $\mathbb{R}^{3}$, and $\mathcal{I}_{\beta}$ is the Riesz potential of order $\beta$ on $\mathbb{R}^{4}$, $\beta\in[0,4]$, defined by
$I_{\beta}(g)(x,t):=\int_{\mathbb{R}^{4}}\frac{|g(y,s)|}{\delta_{p}^{5-\beta}((x,t),(y,s))}dyds.$
Applying the Riesz potential estimates (for a reference on the Riesz potential estimates, see e.g. [45]), we conclude that
$\|\nabla{\widecheck{\bf d}}\|_{M^{\frac{3(1-\alpha)}{1-2\alpha},3(1-\alpha)}(\mathbb{R}^{4})}\leq C\|\mathsf{F}\|_{M^{\frac{3}{2},3(1-\alpha)}(\mathbb{R}^{4})}\leq C(1+\varepsilon_{0}).$ (66)
Since ${{\bf{d}}}-{\widecheck{\bf d}}$ is caloric in $\mathcal{Q}_{\frac{r_{0}}{2}}(z_{0})$,
$\|\nabla({{\bf{d}}}-{\widecheck{\bf d}})\|_{L^{p}(\mathcal{Q}_{\frac{r_{0}}{2}(z_{0})})}\leq C(p,r_{0},\varepsilon_{0}).$ (67)
On the other hand, let $\widecheck{{\bf{v}}}$ solve the following Stokes equations
$\left\{\begin{array}[]{l}\partial_{t}\widecheck{{\bf{v}}}-\Delta\widecheck{{\bf{v}}}+\nabla\widecheck{\pi}\\ =-\nabla\cdot[\eta^{2}({{\bf{u}}}\otimes{{\bf{u}}})+(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}|\nabla{{\bf{d}}}|^{2}{\rm I}_{3})-(F({{\bf{d}}})-(F({{\bf{d}}}))_{z_{0},r_{0}}){\rm I}_{3}]\\ =:-\mathop{\rm div}\nolimits\mathsf{X},\\ \\ \nabla\cdot\widecheck{{\bf{v}}}=0,\\ \widecheck{{\bf{v}}}(\cdot,0)=0.\end{array}\right$. (68)
Now we apply the Oseen kernel for estimations of $\widecheck{{\bf{v}}}$; for references about the Oseen kernel, we refer the readers to e.g. [78, 59, 58] Particularly, by using the decaying property of the Oseen kernel (see Proposition 11.1 in [58]), then with a computation similar to (65), we can arrive at
$|\widecheck{{\bf{v}}}(x,t)|\leq C\int_{0}^{t}\int_{\mathbb{R}^{3}}\frac{|{\mathsf{X}}(y,s)|}{\delta_{p}^{4}((x,t),(y,s))}dyds\leq C\mathcal{I}_{1}(|\mathsf{X}|)(x,t).$ (69)
Note that
$\displaystyle\|\mathsf{X}\|_{M^{\frac{3}{2},3(1-\alpha)}}\leq C(1+\varepsilon_{0}).$
This yields that
$\|\widecheck{{\bf{v}}}\|_{M^{\frac{3(1-\alpha)}{1-2\alpha},3(1-\alpha)}(\mathbb{R}^{4})}\leq C\|\mathsf{X}\|_{M^{\frac{3}{2},3(1-\alpha)}(\mathbb{R}^{4})}\leq C(1+\varepsilon_{0}).$ (70)
Meanwhile, we have $({{\bf{v}}}-\widecheck{{\bf{v}}})$ solves the linear homogeneous Stokes equation in $\mathcal{Q}_{\frac{r_{0}}{2}}(z_{0})$ and
$\|{{\bf{v}}}-\widecheck{{\bf{v}}}\|_{L^{p}(\mathcal{Q}_{\frac{r_{0}}{4}}(z_{0}))}\leq C(p,r_{0},\varepsilon_{0}).$ (71)
Bootstrapping $\alpha$ in the previous estimates (66), (67), (70) and (71) to get $\alpha\uparrow\frac{1}{2}$, and then by the embedding theorem for the Morrey spaces, we obtain that
$\|({{\bf{v}}},\nabla{{\bf{d}}})\|_{L^{p}(\mathcal{Q}_{\frac{r_{0}}{4}}(z_{0}))}\leq C(p,r_{0},\varepsilon_{0}).$
The estimate for ${\pi}$ comes directly from the standard $L^{p}$ estimate for the Poisson equation:
$\Delta{\pi}=\mathop{\rm div}\nolimits^{2}[{{\bf{u}}}\otimes{{\bf{u}}}+(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}|\nabla{{\bf{d}}}|^{2}{\rm I}_{3})-(F({{\bf{d}}})-(F({{\bf{d}}}))_{z_{0},r_{0}})].$
5.3 The A-B-C-D Lemma and Criteria for the regular points
Now we are ready to verify the first criteria for the regular points:
Proposition 7.
If $z_{0}$ satisfies the assumption in Lemma 5, then $z_{0}$ is a regular point.
Proof.
By Lemma 6, we have that $({{\bf{u}}},{\pi},\nabla{{\bf{d}}})\in L^{p}(\mathcal{Q}_{r_{0}/4}(z_{0}))$ for $1<p<\infty$. Applying the regularity estimate for the generalized Stokes system (cf. [51, Lemma 2.2]) to the ${\bf{v}}$ equation in (11)1, we get that ${\bf{v}}\in C^{\alpha}(\mathcal{Q}_{\frac{r_{0}}{8}}(z_{0}))$ for some $\alpha\in(0,1)$. On the other hand, the ${{\bf{d}}}$ equation reads
$\partial_{t}{{\bf{d}}}-\Delta{{\bf{d}}}=-({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{d}}}-f({{\bf{d}}})\in L^{p}(\mathcal{Q}_{\frac{r_{0}}{4}}(z_{0})).$
By the standard $W^{2,p}$ estimate we can get that $\|(\partial_{t}{{\bf{d}}},\nabla^{2}{{\bf{d}}})\|_{L^{p}(\mathcal{Q}_{\frac{r_{0}}{8}}(z_{0}))}<C(p,r_{0},\varepsilon_{0})$. And the Sobolev embedding implies $\nabla{{\bf{d}}}\in C^{\alpha}(\mathcal{Q}_{\frac{r_{0}}{8}}(z_{0})),$ thus $z_{0}$ is a regular point. ∎
In fact we can improve Proposition 7 into the following theorem:
Theorem 8.
There exists $\varepsilon_{1}>0$ such that if $({{\bf{u}}},P,{{\bf{d}}})$ is a suitable weak solution to (4), which satisfies, for $z_{0}\in\mathbb{T}^{3}\times(0,\infty)$,
$\limsup_{r\to 0}\frac{1}{r}\int_{\mathcal{Q}_{r}(z_{0})}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})\leq\varepsilon_{1}^{2},$ (72)
then $({{\bf{u}}},\nabla{{\bf{d}}})$ is bounded near $z_{0}$.
For simplicity, we assume that $z_{0}=0$. We generalize the dimensionless estimates in [11, 65] with the following quantities:
$\displaystyle A(r)$ $\displaystyle:=\sup_{-r^{2}\leq t\leq 0}r^{-1}\int_{B_{r}(0)\times\{t\}}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}),$
$\displaystyle B(r)$ $\displaystyle:=r^{-1}\int_{\mathcal{Q}_{r}(0)}\left(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2}\right),$
$\displaystyle C(r)$ $\displaystyle:=r^{-2}\int_{Q_{r}(0)}(|{{\bf{v}}}|^{3}+|\nabla{{\bf{d}}}|^{3}),$
$\displaystyle D(r)$ $\displaystyle:=r^{-2}\int_{\mathcal{Q}_{r}(0)}|{\pi}|^{\frac{3}{2}},$
To prove Theorem 8, we will need some preparations. The first step is to note that we have the following interpolation lemma.
Lemma 9.
For $v\in H^{1}(\mathbb{T}^{3})$, we have
$\int_{B_{r}(0)}|v|^{q}(x,t){\,\mathrm{d}x}\lesssim\left(\int_{B_{r}(0)}|\nabla v|^{2}(x,t){\,\mathrm{d}x}\right)^{\frac{q}{2}-a}\left(\int_{B_{r}(0)}|v|^{2}(x,t)\right)^{a}$ (73)
$+r^{3(1-\frac{q}{2})}\left(\int_{B_{r}(0)}|v|^{2}(x,t){\,\mathrm{d}x}\right)^{\frac{q}{2}},$
for every $B_{r}(0)\subset\mathbb{T}^{3}$, $2\leq q\leq 6$, $a=\frac{3}{2}(1-\frac{q}{6}).$
With application of Lemma 9 we can establish the following result.
Lemma 10.
For any ${{\bf{u}}}\in L^{\infty}([-\rho^{2},0];L^{2}(B_{\rho}(0)))\cap L^{2}([-\rho^{2},0];H^{1}(B_{\rho}(0)))$, ${{\bf{d}}}\in L^{\infty}([-\rho^{2},0];H^{1}(B_{\rho}(0)))\cap L^{2}([-\rho^{2},0];H^{2}(B_{\rho}(0)))$, it holds that for any $0<r\leq\rho$,
$C(r)\lesssim\left(\frac{r}{\rho}\right)^{3}A^{\frac{3}{2}}(\rho)+\left(\frac{\rho}{r}\right)^{3}A^{\frac{3}{4}}(\rho)B^{\frac{3}{4}}(\rho).$ (74)
Proof.
By Poincaré’s inequality, we can show that for $0<r\leq\rho$,
$\displaystyle\int_{B_{r}(0)}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})$
$\displaystyle\lesssim$ $\displaystyle\int_{B_{r}(0)}\left(\left||{{\bf{v}}}|^{2}-(|{{\bf{v}}}|^{2})\rho\right|+\left||\nabla{{\bf{d}}}|^{2}-(|\nabla{{\bf{d}}}|^{2})_{\rho}\right|\right){\,\mathrm{d}x}+\left(\frac{r}{\rho}\right)^{3}\int_{B_{\rho}(0)}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}){\,\mathrm{d}x}$
$\displaystyle\lesssim$ $\displaystyle\rho\int_{B_{\rho}(0)}(|{{\bf{v}}}||\nabla{{\bf{v}}}|+|\nabla{{\bf{d}}}||\nabla^{2}{{\bf{d}}}|){\,\mathrm{d}x}+\left(\frac{r}{\rho}\right)^{3}\int_{B_{\rho}(0)}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}){\,\mathrm{d}x}$
$\displaystyle\lesssim$ $\displaystyle\rho^{\frac{3}{2}}\left(\rho^{-1}\int_{B_{\rho}(0)}|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}{\,\mathrm{d}x}\right)^{\frac{1}{2}}\left(\int_{B_{\rho}(0)}|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2}\right)^{\frac{1}{2}}$
$\displaystyle+\left(\frac{r}{\rho}\right)^{3}\int_{B_{\rho}(0)}\left(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}\right){\,\mathrm{d}x}$
$\displaystyle\lesssim$ $\displaystyle\rho^{\frac{3}{2}}A^{\frac{1}{2}}(\rho)\left(\int_{B_{\rho}(0)}|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2}{\,\mathrm{d}x}\right)^{\frac{1}{2}}+\left(\frac{r}{\rho}\right)^{3}A(\rho).$
Applying Lemma 9 with $q=3$, $a=\frac{3}{4}$ we can get
$\displaystyle\int_{B_{r}(0)}(|{{\bf{v}}}|^{3}+|\nabla{{\bf{d}}}|^{3}){\,\mathrm{d}x}$
$\displaystyle\lesssim$ $\displaystyle\rho^{\frac{4}{3}}\left(\int_{B_{r}(0)}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2}){\,\mathrm{d}x}\right)^{\frac{3}{4}}\left(\rho^{-1}\int_{B_{r}(0)}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}){\,\mathrm{d}x}\right)^{\frac{3}{4}}$
$\displaystyle+r^{-\frac{3}{2}}\left(\int_{B_{r}(0)}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}){\,\mathrm{d}x}\right)^{\frac{3}{2}}$
$\displaystyle\lesssim$ $\displaystyle\rho^{\frac{3}{4}}A^{\frac{3}{4}}(\rho)\left(\int_{B_{r}(0)}|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2}\right)^{\frac{3}{4}}+r^{-\frac{3}{2}}\left(\int_{B_{r}(0)}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})\right)^{\frac{3}{2}}$
$\displaystyle\lesssim$ $\displaystyle\left(\rho^{\frac{3}{4}}+\frac{\rho^{\frac{9}{4}}}{r^{\frac{3}{2}}}\right)\left(\int_{B_{r}(0)}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})\right)^{\frac{3}{4}}A^{\frac{3}{4}}(\rho)+\left(\frac{r}{\rho}\right)^{3}A^{\frac{3}{2}}(\rho),$
where in the last we apply the previous inequality. Integrating the inequality over $[-r^{2},0]$ and by Hölder’s inequality, we have
$\displaystyle C(r)$ $\displaystyle=\frac{1}{r^{2}}\int_{\mathcal{Q}_{r}(0)}(|{{\bf{v}}}|^{3}+|\nabla{{\bf{d}}}|^{3})$
$\displaystyle\lesssim\left(\frac{r}{\rho}\right)^{3}A^{\frac{3}{2}}(\rho)+\left(\rho^{\frac{3}{4}}+\frac{\rho^{\frac{9}{4}}}{r^{\frac{3}{2}}}\right)\int_{-r^{2}}^{0}\left(\int_{B_{r}(0)}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})\right)^{\frac{3}{4}}A^{\frac{3}{4}}(\rho)$
$\displaystyle\lesssim\left(\frac{r}{\rho}\right)^{3}A^{\frac{3}{2}}(\rho)+r^{-\frac{3}{2}}\rho^{\frac{3}{4}}\left(\rho^{\frac{3}{4}}+\frac{\rho^{\frac{9}{4}}}{r^{\frac{3}{2}}}\right)A^{\frac{3}{4}}(\rho)B^{\frac{3}{4}}(\rho)$
$\displaystyle=\left(\frac{r}{\rho}\right)^{3}A^{\frac{3}{2}}(\rho)+\left[\left(\frac{\rho}{r}\right)^{\frac{3}{2}}+\left(\frac{\rho}{r}\right)^{3}\right]A^{\frac{3}{4}}(\rho)B^{\frac{3}{4}}(\rho)$
$\displaystyle\lesssim\left(\frac{r}{\rho}\right)^{3}A^{\frac{3}{2}}(\rho)+\left(\frac{\rho}{r}\right)^{3}A^{\frac{3}{4}}(\rho)B^{\frac{3}{4}}(\rho).$
Next, we estimate the pressure function to get
Lemma 11.
Under the same assumption of Lemma 10, it holds that for any $0<r<\rho/2$,
$D(r)\lesssim\frac{r}{\rho}D(\rho)+\left(\frac{\rho}{r}\right)^{2}\left[A^{\frac{3}{4}}(\rho)B^{\frac{3}{4}}(\rho)+\rho^{\frac{3}{2}}B^{\frac{3}{4}}(\rho)+\rho^{3}\right].$ (75)
Proof.
Taking the divergence of (11)1 yields
$\displaystyle-\Delta{\pi}=$ $\displaystyle\mathop{\rm div}\nolimits^{2}\Big[{{\bf{v}}}\otimes{{\bf{v}}}+\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}(\nabla{{\bf{d}}}:\nabla{{\bf{d}}}){\rm I}_{3}+{{\bf{z}}}\otimes{{\bf{v}}}+{{\bf{v}}}\otimes{{\bf{z}}}+{{\bf{z}}}\otimes{{\bf{z}}}-F({{\bf{d}}}){\rm I}_{3}\Big]$
$\displaystyle=$ $\displaystyle\mathop{\rm div}\nolimits^{2}\Big[({{\bf{v}}}-({{\bf{v}}})_{\rho})\otimes({{\bf{v}}}-({{\bf{v}}})_{\rho})+(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho})\odot(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho})$
$\displaystyle-\frac{1}{2}(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}):(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}){\rm I}_{3}$
$\displaystyle+(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho})\odot(\nabla{{\bf{d}}})_{\rho}+(\nabla{{\bf{d}}})_{\rho}\odot(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho})$
$\displaystyle-\frac{1}{2}(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}):(\nabla{{\bf{d}}})_{\rho}{\rm I}_{3}$
$\displaystyle-\frac{1}{2}(\nabla{{\bf{d}}})_{\rho}\odot(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}){\rm I}_{3}$
$\displaystyle+{{\bf{z}}}\otimes({{\bf{v}}}-({{\bf{v}}})_{\rho})+({{\bf{v}}}-({{\bf{v}}})_{\rho})\otimes{{\bf{z}}}+{{\bf{z}}}\otimes{{\bf{z}}}-F({{\bf{d}}}){\rm I}_{3}\Big]$
$\displaystyle:=$ $\displaystyle\mathop{\rm div}\nolimits^{2}\mathcal{G}\text{ in }B_{\rho}.$
Here $({{\bf{v}}})_{\rho}$ denotes the average of ${{\bf{v}}}$ over $B_{\rho}$. Let $\eta\in C_{0}^{\infty}(\mathbb{T}^{3})$ be a cut-off function of $B_{\rho/2}$ such that $|\nabla\eta|\lesssim\rho^{-1}$. Define an auxiliary function by
$\displaystyle\widecheck{\pi}$ $\displaystyle(x)=-\int_{\mathbb{T}^{3}}\nabla_{y}^{2}G(x-y):\eta^{2}(y)\mathcal{G}(y){\,\mathrm{d}y},$
where $G$ is the corresponding kernel of Laplacian on $\mathbb{T}^{3}$. Then we have
$\Delta\widecheck{\pi}=\mathop{\rm div}\nolimits^{2}\mathcal{G}\text{ in }B_{\rho/2},$
and
$\Delta({\pi}-\widecheck{\pi})=0\text{ in }B_{\rho/2}.$
From the singular integral theory (Calderon–Zygmund theory) and the $L^{\infty}$ bound on $({{\bf{z}}},{{\bf{d}}})$ we can infer that
$\displaystyle\int_{\mathbb{T}^{3}}|\widecheck{\pi}|^{\frac{3}{2}}{\,\mathrm{d}x}\lesssim$ $\displaystyle\int_{\mathbb{T}^{3}}\eta^{3}\left(|{{\bf{v}}}-({{\bf{v}}})_{\rho}|^{3}+|\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}|^{3}\right)$
$\displaystyle+|(\nabla{{\bf{d}}})_{\rho}|^{\frac{3}{2}}\int_{\mathbb{T}^{3}}\eta^{3}|\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}|^{\frac{3}{2}}+\int_{\mathbb{T}^{3}}\eta^{3}|{{\bf{v}}}-({{\bf{v}}})_{\rho}|^{\frac{3}{2}}+\int_{\mathbb{T}^{3}}\eta^{3}$
$\displaystyle\lesssim$ $\displaystyle\int_{B_{\rho}}(|{{\bf{v}}}-({{\bf{v}}})_{\rho}|^{3}+|\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}|^{3})+|(\nabla{{\bf{d}}})_{\rho}|^{\frac{3}{2}}\int_{B_{\rho}}|\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}|^{\frac{3}{2}}$
$\displaystyle+\int_{B_{\rho}}(|{{\bf{v}}}-({{\bf{v}}})_{\rho}|^{\frac{3}{2}}+1)$
Notice that ${\pi}-\widecheck{\pi}$ is harmonic in $B_{\rho/2}$, and we have that for $0<r\leq\rho/2$,
$\displaystyle\frac{1}{r^{2}}\int_{B_{r}}|{\pi}-\widecheck{\pi}|^{\frac{3}{2}}$ $\displaystyle\lesssim\left(\frac{r}{\rho}\right)\frac{1}{\rho^{2}}\int_{B_{\rho/2}}|{\pi}-\widecheck{\pi}|^{\frac{3}{2}}$
$\displaystyle\lesssim\left(\frac{r}{\rho}\right)\left[\frac{1}{\rho^{2}}\int_{B_{\rho/2}}|{\pi}|^{\frac{3}{2}}+\frac{1}{\rho^{2}}\int_{B_{\rho/2}}|\widecheck{\pi}|^{\frac{3}{2}}\right].$
Integrating the expression above over $[-r^{2},0]$, we can show that
$\displaystyle\frac{1}{r^{2}}\int_{\mathcal{Q}_{r}(0)}|{\pi}|^{\frac{3}{2}}\lesssim$ $\displaystyle\left(\frac{r}{\rho}\right)\frac{1}{\rho^{2}}\int_{\mathcal{Q}_{\rho}}|{\pi}|^{\frac{3}{2}}+\left(\frac{\rho}{r}\right)^{2}\frac{1}{\rho^{2}}\int_{\mathcal{Q}_{\rho}}|\widecheck{\pi}|^{\frac{3}{2}}$
$\displaystyle\lesssim$ $\displaystyle\left(\frac{r}{\rho}\right)D(\rho)+\left(\frac{\rho}{r}\right)^{2}\Big[\frac{1}{\rho^{2}}\int_{\mathcal{Q}_{\rho}}|{{\bf{v}}}-({{\bf{v}}})_{\rho}|^{3}+|\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}|^{3}$
$\displaystyle+\frac{1}{\rho^{2}}(\sup_{-\rho^{2}\leq t\leq 0}|(\nabla{{\bf{d}}}(t))_{\rho}|^{\frac{3}{2}})\int_{\mathcal{Q}_{\rho}}|\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}|^{\frac{3}{2}}$
$\displaystyle+\frac{1}{\rho^{2}}\int_{\mathcal{Q}_{\rho}}|{{\bf{v}}}-({{\bf{v}}})_{\rho}|^{\frac{3}{2}}+\rho^{3}\Big].$
Now we apply the interpolation inequality
$\displaystyle\frac{1}{\rho^{2}}\int_{\mathcal{Q}_{\rho}}(|{{\bf{v}}}-({{\bf{v}}})_{\rho}|^{3}+|\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}|^{3})$
$\displaystyle\lesssim\left(\sup_{-\rho^{2}\leq t\leq 0}\frac{1}{\rho}\int_{B_{\rho}}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})\right)^{\frac{3}{4}}\left(\frac{1}{\rho}\int_{\mathcal{Q}_{\rho}}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})\right)^{\frac{3}{4}}$
$\displaystyle=A^{\frac{3}{4}}(\rho)B^{\frac{3}{4}}(\rho),$
and by Hölder’s inequality and Poincaré’s inequality, we can verify that
$\displaystyle\frac{1}{\rho^{2}}(\sup_{-\rho^{2}\leq t\leq 0}|(\nabla{{\bf{d}}})_{\rho}|^{\frac{3}{2}})\int_{\mathcal{Q}_{\rho}}|\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}|^{\frac{3}{2}}$
$\displaystyle\lesssim$ $\displaystyle\rho^{2}\left(\sup_{-\rho^{2}\leq t\leq 0}\int_{B_{\rho}}|\nabla{{\bf{d}}}|^{2}\right)^{\frac{3}{4}}\left(\frac{1}{\rho}\int_{Q_{\rho}}|\nabla^{2}{{\bf{d}}}|^{2}\right)^{\frac{3}{4}}$
$\displaystyle\leq$ $\displaystyle A^{\frac{3}{4}}(\rho)B^{\frac{3}{4}}(\rho),$
and
$\displaystyle\frac{1}{\rho^{2}}\int_{\mathcal{Q}_{\rho}}|{{\bf{v}}}-({{\bf{v}}})_{\rho}|^{\frac{3}{2}}\lesssim\frac{1}{\rho^{2}}\left(\int_{\mathcal{Q}_{\rho}}|{{\bf{v}}}-({{\bf{v}}})_{\rho}|^{2}\right)^{\frac{3}{4}}\lesssim\rho^{\frac{3}{2}}B^{\frac{3}{4}}(\rho).$
Now we have the interpolation estimate (74) and the pressure estimate (75), the proof of Theorem 8 is based on a dichotomy argument:
Proof of Theorem 8.
. For $\theta\in(0,\frac{1}{2})$ and $\rho\in(0,1)$, let $\varphi\in C_{0}^{\infty}(\mathcal{Q}_{\theta\rho}(0))$ be a function such that
$\varphi=1\text{ in }\mathcal{Q}_{\theta\rho/2}(0),\qquad|\nabla\varphi|\lesssim(\theta\rho)^{-1},\qquad|\nabla^{2}\varphi|+|\partial_{t}\varphi|\lesssim(\theta\rho)^{-2}.$
Applying the local energy inequality in (15) with $\varphi^{2}$, the $L^{\infty}$-bounded of $({{\bf{z}}},{{\bf{d}}})$, and integrating by parts, we obtain that
$\displaystyle\sup_{-(\theta\rho)^{2}\leq t\leq 0}\int_{\mathbb{T}^{3}}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})\varphi^{2}+\int_{\mathbb{T}^{3}\times[-(\theta\rho)^{2},0]}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})\varphi^{2}$
$\displaystyle\lesssim$ $\displaystyle\int_{\mathbb{T}^{3}\times[-(\theta\rho)^{2},0]}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})(|\partial_{t}\varphi|+|\nabla\varphi|^{2}+|\nabla^{2}\varphi|)$
$\displaystyle+\int_{\mathbb{T}^{3}\times[-(\theta\rho)^{2},0]}[(|{{\bf{v}}}|^{2}-(|{{\bf{v}}}|^{2})_{\rho})+|{\pi}||{{\bf{v}}}|]|\nabla\varphi|$
$\displaystyle+\int_{\mathbb{T}^{3}\times[-(\theta\rho)^{2},0]}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})\varphi^{2}+|\nabla{\bf{d}}|^{2}|{{\bf{v}}}||\nabla\varphi|$
$\displaystyle+\int_{\mathbb{T}^{3}\times[-(\theta\rho)^{2},0]}|{{\bf{z}}}|^{2}|\nabla\varphi|$
$\displaystyle+\int_{\mathbb{T}^{3}\times[-(\theta\rho)^{2},0]}\varphi^{2}.$
Hence we obtain that
$\displaystyle A\left(\frac{1}{2}\theta\rho\right)+B\left(\frac{1}{2}\theta\rho\right)$
$\displaystyle=$ $\displaystyle\sup_{-\left(\frac{\theta\rho}{2}\right)^{2}\leq t\leq 0}\frac{2}{\theta\rho}\int_{B_{\frac{\theta\rho}{2}}}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})+\frac{2}{\theta\rho}\int_{\mathcal{Q}_{\frac{\theta\rho}{2}}(0)}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})$
$\displaystyle\lesssim$ $\displaystyle\sup_{-(\theta\rho)^{2}\leq t\leq 0}\frac{1}{\theta\rho}\int_{\mathbb{T}^{3}}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})\varphi^{2}+\frac{1}{\theta\rho}\int_{\mathbb{T}^{3}\times[-(\theta\rho)^{2},0]}(|\nabla{{\bf{u}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})\varphi^{2}$
Putting together all the estimates, we have
$\displaystyle A(\frac{1}{2}\theta\rho)+B(\frac{1}{2}\theta\rho)$
$\displaystyle\lesssim$ $\displaystyle\left[C^{\frac{2}{3}}(\theta\rho)+A^{\frac{1}{2}}(\theta\rho)B^{\frac{1}{2}}(\theta\rho)C^{\frac{1}{3}}(\theta\rho)+C^{\frac{1}{3}}(\theta\rho)D^{\frac{2}{3}}(\theta\rho)+(\theta\rho)^{3}\right]$
$\displaystyle\lesssim$ $\displaystyle C^{\frac{2}{3}}(\theta\rho)+A(\theta\rho)B(\theta\rho)+D^{\frac{4}{3}}(\theta\rho)+(\theta\rho)^{3}.$
So that
$\displaystyle A^{\frac{3}{2}}(\frac{1}{2}\theta\rho)$ $\displaystyle\lesssim C(\theta\rho)+A^{\frac{3}{2}}(\theta\rho)B^{\frac{3}{2}}(\theta\rho)+D^{2}(\theta\rho)+(\theta\rho)^{\frac{9}{2}}.$
On the other hand, from Lemma 10 and 11 we can conclude that
$\displaystyle C(\theta\rho)$ $\displaystyle\lesssim\theta^{3}A^{\frac{3}{2}}(\rho)+\theta^{-3}A^{\frac{3}{4}}(\rho)B^{\frac{3}{4}}(\rho),$
$\displaystyle D^{2}(\theta\rho)$ $\displaystyle\lesssim\theta^{2}D^{2}(\rho)+\theta^{-4}(A^{\frac{3}{2}}(\rho)B^{\frac{3}{2}}(\rho)+\rho^{3}B^{\frac{3}{2}}(\rho)+\rho^{6}),$
and it is easy to see that
$A^{\frac{3}{2}}(\theta\rho)B^{\frac{3}{2}}(\theta\rho)\lesssim\theta^{-3}A^{\frac{3}{2}}(\rho)B^{\frac{3}{2}}(\rho).$
Therefore we conclude that for $0<\theta<\frac{1}{2}$,
$\displaystyle A^{\frac{3}{2}}(\frac{1}{2}\theta\rho)+D^{2}(\frac{1}{2}\theta\rho)$
$\displaystyle\leq c\left[\theta^{2}D^{2}(\rho)+\theta^{-4}(A^{\frac{3}{2}}(\rho)B^{\frac{3}{2}}(\rho)+\rho^{3}B^{\frac{3}{2}}(\rho)+\rho^{6})+\theta^{3}A^{\frac{3}{2}}(\rho)+\theta^{-8}A^{\frac{3}{2}}(\rho)B^{\frac{3}{2}}(\rho)+\theta^{2}+(\theta\rho)^{\frac{9}{2}}\right]$
$\displaystyle\leq c(\theta^{2}+\theta^{-8}B^{\frac{3}{2}}(\rho))(A^{\frac{3}{2}}(\rho)+D^{2}(\rho))+c(\theta^{-4}\rho^{3}B^{\frac{3}{2}}(\rho)+\theta^{-4}\rho^{6}+\theta^{2}+(\theta\rho)^{\frac{9}{2}}).$
For $\varepsilon_{0}$ given by Lemma 5, let $\theta_{0}\in(0,\frac{1}{2})$ be such that
$c\theta_{0}=\min\{\frac{1}{4},\frac{1}{8}\varepsilon_{1}^{2}\}.$
Since
$\limsup_{r\to 0}\frac{1}{r}\int_{\mathcal{Q}_{r}}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})<\varepsilon_{1}^{2}$
we can choose $\rho_{0}>0$ such that
$c\theta_{0}^{-8}B^{\frac{3}{2}}(\rho)\leq\frac{1}{4}\qquad\forall 0<\rho<\rho_{0},$
and
$c(\theta_{0}^{-4}\rho^{3}B^{\frac{3}{2}}(\rho)+\theta_{0}^{-4}\rho^{6}+\theta_{0}^{2}+(\theta_{0}\rho)^{\frac{9}{2}})\leq\frac{1}{2}\varepsilon_{1}^{2}\qquad\forall 0<\rho<\rho_{0}.$
Therefore we obtained that
$A^{\frac{3}{2}}(\frac{1}{2}\theta_{0}\rho)+D^{2}(\frac{1}{2}\theta_{0}\rho)\leq\frac{1}{2}(A^{\frac{3}{2}}(\rho)+D^{2}(\rho))+\frac{1}{2}\varepsilon_{1}^{2},\qquad\forall 0<\rho<\rho_{0}.$
Iterating this inequality yields that
$A^{\frac{3}{2}}((\frac{1}{2}\theta_{0})^{k}\rho)+D^{2}((\frac{1}{2}\theta_{0})^{k}\rho)\leq\frac{1}{2^{k}}(A^{\frac{3}{2}}(\rho)+D^{2}(\rho))+\varepsilon_{1}^{2}$
holds for all $0<\rho<\rho_{0}$ and $k\geq 1$. Furthermore, we can obtain from (74) that
$\displaystyle C((\frac{1}{2}\theta_{0})^{k}\rho)$ $\displaystyle\leq c[(\frac{1}{2}\theta_{0})^{3}A^{\frac{3}{2}}((\frac{1}{2}\theta_{0})^{k-1}\rho)+(\frac{1}{2}\theta_{0})^{-3}A^{\frac{3}{4}}((\frac{1}{2}\theta_{0})^{k-1}\rho)B^{\frac{3}{4}}((\frac{1}{2}\theta_{0})^{k-1}\rho)]$
$\displaystyle\leq c[(\frac{1}{2}\theta_{0})^{3}+(\frac{1}{2}\theta_{0})^{-3}\varepsilon_{1}^{\frac{3}{2}}][(\frac{1}{2})^{k-1}(A^{\frac{3}{2}}(\rho)+D^{2}(\rho))+\varepsilon_{1}^{2}]$
holds for all $0<\rho<\rho_{0}$ and $k\geq 1$. Then we can arrive at
$\limsup_{k\to\infty}\left[C((\frac{1}{2}\theta_{0})^{k}\rho)+D^{2}((\frac{1}{2}\theta_{0})^{k}\rho)\right]\leq c[1+(\frac{1}{2}\theta_{0})^{3}+(\frac{1}{2}\theta_{0})^{-3}\varepsilon_{1}^{\frac{3}{2}}]\varepsilon_{1}^{2}\leq\frac{1}{2}\varepsilon_{0}^{3}$
holds for all $\rho\in(0,\rho_{0})$ provided that $\varepsilon_{1}=\varepsilon_{1}(\theta_{0},\varepsilon_{0})>0$ is chosen sufficiently small. Hence there exists a $k_{0}>0$ such that
$C((\frac{1}{2}\theta_{0})^{k_{0}}\rho)+D((\frac{1}{2}\theta_{0})^{k_{0}}\rho)\leq\varepsilon_{0}^{3}$
which satisfies the assumption of Lemma 5, and as a consequence, $z_{0}=0$ is a regular point. ∎
5.4 The Hausdorff measure estimate
Now we have all the ingredients to prove Theorem 2. It is based on a standard covering argument which we detailed below for the reader’s convenience.
Proof of Theorem 2.
If $z\in\Sigma$, then by Theorem 8,
$\limsup_{r\to 0}\frac{1}{r}\int_{\mathcal{Q}_{r}(z)}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})\geq\varepsilon_{1}.$
Let $V$ be a neighborhood of $\Sigma$ and $\eta>0$ such that for all $z\in\Sigma$, then there exists a $r<\eta$ such that $\mathcal{Q}_{\eta}(z)\subset V$ and
$\frac{1}{r}\int_{\mathcal{Q}_{r}(z)}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})\geq\varepsilon_{1}.$
By Vitali’s covering lemma, we can find a pairwisely disjoint collection of $\mathcal{Q}_{r_{i}}(z_{i})$ such that
$\Sigma\subset\bigcup_{i=1}^{\infty}\mathcal{Q}_{5r_{i}}(z_{i}).$
Hence,
$\mathcal{P}_{5\eta}^{1}(\Sigma)$ $\leq\sum_{i=1}^{\infty}5r_{i}\leq\frac{5}{\varepsilon_{1}}\sum_{i=1}^{\infty}\int_{\mathcal{Q}_{r_{i}}(z_{i})}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})$ (76)
$\leq\frac{5}{\varepsilon_{1}}\int_{\bigcup_{i}\mathcal{Q}_{r_{i}}(z_{i})}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})$
$\leq\frac{5}{\varepsilon_{1}}\int_{V}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})<\infty.$
Then we can conclude that $\Sigma$ is of zero Lebesgue measure $\,\mathrm{d}x\,\mathrm{d}t$. Furthermore, we can choose $V$ with arbitrarily small measure, then by the absolute continuity of integral for $|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2}$, we can conclude from (76) that $\mathcal{P}^{1}(\Sigma)=0$. ∎
Declarations
  • The authors declare that no funds, grants, or other support were received during the preparation of this manuscript.
  • The authors have no relevant financial or non-financial interests to disclose.
  • The authors have no competing interests to declare that are relevant to the content of this article.
  • All authors certify that they have no affiliations with or involvement in any organization or entity with any financial interest or non-financial interest in the subject matter or materials discussed in this manuscript.
Appendix A Proofs of some preliminary results
For the reader’s convenience, we include in this appendix additional details for several results whose full derivations were too lengthy to present in the main text.
We begin by providing a proof of Lemma 1 from Section 3.2. To this end, we first establish the following preliminary result, which is essential to the proof.
Lemma 12 (A preliminary result).
Assuming that
$W(\omega)\in C^{\frac{1}{2}-\varepsilon}([0,\infty);D(A^{\frac{1}{4}+\beta}))\mbox{ $\displaystyle\mathbb{P}$-a.s. $\displaystyle\omega$, }$ (77)
for some $\beta>\varepsilon>0$, it follows that (13) holds.
Proof of Lemma 12.
From (10), we derive in the mild form
${{\bf{z}}}(t)=\int_{0}^{t}e^{-(t-s)A}\,\mathrm{d}W(s)=\sum_{i}^{\infty}\int_{0}^{t}e^{-(t-s)A}b_{i}\,\mathrm{d}W_{i}(s).$
By the Doob inequality and the Burkholder–Davis–Gundy inequality, for arbitrary p, $\widetilde{\delta}$, and $\varepsilon$, such that $p>1$, $\widetilde{\delta}>0$, and $\varepsilon>0$, we have the following estimation
$\mathbb{E}\left[\sup_{0\leq t\leq T}\|{{\bf{z}}}(t)\|_{D(A^{\widetilde{\delta}+\frac{1}{2}-\varepsilon})}^{p}\right]$ (78)
$=$ $\mathbb{E}\left[\sup_{0\leq t\leq T}\|A^{\widetilde{\delta}+\frac{1}{2}-\varepsilon}{{\bf{z}}}(t)\|_{L^{2}(\mathbb{T}^{3})}^{p}\right]$
$\leq$ $C(p)\mathbb{E}\left(\sum_{i=1}^{\infty}\gamma_{i}\int_{0}^{T}\|A^{\frac{1}{2}-\varepsilon}e^{-(T-s)A}A^{\widetilde{\delta}}b_{i}\|_{L^{2}(\mathbb{T}^{3})}^{2}\,\mathrm{d}s\right)^{\frac{p}{2}}$
$\leq$ $C(p)\mathbb{E}\left(\sum_{i=1}^{\infty}\gamma_{i}\int_{0}^{T}\|A^{\frac{1}{2}-\varepsilon}e^{-(T-s)A}\|^{2}_{\mathcal{B}(L^{2}(\mathbb{T}^{3}))}\|A^{\widetilde{\delta}}b_{i}\|_{L^{2}(\mathbb{T}^{3})}^{2}\,\mathrm{d}s\right)^{\frac{p}{2}}$
$\leq$ $C(p)\left(\sum_{i=1}^{\infty}\gamma_{i}\int_{0}^{T}\frac{1}{(T-s)^{1-2\varepsilon}}ds\right)^{\frac{p}{2}}$
[…]
${{\bf{z}}}(\omega)\in L^{\infty}(0,T;D(A^{{\widetilde{\delta}}+\frac{1}{2}-\varepsilon}))\mbox{ for $\displaystyle\mathbb{P}$- a.s. $\displaystyle\omega$}.$ (79)
Now we can set $\widetilde{\delta}={1}/{4}+\beta$ for some $\beta>\varepsilon>0$, which then together with the Sobolev embedding implies that
$D(A^{{\widetilde{\delta}}+\frac{1}{2}-\varepsilon})=D(A^{\frac{3}{4}+\beta-\varepsilon})\subset H^{\frac{3}{2}+2\beta-2\varepsilon}(\mathbb{T}^{3})\subset L^{\infty}(\mathbb{T}^{3}).$
Combining this with (79), we conclude that (13) holds whenever (77) is satisfied. ∎
With Lemma 12 established, it remains to verify that the noise assumption given in (8) indeed implies (77), thereby completing the proof of Lemma 1. From (8) and (7) we have
$W(t)\in C^{\beta}\left([0,T];D(A^{\delta})\right)\quad\mathbb{P}-a.s.,$ (80)
for some $\delta>3/4$ and every $\beta\in(0,1/2)$. Hence, Hence, there exist parameters $\beta$ and $\varepsilon$ such that $\beta>\varepsilon>0$, $\beta\geq{1}/{2}-\varepsilon$ and $\delta\geq 1/4+\beta$; for example we may choose $\varepsilon$ to be $1/4$ and $\beta$ to be in $(1/4,1/2)$. With such a choice, (80) implies (77), and therefore Lemma 12 applies under assumption (8). This completes the proof of Lemma 1.
Appendix B Energy equality for the approximation sequence
For the reader’s convenience, we present detailed calculations leading to the global energy estimates for the approximating sequence of the modified Ericksen–Leslie equations. These computations motivate the strategy for obtaining uniform estimates independent of the approximation parameter (see the proof of Lemma 3).
Lemma 13 (Energy equality for the approximating equations).
Fix $\omega\in\Omega$ such that ${{\bf{z}}}(\omega)\in L_{\rm loc}^{\infty}(\mathbb{T}^{3}\times[0,\infty))$. Consider the triple $({{\bf{v}}}^{\sigma}(\omega),{\pi}^{\sigma}(\omega),{{\bf{d}}}^{\sigma}(\omega))$, a classical strong (in the PDE sense) solution to (17), then we have the following equality:
$\frac{{\,\mathrm{d}}}{{\,\mathrm{d}}t}\int_{\mathbb{T}^{3}}\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}^{\sigma}|^{2}}{2}+F({{\bf{d}}}^{\sigma})\right)\varphi+\int_{\mathbb{T}^{3}}\left(|\nabla{{\bf{v}}}^{\sigma}|^{2}+|\Delta{{\bf{d}}}^{\sigma}|^{2}+|f({{\bf{d}}})^{\sigma}|^{2}\right)\varphi$ (81)
$=$ $\int_{\mathbb{T}^{3}}\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}^{\sigma}|^{2}}{2}+F({{\bf{d}}}^{\sigma})\right)\partial_{t}\varphi+\int_{\mathbb{T}^{3}}[({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\cdot\nabla{{{\bf{v}}}}^{\sigma}]\cdot{{{\bf{z}}}}\varphi$
$+\int_{\mathbb{T}^{3}}\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}+{{\bf{v}}}^{\sigma}\cdot{{\bf{z}}}\right)({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\cdot\nabla\varphi+{\pi}^{\sigma}{{\bf{v}}}^{\sigma}\cdot\nabla\varphi+\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}^{\sigma}|^{2}}{2}\right)\Delta\varphi$
$+\int_{\mathbb{T}^{3}}\left(\nabla{{\bf{d}}}^{\sigma}\odot\nabla{{\bf{d}}}^{\sigma}-\frac{1}{2}|\nabla{{\bf{d}}}^{\sigma}|^{2}{\rm I}_{3}\right):\nabla^{2}\varphi$
$+\int_{\mathbb{T}^{3}}({\bf{z}}\cdot\nabla)\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma}))\varphi+\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}}^{\sigma})\cdot\nabla{{\bf{d}}}^{\sigma}]\cdot(\nabla\varphi\cdot\nabla){{\bf{d}}}^{\sigma}$
$-\int_{\mathbb{T}^{3}}f({{\bf{d}}}^{\sigma})\cdot(\nabla\varphi\cdot\nabla){{\bf{d}}}^{\sigma}{-}2\int_{\mathbb{T}^{3}}\nabla f({{\bf{d}}}^{\sigma}):\nabla{{\bf{d}}}^{\sigma}\varphi.$
Proof.
We first note that all the subsequent calculations are valid as there exists a classical solution to (17) for every fixed $\sigma$, according to (18).
We take the inner product of (17)1 with respect to ${{\bf{v}}}^{\sigma}\varphi$, and proceed to analyze each term separately. We begin with the following:
$\int_{\mathbb{T}^{3}}\partial_{t}{{\bf{v}}}^{\sigma}\cdot{{\bf{v}}}^{\sigma}\varphi=\frac{\,\mathrm{d}}{{\,\mathrm{d}t}}\int_{\mathbb{T}^{3}}\frac{|{{\bf{v}}^{\sigma}}|^{2}}{2}\varphi-\int_{\mathbb{T}^{3}}\frac{|{{\bf{v}}^{\sigma}}|^{2}}{2}\partial_{t}\varphi.$
For the second term using integration by parts we have
$\displaystyle\int_{\mathbb{T}^{3}}-\Delta{{\bf{v}}}^{\sigma}\cdot{{\bf{v}}}^{\sigma}\varphi$ $\displaystyle=\int_{\mathbb{T}^{3}}\nabla{{\bf{v}}}^{\sigma}:\nabla({{\bf{v}}}^{\sigma}\varphi)$
$\displaystyle=\int_{\mathbb{T}^{3}}|\nabla{{\bf{v}}}^{\sigma}|^{2}\varphi+\int_{\mathbb{T}^{3}}\nabla{{\bf{v}}}^{\sigma}:({{\bf{v}}}^{\sigma}\otimes\nabla\varphi)$
$\displaystyle=\int_{\mathbb{T}^{3}}|\nabla{{\bf{v}}}^{\sigma}|^{2}\varphi+\int_{\mathbb{T}^{3}}\nabla\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}\right)\cdot\nabla\varphi$
$\displaystyle=\int_{\mathbb{T}^{3}}|\nabla{{\bf{v}}}^{\sigma}|^{2}\varphi-\int_{\mathbb{T}^{3}}\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}\Delta\varphi.$
For the third term, we have
$\displaystyle\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}}^{\sigma})\cdot\nabla({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])]\cdot{{\bf{v}}}^{\sigma}\varphi$
$\displaystyle=$ $\displaystyle-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}}^{\sigma})\cdot({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])]\cdot{{\bf{v}}}^{\sigma}\nabla\varphi-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}}^{\sigma})\cdot({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])]\cdot\nabla{{\bf{v}}}^{\sigma}\varphi$
$\displaystyle=$ $\displaystyle-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}}^{\sigma})\cdot({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])]\cdot{{\bf{v}}}^{\sigma}\nabla\varphi-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])]\cdot\nabla{{\bf{v}}}^{\sigma}\cdot{{\bf{z}}}\varphi$
$\displaystyle-\int_{\mathbb{T}^{3}}[{{\bf{v}}}^{\sigma}\cdot({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])]\cdot\nabla{{\bf{v}}}^{\sigma}\varphi$
$\displaystyle=$ $\displaystyle-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])]\cdot\nabla{{\bf{v}}}^{\sigma}\cdot{{\bf{z}}}\varphi-\int_{\mathbb{T}^{3}}[{{\bf{z}}}\cdot({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])]\cdot{{\bf{v}}}^{\sigma}\nabla\varphi$
$\displaystyle-\int_{\mathbb{T}^{3}}[{{\bf{v}}}^{\sigma}\cdot({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])]\cdot{{\bf{v}}}^{\sigma}\nabla\varphi-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])]\cdot\nabla\left(\dfrac{|{{\bf{v}}}^{\sigma}|^{2}}{2}\right)\varphi$
$\displaystyle=$ $\displaystyle-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])]\cdot\nabla{{\bf{v}}}^{\sigma}\cdot{{\bf{z}}}\varphi-\int_{\mathbb{T}^{3}}[{{\bf{z}}}\cdot({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])]\cdot{{\bf{v}}}^{\sigma}\nabla\varphi$
$\displaystyle-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])]\cdot\left(\dfrac{|{{\bf{v}}}^{\sigma}|^{2}}{2}\right)\nabla\varphi$
$\displaystyle=$ $\displaystyle-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])]\cdot\nabla{{\bf{v}}}^{\sigma}\cdot{{\bf{z}}}\varphi-\int_{\mathbb{T}^{3}}\left[{{\bf{z}}}\cdot{{\bf{v}}}^{\sigma}+\left(\dfrac{|{{\bf{v}}}^{\sigma}|^{2}}{2}\right)\right]({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\nabla\varphi$
Next, we have
$\int_{\mathbb{T}^{3}}\nabla{\pi}^{\sigma}\cdot{{\bf{v}}}^{\sigma}\varphi=-\int_{\mathbb{T}^{3}}{\pi}^{\sigma}{{\bf{v}}}^{\sigma}\cdot\nabla\varphi.$
For the term on the right-hand side, we have:
$\displaystyle\int_{\mathbb{T}^{3}}[{{\bf{v}}}^{\sigma}\cdot\nabla]\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]\cdot(-\Delta{{\bf{d}}}^{\sigma}+f({{\bf{d}}}^{\sigma}))\varphi$
$\displaystyle=$ $\displaystyle\underbrace{-\int_{\mathbb{T}^{3}}[{{\bf{v}}}^{\sigma}\cdot\nabla]\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]\cdot\Delta{{\bf{d}}}^{\sigma}\varphi}_{\mathcal{I}}\underbrace{+\int_{\mathbb{T}^{3}}[{{\bf{v}}}^{\sigma}\cdot\nabla]\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]\cdot f({{\bf{d}}}^{\sigma})\varphi}_{\mathcal{II}}.$ (82)
Differentiating (17)3 gives
$\partial_{t}\nabla{{\bf{d}}}^{\sigma}+\nabla\left[({{\bf{z}}}+{{\bf{v}}}^{\sigma})\cdot\nabla\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]\right]=\Delta(\nabla{{\bf{d}}}^{\sigma})-\nabla(f({{\bf{d}}}^{\sigma})).$ (83)
We take the inner product of (83) with $\nabla{{\bf{d}}}^{\sigma}\varphi$, and proceed to analyze each term separately. We begin with the following:
$\int_{\mathbb{T}^{3}}\partial_{t}(\nabla{{\bf{d}}}^{\sigma}):\nabla{{\bf{d}}}^{\sigma}\varphi=\frac{\,\mathrm{d}}{{\,\mathrm{d}t}}\int_{\mathbb{T}^{3}}\frac{1}{2}|\nabla{{\bf{d}}}^{\sigma}|^{2}\varphi-\int_{\mathbb{T}^{3}}\frac{1}{2}|\nabla{{\bf{d}}}^{\sigma}|^{2}\partial_{t}\varphi.$
Secondly, we have
$\displaystyle\int_{\mathbb{T}^{3}}\nabla[({{\bf{z}}}+{{\bf{v}}}^{\sigma})\cdot\nabla\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]]:\nabla{{\bf{d}}}^{\sigma}\varphi$
$\displaystyle=$ $\displaystyle\int_{\mathbb{T}^{3}}-[({{\bf{z}}}+{{\bf{v}}}^{\sigma})\cdot\nabla]\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]\cdot\Delta{{\bf{d}}}^{\sigma}\varphi-[({{\bf{z}}}+{{\bf{v}}}^{\sigma})\cdot\nabla]\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]\cdot[\nabla\varphi\cdot\nabla]{{\bf{d}}}^{\sigma}$
$\displaystyle=$ $\displaystyle-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}}^{\sigma})\cdot\nabla]\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]\cdot\Delta{{\bf{d}}}^{\sigma}\varphi-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}}^{\sigma})\cdot\nabla]\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]\cdot[\nabla\varphi\cdot\nabla]{{\bf{d}}}^{\sigma}$
$\displaystyle=$ $\displaystyle\underbrace{-\int_{\mathbb{T}^{3}}[{{\bf{z}}}\cdot\nabla]\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]\cdot\Delta{{\bf{d}}}^{\sigma}\varphi}_{\mathcal{III}}\underbrace{-\int_{\mathbb{T}^{3}}[{{\bf{v}}}^{\sigma}\cdot\nabla]\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]\cdot\Delta{{\bf{d}}}^{\sigma}\varphi}_{\mathcal{IV}}$
$\displaystyle-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}}^{\sigma})\cdot\nabla]\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]\cdot[\nabla\varphi\cdot\nabla]{{\bf{d}}}^{\sigma}.$ (84)
Finally, for the right-hand side term, we obtain the following after applying integration by parts:
$\int_{\mathbb{T}^{3}}\nabla(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})):\nabla{{\bf{d}}}^{\sigma}\varphi$ (85)
$=$ $-\int_{\mathbb{T}^{3}}|\Delta{{\bf{d}}}^{\sigma}|^{2}\varphi-\int_{\mathbb{T}^{3}}\Delta{{\bf{d}}}^{\sigma}\cdot(\nabla\varphi\cdot\nabla{{\bf{d}}}^{\sigma})-\int_{\mathbb{T}^{3}}\nabla f({{\bf{d}}}^{\sigma}):\nabla{{\bf{d}}}^{\sigma}\varphi$
$=$ $-\int_{\mathbb{T}^{3}}|\Delta{{\bf{d}}}^{\sigma}|^{2}\varphi+\int_{\mathbb{T}^{3}}\frac{1}{2}|\nabla{{\bf{d}}}^{\sigma}|^{2}\Delta\varphi+\int_{\mathbb{T}^{3}}\left(\nabla{{\bf{d}}}^{\sigma}\odot\nabla{{\bf{d}}}^{\sigma}-\frac{1}{2}|\nabla{{\bf{d}}}^{\sigma}|^{2}{\rm I}_{3}\right):\nabla^{2}\varphi$
$-\int_{\mathbb{T}^{3}}\nabla f({{\bf{d}}}^{\sigma}):\nabla{{\bf{d}}}^{\sigma}\varphi.$
We now return to (17)3 and take the inner product of both sides with $f({\bf{d}}^{\sigma})\varphi$. This yields:
$\displaystyle\int_{\mathbb{T}^{3}}\partial_{t}{{\bf{d}}}^{\sigma}\cdot f({{\bf{d}}}^{\sigma})\varphi$ $\displaystyle=\frac{\,\mathrm{d}}{{\,\mathrm{d}t}}\int_{\mathbb{T}^{3}}F({{\bf{d}}}^{\sigma})\varphi-\int_{\mathbb{T}^{3}}F({{\bf{d}}}^{\sigma})\partial_{t}\varphi,$
$\displaystyle\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}}^{\sigma})\cdot\nabla\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]]\cdot f({{\bf{d}}}^{\sigma})\varphi$ $\displaystyle=\underbrace{\int_{\mathbb{T}^{3}}({\bf{z}}\cdot\nabla\Phi_{\sigma}[{{\bf{d}}}^{\sigma}])\cdot f({\bf{d}}^{\sigma})\varphi}_{\mathcal{V}}\underbrace{+\int_{\mathbb{T}^{3}}({\bf{v}}\cdot\nabla\Phi_{\sigma}[{{\bf{d}}}^{\sigma}])\cdot f({\bf{d}}^{\sigma})\varphi}_{\mathcal{VI}},$
$\displaystyle\int_{\mathbb{T}^{3}}(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma}))\cdot f({{\bf{d}}}^{\sigma})\varphi$ $\displaystyle=-\int_{\mathbb{T}^{3}}\nabla{{\bf{d}}}^{\sigma}:\nabla f({{\bf{d}}}^{\sigma})\varphi+f({{\bf{d}}}^{\sigma})\cdot(\nabla\varphi\cdot\nabla{{\bf{d}}}^{\sigma})+|f({{\bf{d}}}^{\sigma})|^{2}\varphi.$
Combining all the above identities, we obtain (81); particularly, we utilize the following cancellations
$\displaystyle\mathcal{I}-\mathcal{IV}=0,$
$\displaystyle\mathcal{II}-\mathcal{VI}=0.$
Moreover, note that we have
$\mathcal{III}+\mathcal{V}=-\int_{\mathbb{T}^{3}}({\bf{z}}\cdot\nabla)\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma}))\varphi$
which then yields the seventh term on the right-hand-side of (81). ∎
We remark that the above computations remain valid if we choose $\varphi$ to be a constant function and the same calculation will result in the following global energy equality:
Corollary 14.
Under the same assumptions as in Lemma 13, it holds that
$\frac{\,\mathrm{d}}{{\,\mathrm{d}t}}\int_{\mathbb{T}^{3}}\frac{1}{2}(|{{\bf{v}}}^{\sigma}|^{2}+|\nabla{{\bf{d}}}^{\sigma}|^{2})+F({{\bf{d}}}^{\sigma}){\,\mathrm{d}x}+\int_{\mathbb{T}^{3}}|\nabla{{\bf{v}}}^{\sigma}|^{2}{\,\mathrm{d}x}$ (86)
$+\int_{\mathbb{T}^{3}}|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})|^{2}{\,\mathrm{d}x}$
$=$ $\int_{\mathbb{T}^{3}}[({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\cdot\nabla{{\bf{v}}}^{\sigma}]\cdot{{\bf{z}}}{\,\mathrm{d}x}+\int_{\mathbb{T}^{3}}{{\bf{z}}}\cdot\nabla\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})){\,\mathrm{d}x}.$
Appendix C Details of derivations of estimations for the approximation sequence
In this section, we present the detailed steps leading from (24) to (25), as part of the proof of Lemma 3. This exposition is included to facilitate the reader’s understanding of the underlying computations.
Firstly, observing that $\|\nabla^{2}{{\bf{d}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}=\|\Delta{{\bf{d}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}$, from (24) we have
$\frac{\,\mathrm{d}}{{\,\mathrm{d}t}}\int_{\mathbb{T}^{3}}\frac{1}{2}(|{{\bf{v}}}^{\sigma}|^{2}+|\nabla{{\bf{d}}}^{\sigma}|^{2})+F({{\bf{d}}}^{\sigma}){\,\mathrm{d}}x+|\nabla{{\bf{v}}}^{\sigma}|^{2}_{L^{2}(\mathbb{T}^{3})}$ (87)
$+\dfrac{1}{8}\int_{\mathbb{T}^{3}}|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})|^{2}{\,\mathrm{d}x}$
$+C\int_{\mathbb{T}^{3}}[|\Delta{{\bf{d}}}^{\sigma}|^{2}+|f({{\bf{d}}}^{\sigma})|^{2}+|\nabla{{\bf{d}}}^{\sigma}|^{2}|{{\bf{d}}}^{\sigma}|^{2}+|(\nabla{{\bf{d}}}^{\sigma})^{\top}{{\bf{d}}}^{\sigma}|^{2}]{\,\mathrm{d}x}$
$\leq C\left(\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}+{\dfrac{1}{2}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}}\right)+C\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}$
$\quad+\dfrac{6}{4}\int_{\mathbb{T}^{3}}|\nabla{{\bf{d}}}^{\sigma}|^{2}{\,\mathrm{d}x}.$
We define
$\displaystyle U(t):=\int_{\mathbb{T}^{3}}\frac{1}{2}(|{{\bf{v}}}^{\sigma}(t)|^{2}+|\nabla{{\bf{d}}}^{\sigma}(t)|^{2})+F({{\bf{d}}}^{\sigma}(t)){\,\mathrm{d}}x,$ (88)
Combining this with (87) and the fact that $F({{\bf{d}}}^{\sigma}(t))$ is nonnegative for every $t$, we obtain
$\frac{\,\mathrm{d}}{{\,\mathrm{d}t}}U(t)$ $\leq C\left(\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}+\dfrac{1}{2}\right)U(t)+C\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4},$ (89)
which implies that
$\frac{\,\mathrm{d}}{{\,\mathrm{d}t}}U(t)-C\left(\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}+\dfrac{1}{2}\right)U(t)$ $\leq C\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}.$ (90)
Multiplying both sides of the above inequality by
$\exp\left({-C\int^{t}_{0}\left(\|{{\bf{z}}}(s)\|_{L^{4}(\mathbb{T}^{3})}^{4}+\dfrac{1}{2}\right)ds}\right),$
we obtain
$\frac{\,\mathrm{d}}{{\,\mathrm{d}t}}\left(e^{{-C\int^{t}_{0}\left(\|{{\bf{z}}}(s)\|_{L^{4}(\mathbb{T}^{3})}^{4}+\frac{1}{2}\right)ds}}U(t)\right)$ $\leq Ce^{{-C\int^{t}_{0}(\|{{\bf{z}}}(s)\|_{L^{4}(\mathbb{T}^{3})}^{4}+\frac{1}{2})ds}}\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}\leq C\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}.$
Integrating both sides of the above inequality from 0 to t, for any $t\in(0,T]$, and then taking the supremum over $[0,T]$, we have
$\sup_{0\leq t\leq T}U(t)$ $\leq e^{{C\int^{T}_{0}\left(\|{{\bf{z}}}(s)\|_{L^{4}(\mathbb{T}^{3})}^{4}+\frac{1}{2}\right)ds}}\left(U(0)+C\int_{0}^{T}\|{{\bf{z}}}(t)\|_{L^{4}(\mathbb{T}^{3})}^{4}dt\right)$ (91)
Returning to (87), we integrate both sides over the interval $[0,t]$ and subsequently take the supremum over $t\in[0,T]$, which yields
$\sup_{t\in[0,T]}U(t)+\int_{0}^{T}\left(|\nabla{{\bf{v}}}^{\sigma}(t)|^{2}_{L^{2}(\mathbb{T}^{3})}+\dfrac{1}{8}\int_{\mathbb{T}^{3}}|\Delta{{\bf{d}}}^{\sigma}(t)-f({{\bf{d}}}^{\sigma}(t))|^{2}{\,\mathrm{d}x}\right)\,\mathrm{d}t$ (92)
$+C\int_{0}^{T}\int_{\mathbb{T}^{3}}[|\Delta{{\bf{d}}}^{\sigma}(t)|^{2}+|f({{\bf{d}}}^{\sigma}(t))|^{2}+|\nabla{{\bf{d}}}^{\sigma}(t)|^{2}|{{\bf{d}}}^{\sigma}(t)|^{2}]{\,\mathrm{d}x}\,\mathrm{d}t$
$+C\int_{0}^{T}\int_{\mathbb{T}^{3}}|(\nabla{{\bf{d}}}^{\sigma}(t))^{\top}{{\bf{d}}}^{\sigma}(t)|^{2}{\,\mathrm{d}x}\,\mathrm{d}t$
$\leq C\int_{0}^{T}\int_{\mathbb{T}^{3}}\left(\|{{\bf{z}}}(t)\|_{L^{4}(\mathbb{T}^{3})}^{4}+{\dfrac{1}{2}\|{{\bf{v}}}^{\sigma}(t)\|_{L^{2}(\mathbb{T}^{3})}^{2}}\right){\,\mathrm{d}x}\,\mathrm{d}t$
$+C\int_{0}^{T}\left(\|{{\bf{z}}}(t)\|_{L^{4}(\mathbb{T}^{3})}^{4}\|{{\bf{v}}}(t)^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+\dfrac{6}{4}\int_{\mathbb{T}^{3}}|\nabla{{\bf{d}}}^{\sigma}(t)|^{2}{\,\mathrm{d}x}\right)\,\mathrm{d}t.$
Combining this with (91), we obtain
$\int_{0}^{T}|\nabla{{\bf{v}}}^{\sigma}(t)|^{2}_{L^{2}(\mathbb{T}^{3})}\,\mathrm{d}t+C\int_{0}^{T}\int_{\mathbb{T}^{3}}|\Delta{{\bf{d}}}^{\sigma}(t)|^{2}{\,\mathrm{d}x}\,\mathrm{d}t$ (93)
$\leq C\int_{0}^{T}\left(\|{{\bf{z}}}(t)\|_{L^{4}(\mathbb{T}^{3})}^{4}+{\dfrac{1}{2}\|{{\bf{v}}}^{\sigma}(t)\|_{L^{2}(\mathbb{T}^{3})}^{2}}\right)\,\mathrm{d}t$
$\quad+C\int_{0}^{T}\left(\|{{\bf{z}}}(t)\|_{L^{4}(\mathbb{T}^{3})}^{4}\|{{\bf{v}}}(t)^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+\dfrac{6}{4}\int_{\mathbb{T}^{3}}|\nabla{{\bf{d}}}^{\sigma}(t)|^{2}{\,\mathrm{d}x}\right)\,\mathrm{d}t$
$\leq C\int_{0}^{T}\|{{\bf{z}}}(t)\|_{L^{4}(\mathbb{T}^{3})}^{4}\,\mathrm{d}t+CT\sup_{0\leq t\leq T}U(t)+C\sup_{0\leq t\leq T}U(t)\int_{0}^{T}\|{{\bf{z}}}(t)\|_{L^{4}(\mathbb{T}^{3})}^{4}\,\mathrm{d}t$
$\leq C\|{{\bf{z}}}(t)\|_{L^{4}(Q_{T})}^{4}+e^{{C\int^{T}_{0}\left(\|{{\bf{z}}}(s)\|_{L^{4}(\mathbb{T}^{3})}^{4}+\frac{1}{2}\right)ds}}\left(U(0)+C\|{{\bf{z}}}(t)\|_{L^{4}(Q_{T})}^{4}dt\right),$
which then implies (25) as desired.